### v1

# Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing

Abstract
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products, e.g., mRNA vaccines, in pharmaceutical manufacturing. While extensive efforts have been dedicated to shift the pharmaceutical industry towards continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode. This article proposes the first mechanistic model for a complete continuous lyophilization process, which includes freezing, primary drying, and secondary drying. The state-of-the-art lyophilization technology is considered, in which vials are suspended and moved continuously through the process. The model can describe the evolution of several critical process parameters, namely the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire lyophilization process. The model is also demonstrated for several applications related to process design and optimization. Ultimately, the framework and results presented in this work can serve as a solid foundation to guide the design and development of future continuous lyophilization processes.
**footnotetext: Corresponding author. Email: braatz@mit.edu
Keywords— Lyophilization, Freeze Drying, Continuous Manufacturing, Continuous Lyophilization, Pharmaceutical Manufacturing, Suspended Vials
Introduction
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products in pharmaceutical manufacturing [1]. Various promising applications of lyophilization have been proven, including its recent applications that provide long-term stability for mRNA vaccines (e.g., for COVID-19) to allow for higher storage temperature [2, 3]. Advancement in lyophilization technology could therefore play a crucial role in the future of the pharmaceutical industry. A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying. During freezing, the product and liquid (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with a small amount of liquid (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals. In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum. Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
Although the current trends in the pharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in a batch mode [4, 5], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [6, 5]. Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs. A number of continuous lyophilization concepts for pharmaceutical products have been proposed, including spray freeze drying [7] and spin freezing [4] (see a comprehensive review of continuous lyophilization technologies in Ref. [5]). A novel lyophilization technology has been recently proposed by Ref. [6], in which vials are suspended and continuously moved along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes. Since the phenomena in lyophilization are underpinned by several theories in the area of heat and mass transfer, models for lyophilization are mostly physics-based rather than data-driven. Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [8, 9, 10, 11, 12], primary drying [13, 14, 15, 16, 17, 8, 18, 19, 20, 21, 22], and secondary drying [13, 14, 15, 16, 17, 23, 21, 24, 25]. Only a few models for continuous lyophilization are available. For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by Ref. [26]. A model for the freezing step was also developed for spin freezing [27]. For spray freeze drying, Ref. [28] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases. Nevertheless, there is no model for the suspended-vial configuration. Besides, none of the published models for continuous lyophilization considers a complete lyophilization cycle (all three steps), which is critical for optimization and control of the entire continuous operation.
This article presents the first mechanistic model for continuous lyophilization of suspended vials. The model is developed to capture important phenomena of the entire lyophilization process, including cooling and stochastic/controlled ice nucleation during the freezing step, heat and mass transfer with sublimation during the primary drying step, and heat transfer with desorption during the secondary drying step. The proposed model is validated and employed to predict the evolution of important process parameters, with several applications associated with process design and optimization demonstrated.
This article is organized as follows. Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies. Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature. Section 4 validates the proposed model with the experimental data from the relevant literature. Section 5 showcases four case studies to demonstrate various applications of our model. Finally, Section 6 summarizes the study and briefly discusses possible future work.
Process Description
This work considers lyophilization of unit doses, in which the product is introduced into a vial prior to being lyophilized. This type of lyophilization, compared to lyophilization of bulk material, is more preferable in (bio)pharmaceutical applications due to its accurate dosage and better control of sterility [5]. Therefore, results and discussion presented in this article are based entirely on lyophilization of unit doses.
Batch lyophilization
The majority of lyophilization processes in the pharmaceutical industry has been operated in a batch mode for many decades. The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig. 1A). The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps.
Figure 1: (A) Conventional batch lyophilization of unit doses. A number of vials are placed on the cooling/heating shelf. (B) Continuous lyophilization of suspended vials. A number of vials are suspended and moved continuously along the lyophilizer.
Several drawbacks associated with batch lyophilization are extensively discussed in Refs. [6, 5]. Typical issues of any batch process include batch-to-batch variability, quality control, process downtime, and operational flexibility. Another disadvantage that is specific to the configuration illustrated in Fig. 1A is non-uniformity in heat transfer. For example, vials located on the side and corner of the shelf are typically affected by thermal radiation from the chamber walls, whereas the effect is almost negligible for center vials. Variability in heat transfer conditions could result in different crystal structures and drying times, hence variation in the final product quality. Consequently, process control and optimization are not straightforward. These known issues motivate the development of a continuous lyophilization concept that can be practically implemented in an industrial scale.
Continuous lyophilization
A comprehensive review of continuous lyophilization technologies is given in Ref. [5]. To summarize, the current continuous lyophilization technologies can be divided into four main categories. The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30]. Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with high surface-to-volume ratio, greatly improving heat and mass transfer in the process [7, 28]. Both aforementioned approaches are proposed for lyophilization of bulk material. For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [4].
Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by Ref. [6]. This technology employs the suspended-vial configuration as illustrated in Fig. 1B, where all vials are suspended and continuously moved through the freezing and drying chambers, without any contact between the vials and shelf. This suspended-vial technology has several benefits over the existing ones. First, heat transfer in the equipment is uniform because every vial experiences the exact same heat transfer conditions along the way. In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation. Second, the process does not involve with complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control. Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications. Finally, with the automatic filling and load-lock systems, the process is fully continuous. Therefore, our mechanistic model developed in this work is based on this suspended-vial configuration. We refer to Ref. [6] for detailed information and discussion related to the equipment design.
Mechanistic Modeling
This section develops a mechanistic model for continuous lyophilization based on the suspended-vial configuration (see Fig. 1B for a simple schematic and Ref. [6] for detailed discussion). The section starts by discussing the existing models in the literature and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
Review of the existing models
Mechanistic modeling for conventional batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry. Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials. Although there are some previous studies briefly discussing the literature on mechanistic modeling of lyophilization [18, 21], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation. Hence, this section firstly gives a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes the optimal model strategies to be used. Note that published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
Figure 2: Modeling strategies for the freezing, primary drying, and secondary drying steps in lyophilization. The strategies used in this article are highlighted.
Figure 2 summarizes the modeling strategies for the three different steps in lyophilization, including freezing, primary drying, and secondary drying. Published models mostly focus on primary drying because it is recognized as the most time-consuming, hazardous, and expensive step, and thus the first target for optimization and improvement. A variety of modeling strategies are available, ranging from the simplest lumped capacity model to the high-fidelity multidimensional model. Models for secondary drying have also been well developed as the final product quality (i.e., moisture content) is governed by this step. The mechanistic modeling of freezing receives the least attention in the lyophilization literature due to its complicated behavior, including subcooling and stochastic ice nucleation. As a result, the number of available models for freezing is more limited compared to that of the drying models.
Modeling the freezing step is most complicated among the three steps of localization; the blue column in Fig. 2 highlights modeling strategies for freezing. Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights. A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31]. Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails subcooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point. A more common but complicated strategy is to consider the subcooling period and incorporate the heat transfer associated with ice nucleation as done in Ref. [8]. Recently, Ref. [11] introduced a state-of-the-art freezing model which incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization. The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [12]. The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the mass transfer resistance during sublimation in primary drying. While some recent studies have built mechanistic models with the goal of predicting pore size distribution [10, 32], how to best model the specific complex interacting phenomena remains an open research question.
Many models for primary drying are available in the literature (see the green column in Fig. 2). The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system. In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and the temperature gradient in the vertical direction is substantial. A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 8, 35]. By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [35], and so the model can be implemented easily. However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer. A clever way of incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref. [18]). Of course, there are a number of models considering detailed heat and mass transfer, both in 1D [14, 15] and in higher dimensions [16]. Nevertheless, implementing those high-fidelity model for practical applications, e.g., state estimation and control, is not easy due to its high complexity, computational time, and number of parameters [18, 21]. Besides, the accuracy of complex models is not significantly different from 1D or simplified models in most cases [21]. Consequently, these high-fidelity models are not commonly used. Note that it is widely acknowledged that the amount of liquid removed via sublimation during primary drying is much higher than that of desorption, and so the effect of desorption can be omitted [15, 16, 18, 21].
Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation. Several modeling strategies are available, including a lumped capacity model [1], 1D heat transfer [25], 1D heat and mass transfer [15], and multidimensional models [16]. In the study by Ref. [25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors. This observation is similar to the primary drying case, where a vertical thermal gradient is usually dominant.
Here we select modeling strategies for our continuous lyophilization model that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum. From the above discussion, the selected strategies are (1) freezing with nucleation and subcooling for the freezing step, (2) 1D heat transfer with mass transfer incorporated as a boundary condition for the primary drying step, and (3) 1D heat transfer with bound water desorption for the secondary drying step, which are also highlighted in Fig. 2. The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration. Sections 3.2, 3.3, and 3.4 describe the models for freezing, primary drying, and secondary drying, respectively. Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
Figure 3: Schematic diagram showing the mechanistic modeling of suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
Model for freezing
For the design proposed by Ref. [6], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate. Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [36]. With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf. Given that the vial size is not large and heat transfer at all surfaces is almost identical, it is reasonable to use a lumped capacity model for the freezing step [11].
Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification, and (5) final cooling.
3.2.1 Preconditioning
In our model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$. Preconditioning entails cooling the product in a vial such that the product temperature is uniform at the target value below the freezing point (subcooling). In terms of modeling, this is the simplest step because there is no phase change and nucleation, and so only sensible heat is important.
The energy balance equation for the liquid solution in a vial is
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}},$ (1)
where $T(t)$ is the temperature, $t$ is time, $m$ is the total mass, $C_{p}$ is the specific heat capacity (per mass), $Q$ is the total heat transfer between the vial surface and environment, the subscripts ‘w’ and ‘s’ refer to the water (solvent) and solid (solute) phases, and the subscripts ‘s1’, ‘s2’ and ‘s3’ denote the top, bottom, and side of the vial, respectively. The initial conditions for (1) are
$T(t_{0})=T_{0},$ (2)
where $T_{0}$ is the initial product temperature. During the freezing step in general, the amount of water changes with time due to phase transition, and so we denote $m_{\textrm{w0}}$ as the initial mass of water at $t_{0}$. On the other hand, the amount of solid does not change, meaning that $m_{\textrm{s}}$ is a constant. However, during the preconditioning step, there is no phase transition at all, so
$m_{\textrm{w}}(t_{0}\leq t\leq t_{\textrm{f1}})=m_{\textrm{w,0}}.$ (3)
In some cases, experimental data are reported as the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$, where the subscript ‘l’ denotes the liquid phase (solvent + solute). In such cases, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated from $V_{\textrm{l}}$ and the densities $\rho_{\textrm{s}}$, $\rho_{\textrm{w}}$ (see Appendix A for the relevant equations).
The next step is to define the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$, which could vary among systems. For example, based on the data reported in Ref. [6], the bottom shelf is used to cool the gas inside the chamber, which subsequently cools the vials, hence natural convection. Besides, there exists thermal radiation between the vials and chamber walls. Therefore, both natural convection and radiation effects are modeled in this case. Consequently, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are
$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{c}}-T),$ (4)
$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T),$ (5)
$Q_{\textrm{s3}}=h_{\textrm{s3}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{r}\mathcal{F}(T_{\textrm{c}}^{4}-T^{4}),$ (6)
where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, and $\mathcal{F}$ is the transfer factor. In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter. Here there are two important temperatures: chamber wall temperature $T_{\textrm{c}}$ and cold gas temperature $T_{\textrm{g}}$. The top surface (chamber wall) and local gas in that area are not cooled, so we assume the gas temperature is equal to the wall temperature $T_{\textrm{c}}$, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection. Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection. At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and thus $h_{\textrm{s3}}$ takes into account of the convection part only.
As mentioned before, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system. Another design proposed by Ref. [6] is to flush the cryogenic gas directly into the chamber. In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region. Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6).
3.2.2 Vacuum-induced surface freezing
At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until VISF completes at $t_{\textrm{f2}}$. The key idea of VISF is to reduce the total pressure of a system to evaporate the liquid solution, abruptly decreasing the temperature and promoting nucleation. A model for VISF can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref. [31]. First, consider the mass transfer part. The evaporation rate of water at the surface for nonvolatile solute is
$\frac{dm_{\textrm{w}}}{dt}=h_{m}A_{z}(x_{\textrm{w,sat}}-x_{\textrm{w,c}}),$ (7)
where $h_{m}$ is the mass transfer coefficient, $x_{\textrm{w,sat}}$ is the mole fraction of water at the liquid-vapor interface (equilibrium), and $x_{\textrm{w,c}}$ is the mole fraction of water in the chamber (environment). The subscripts ‘sat’ and ‘c’ here denote the equilibrium condition and environment, respectively. To calculate the mass fraction of water for (7), we first assume that there are two gas/vapor components during VISF, namely (1) water and (2) nitrogen or inert gas, denoted by the subscript ‘in’. By assuming the ideal gas law,
$x_{\textrm{w,sat}}=\frac{p_{\textrm{w,sat}}M_{\textrm{w}}}{p_{\textrm{w,sat}}M_{\textrm{w}}+(p_{\textrm{t}}-p_{\textrm{w,sat}})M_{\textrm{in}}},$ (8)
$x_{\textrm{w,c}}=\frac{p_{\textrm{w,c}}M_{\textrm{w}}}{p_{\textrm{w,c}}M_{\textrm{w}}+(p_{\textrm{t}}-p_{\textrm{w,c}})M_{\textrm{in}}},$ (9)
where $p_{\textrm{t}}$ is the total pressure, $p_{\textrm{w}}$ is the partial pressure of water, and $M$ is the molar mass. If the environment contains only nitrogen or inert gas, $p_{\textrm{w,c}}=0$. The saturation pressure $p_{\textrm{w,sat}}$ is a function of temperature, that is [37],
$p_{\textrm{w,sat}}=10^{3}\exp\!\left(\!16.3872-\frac{3885.7}{T-42.98}\right).$ (10)
The initial condition for (7) is simply the final temperature of the preconditioning step.
Next, consider the heat transfer part. The energy balance equation can be modified from (1) as
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}(t)+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (11)
where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization. At low temperature, $\Delta H_{\textrm{vap}}$ can be approximated by
$\Delta H_{\textrm{vap}}=2.257\mbox{$\times$}10^{6}\!\left(\frac{1-T/647.1}{1-373.15/647.1}\right)^{\!\!0.38}.$ (12)
In (11), $m_{\textrm{w}}(t)$ and $Q_{\textrm{s3}}(t)$ are time-dependent because the amount of water (and hence the side surface area $A_{r}$) changes with time. The initial condition for (11) is the final mass of water at the end of the preconditioning step, which is $m_{\textrm{w,0}}$.
Note that, if the VISF process is done properly, the amount of water vaporized is usually very small and so can be safely neglected.
3.2.3 Nucleation
The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$. Our strategy to treat the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref. [11]. In this work, we use the term nucleation to denote the first nucleation where the temperature of a subcooled liquid almost instantaneously increases to the equilibrium/freezing point, whereas the term solidification is used to denote the phase transition from liquid to solid after that first nucleation.
Consider the controlled nucleation case first. The energy balance during nucleation, assuming the process is instantaneous and adiabatic, is
$(T_{\textrm{f,l}}-T_{\textrm{n}})(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})=m_{\textrm{i,n}}\Delta H_{\textrm{fus}},$ (13)
where $T_{\textrm{f,l}}$ is the freezing point of the liquid solution, i.e., the product temperature after nucleation, $T_{\textrm{n}}$ is the nucleation temperature, i.e., the product temperature when nucleation starts, $m_{\textrm{i,n}}$ is the mass of ice formed immediately after nucleation, the subscript ‘i’ denotes the solid phase (ice), and $\Delta H_{\textrm{fus}}$ is the heat of fusion. With the presence of a non-volatile solute, the freezing-point depression is
$T_{\textrm{f,w}}-T_{\textrm{f,l}}=\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}-m_{\textrm{i,n}}}\right),$ (14)
where $T_{\textrm{f,l}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute. The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (13) and (14) simultaneously, both of which become the initials conditions for the next (solidification) step.
Finally, define $t_{\textrm{f3}}$ as the time when the nucleation process completes. Since the nucleation process is nearly instantaneous as explained by Ref. [11], $t_{\textrm{f3}}$ is set to $t_{\textrm{f2}}$, with the relations
$m_{\textrm{w}}(t_{\textrm{f3}})=m_{\textrm{w}}(t_{\textrm{f2}})-m_{\textrm{i,n}},$ (15)
$m_{\textrm{i}}(t_{\textrm{f2}})=0,$ (16)
$m_{\textrm{i}}(t_{\textrm{f3}})=m_{\textrm{i,n}},$ (17)
$T(t_{\textrm{f2}})=T_{\textrm{n}},$ (18)
$T(t_{\textrm{f3}})=T_{\textrm{f,l}}.$ (19)
For the uncontrolled nucleation case (no VISF), the entire calculation is Section 3.2.2 is skipped. By nature, ice nucleation is stochastic and interpreted as a Poisson process. The rate constant of the Poison process $\lambda$ is [32]
$\lambda=k_{\textrm{n}}(T_{\textrm{f,l}}-T)^{b_{\textrm{n}}}V_{\textrm{l}},$ (20)
where $k_{\textrm{n}}$ and $b_{\textrm{n}}$ are the nucleation kinetics parameters. Before nucleation occurs, there is no ice in the system, and hence $T_{\textrm{f,l}}$ can be calculated by
$T_{\textrm{f,w}}-T_{\textrm{f,l}}=\frac{K_{\textrm{f}}m_{\textrm{s}}}{M_{\textrm{s}}m_{\textrm{w}}}.$ (21)
The probability that the first nucleus is formed between time $t$ and $t+\Delta t$ is expressed by
$P=1-\exp(\lambda\Delta t).$ (22)
The probability defined by (22) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$. At this point, the exact same procedure described for the controlled nucleation case, i.e.,(13)–(19), is followed, in which $t_{\textrm{f2}}$ is replaced with $t_{\textrm{f1}}$.
3.2.4 Solidification
After the first nucleation completes, the solidification step starts at $t_{\textrm{f3}}$. During this step, the product temperature is equal to the freezing point of the solution, which follows the freezing-point depression relation
$T(t)=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}(t)}\right),$ (23)
The energy balance during the freezing step is
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}}+m_{\textrm{i}}(t)C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}(t)+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (24)
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (4), (5), and (6), respectively. During the solidification step, the amount of water and ice changes with time, which affects the volume and surface area of the product, and so $Q_{\textrm{s2}}(t)$ is time-dependent. The relation between the amount of water and ice is
$m_{\textrm{w}}(t)=m_{\textrm{w}}(t_{\textrm{f3}})-m_{\textrm{i}}.$ (25)
The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely (15), (17), and (19).
Finally, define $t_{\textrm{f4}}$ as the time when solidification completes. The criterion for complete solidification is defined as
$m_{\textrm{i}}(t_{\textrm{f4}})=0.95m_{\textrm{w}}(t_{\textrm{f3}}),$ (26)
where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [11].
3.2.5 Cooling
The final step is to ensure that the product temperature is at the desire value before starting the drying process. This cooling step starts at $t_{\textrm{f4}}$ and completes at $t_{\textrm{f5}}$. The model for the final cooling step is relatively simple, with the only energy balance equation
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}}+m_{\textrm{i}}C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}},$ (27)
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (4), (5), and (6), respectively. Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, taking the final conditions of the solidification step. The initial condition for (27) is the final temperature of the solidification step. It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (27) without any significant error.
By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature and amount of substances (e.g., water and ice) during the freezing
Model for primary drying
In conventional lyophilization where vials are placed on the heating shelf, there are three heat transfer mechanisms associated with the bottom surface, namely conduction, convection, and radiation, whereas only radiation dominates at the side and top surfaces [17]. For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf. As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3B). Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$. If the primary drying model is simulated following the freezing step, then set $t_{0}=t_{\textrm{f4}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
The governing equations for the primary drying step consists of (1) the energy balance in the frozen region and (2) the mass balance at the sublimation front. By assuming that the supplied heat is used in the frozen region only, the energy balance for the frozen region is
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=k_{\textrm{f}}\dfrac{\partial^{2}T}{\partial z^{2}}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad S(t)<z<H,$ (28)
where $T(z,t)$ is the temperature, $S(t)$ is the sublimation front/interface position, $k$ is the thermal conductivity, $\rho$ is the density, $C_{p}$ is the heat capacity, $V$ is the volume, $H$ is the height of the product, and the subscript ‘f’ denotes the frozen region. We refer to Appendix A for the calculations of some relevant parameters. The radiative heat transfer from the sidewall $Q_{\textrm{rad}}$ is
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\!\left(T_{\textrm{c}}^{4}-T^{4}\right),$ (29)
where $\mathcal{F}$ is the transfer factor, $\sigma$ is the Stefan-Boltzmann constant, $T_{\textrm{c}}$ is the chamber wall temperature, and $A_{r}=\pi dH$ is the side area of the product. We refer to Section 3.5 for the detailed derivation of (29). The mass balance of water at the sublimation front gives
$\frac{dS}{dt}=\frac{N_{\textrm{w}}}{\rho_{\textrm{f}}-\rho_{\textrm{e}}},$ (30)
where $N_{\textrm{w}}$ is the sublimation flux and $\rho_{\textrm{e}}$ is the effective density of the dried region above the sublimation front. The driving force for mass transfer at the sublimation interface is [17, 1, 22]
$N_{\textrm{w}}=\frac{p_{\textrm{w,sat}}-p_{\textrm{w,c}}}{R_{\textrm{p}}},$ (31)
where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance. The saturation pressure for sublimation can be is described by [22]
$p_{\textrm{w,sat}}=\exp\!\left(\frac{-6139.9}{T}+28.8912\right).$ (32)
The variation of the resistance normally follows the empirical expression [17, 21, 22]
$R_{\textrm{p}}=R_{\textrm{p0}}+\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S},$ (33)
where $R_{\textrm{p}}$ is the mass transfer resistance (aka cake resistance) and $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, $R_{\textrm{p2}}$ are the constants estimated from data.
The partial differential equation (PDE) (28) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H,$ (34)
where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient which combines the effects of thermal radiation and natural convection. At the top surface, the energy balance associated with sublimation and thermal radiation from the above shelf/wall is
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=k_{\textrm{f}}\dfrac{\partial T}{\partial z}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=S(t),$ (35)
where $\Delta H_{\textrm{sub}}$ is the heat of sublimation and $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial.
The initial conditions for (28) and (30) are
$T(z,t_{0})=T_{0},\qquad 0\leq z\leq H,$ (36)
$S(t_{0})=0.$ (37)
If the model is simulated following the freezing step, then set $T_{0}=T(t_{\textrm{f4}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone primary drying simulation.
The primary drying model is simulated until the interface position is equal to the height of the product $S(t)=H$, indicating that there is no frozen material left, which marks the end of the primary drying step $t_{\textrm{d1}}$. In some cases, there might be an additional heating period at the end of primary drying to adjust the temperature and ensure complete sublimation before starting the secondary drying stage. However, this period is trivial to model; only a simple heat equation is required, and so is not described it here.
Model for secondary drying
The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) as done for the primary drying model (Fig. 3C). The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$. If the secondary drying model is simulated following the primary drying step, then set $t_{0}=t_{\textrm{d2}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
The governing equations for the secondary drying step consists of (1) the energy balance in the dried region and (2) the desorption kinetics. The energy balance of the dried product is
$\rho_{\textrm{e}}C_{p,\textrm{e}}\dfrac{\partial T}{\partial t}=k_{\textrm{e}}\dfrac{\partial^{2}T}{\partial z^{2}}+\rho_{\textrm{d}}\Delta H_{\textrm{des}}\dfrac{\partial c_{\textrm{w}}}{\partial t}+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad 0\leq z\leq H,$ (38)
where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the bound water concentration (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the enthalpy of desorption, $Q_{\textrm{rad}}$ is as defined in (29), and the subscript ‘e’ denotes the effective properties considering both solid and gas in the pores. The desorption kinetics of bound water is described by
$\dfrac{\partial c_{\textrm{w}}}{\partial t}=k_{\textrm{d}}(c^{*}_{\textrm{w}}-c_{\textrm{w}}),$ (39)
where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{w}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [14, 15, 21]
$k_{\textrm{d}}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT},$ (40)
where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency) and $E_{\textrm{a}}$ is the activation energy. The above desorption kinetics (39) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [14, 38, 18, 23, 21]. Further simplification that is very common in the literature can be done by setting $c^{*}_{\textrm{w}}=0$. This simplification produces insignificant error as shown in Ref. [15] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [21].
The governing PDE (38) requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H.$ (41)
Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=0.$ (42)
The initial conditions for (38) and (39) are
$T(z,t_{0})=T_{0},\qquad 0\leq z\leq H,$ (43)
$c_{\textrm{w}}(z,t_{0})=c_{\textrm{w,0}},\qquad 0\leq z\leq H.$ (44)
If the model is simulated following the primary drying step, then set $T_{0}=T(z,t_{\textrm{d1}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone secondary drying simulation.
The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step $t_{\textrm{d2}}$.
Modeling heat transfer in suspended vials
Convection and radiation are two key heat transfer mechanisms in the suspended-vial configuration as discussed before. During the freezing step, there could be a combination of natural/forced convection and radiation, depending on how the cryogenic gas is fed to the system. In primary and secondary drying, only natural convection and radiation are important.
Convection can be simply modeled using Newton’s law of cooling. A heat transfer coefficient for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection). Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref. [21].
Thermal radiation is significantly more complicated in terms of modeling. The theories and model equations described below are mainly based on Refs. [31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization. Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; please refer to the aforementioned references for more details. Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively. In general, the net radiant energy receiving by surface 1 can be written as
$Q_{\textrm{rad}}=\sigma A_{1}\mathcal{F}(T_{2}^{4}-T_{1}^{4}),$ (45)
where $A$ is the surface area, $\varepsilon$ is the emissivity, $\mathcal{F}$ is the transfer factor, and the subscripts ‘1’ and ‘2’ denote surfaces 1 and 2, respectively. The transfer factor is dependent on geometry and material type (emissivity). For any two diffuse, gray surfaces that form an enclosure,
$Q_{\textrm{rad}}=\dfrac{\sigma(T_{2}^{4}-T_{1}^{4})}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (46)
where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 ($F_{1-2}=1$) and surface 2 is much larger than surface 1, (46) can be simplified as
$Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}(T^{4}_{2}-T^{4}_{1}).$ (47)
In (47), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Now consider the case when surface 1 is the vial and surface 2 is the chamber walls. In this case, (47) becomes
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}(T^{4}_{\textrm{c}}-T^{4})$ (48)
for the side surface and
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{z}(T^{4}_{\textrm{c}}-T^{4})$ (49)
for the top and bottom surfaces. Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$. The amount of radiative heat for the single-vial case represented by (48) and (49) defines the upper bound on the radiative heat that the vial can receive from the chamber walls. This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among vials. As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$. Therefore, (48) needs to be modified to account for interactions between vials, whereas (49) can be used directly because thermal radiation at the top and bottom surfaces is not influenced by the other vials.
Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials. In batch lyophilization, vials are arranged as an array consisting of many rows, which results in significant differences in heat transfer conditions between the outer and inner vials. In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls. A rigorous way of modeling this complicated radiation is known as the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref. [39]. In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by Ref. [6]. As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials. Therefore, instead of using the radiation network approach, we modify (48) to follow (45) as
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}(T^{4}_{\textrm{c}}-T^{4}),$ (50)
where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph. This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (46). This technique is similar to the hybrid approach presented in Ref. [39]. Note that this approximation is for the suspended-vial configuration, in which heat transfer between vials is not significant; otherwise, the radiation network approach is recommended.
When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; for the temperature of about 300K, the error is about 0.1% [31]. The linearized equation can then be combined with the convection part, with the overall heat transfer coefficient taking into account of both convection and radiation, e.g., (34).
[…]
Default model parameters
The proposed model developed in Sections 3.2, 3.3, and 3.4 require a large number of input parameters. This section defines the default parameter values for our model to be used in the subsequent sections in this work. These parameter values are either obtained from the literature or set based on the typical values used in lyophilization. This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively. Table LABEL:Tab:DefaultParameters shows the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
Table 1: Default model parameters for a complete continuous lyophilization simulation.
Symbol Value Unit Source
Thermophysical properties
[…]
$\Delta H_{\textrm{des}}$ 2.68$\times$106 J/kg [15]
$\Delta H_{\textrm{fus}}$ 3.34$\times$105 J/kg [37]
$\Delta H_{\textrm{sub}}$ 2.84$\times$106 J/kg [18]
$\Delta H_{\textrm{vap}}$ see (12) J/kg [37]
[…]
$\rho_{\textrm{e}}$ 215 kg/m3 [15]
$\rho_{\textrm{d}}$ 212.21 kg/m3 [15]
$\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix A)
$\rho_{\textrm{i}}$ 917 kg/m3 [32]
$\rho_{\textrm{s}}$ 1,587.9 kg/m3 [32]
$\rho_{\textrm{w}}$ 1,000 kg/m3 [12]
Operating conditions
$c_{\textrm{w,0}}$ 0.088 kg water/kg solid [6]
$c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [1]
$p_{\textrm{t}}$ see Fig. 5D Pa
$p_{\textrm{w,c}}$ 3 Pa [6]
$p_{\textrm{w,sat}}$ see (10) and (32) Pa [37, 22]
$T_{0}$ 298.15 K
$T_{\textrm{b}}$ 270 (primary drying) K
295 (secondary drying) K
$T_{\textrm{c}}$ 273 (before VISF) K
240 (after VISF) K
265 (primary drying) K
290 (secondary drying) K
$T_{\textrm{f,w}}$ 273.15 K [12]
$T_{\textrm{g}}$ 268 (before VISF) K
230 (after VISF) K
$T_{\textrm{n}}$ 268 K
Heat and mass transfer
$\mathcal{F}$ 0.78
$f_{\textrm{a}}$ 1.5$\times$10-3 1/s
[…]
$h_{\textrm{m}}$ 6.34$\times$10-3 kg/m2$\cdots$ [31]
[…]
$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [6]
$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [6]
[…]
Product and formulation
$H$ 7.2$\times$10-3 m calculated (see Appendix A)
$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix A)
$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix A)
$V_{\textrm{l}}$ 3$\times$10-6 m3 [6]
$x_{\textrm{s}}$ 0.05 [6]
Vial
$d$ 0.024 m 10R vial
$\epsilon_{\textrm{gl}}$ 0.8 [31]
Others
[…]
$M_{\textrm{in}}$ 28$\times$10-3 kg/mol [40]
$M_{\textrm{s}}$ 0.3423 kg/mol [12]
$M_{\textrm{w}}$ 18$\times$10-3 kg/mol [40]
[…]
Parameters that should be estimated from experimental data include heat transfer coefficients, mass transfer coefficients and cake resistance, and desorption-related parameters (frequency factor and activation energy).
Implementation
In this article, the mechanistic model was implemented in MATLAB R2023a, and so all results and plots were generated using MATLAB. A system of ordinary differential equations (ODEs) was integrated with ode15s. Partial differential equations (PDEs) were discretized spatially to produce a system of ODEs, which was then fed to the ODE solver (see Appendix B for the detailed numerical methods).
Model validation
The mechanistic model presented in Section 3 is validated using the experimental data reported in Ref. [6]. The model is validated for each lyophilization step separately, with the parameters reported in Table 2. The default values defined in Table LABEL:Tab:DefaultParameters are used for parameters not shown in Table 2.
Table 2: Parameters used for the model validation, including Cases 1, 2a, 2b, 3a, and 3b.
Symbol Value Unit Source
Freezing (no VISF)
$T_{0}$ 280 K [6]
$T_{\textrm{c}}$ 272 K estimated from data in Ref. [6]
$T_{\textrm{g}}$ see Fig. 4A K [6]
$T_{\textrm{n}}$ 263.18 K [6]
[…]
Primary drying
$T_{0}$ 231 (Case 2a) K [6]
235 (Case 2b) K [6]
$T_{\textrm{b}}$ 263 (Case 2a) K [6]
313 (Case 2b) K [6]
$T_{\textrm{b}}$ 275 K estimated from data in Ref. [6]
[…]
$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [6]
$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [6]
Secondary drying
$c_{\textrm{w,0}}$ 0.088 (Case 3a) kg water/kg solid [6]
0.075 (Case 3b) kg water/kg solid [6]
$T_{0}$ 273 K [6]
$T_{\textrm{b}}$ 293 K [6]
$T_{\textrm{c}}$ 285 K estimated from data in Ref. [6]
$f_{\textrm{a}}$ 0.42 1/s estimated from data in Ref. [6]
[…]
$V_{\textrm{l}}$ 2$\times$10-6 m3 [6]
Figure 4: Model validation for the lyophilization of suspended vials using the experimental data from Ref. [6]. Panel A shows the model prediction and experimental data for the product temperature during the freezing step. Panel B shows the model prediction and experimental data for the product temperature (assumed to be measured at the bottom surface) during the primary drying step. The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively. Panel C shows the model prediction and experimental data for the bound water concentration during the secondary drying step. The initial concentrations for Cases 3a and 3b are 0.0864 and 0.0739 kg water/kg solid, respectively.
For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig. 4A). During the preconditioning phase, the temperature decreases from its initial value of about 280 K to the nucleation temperature of about 263 K, where nucleation starts. At the first nucleation, the temperature instantaneously increases from 263 K to about the equilibrium freezing point of about 272 K. The temperature slowly decreases during the solidification phase, which takes about 30 min, following the freezing-point depression. Subsequently, the product temperature decreases and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
The model can accurately predict the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig. 4B). The product temperature increases fast at the beginning because the driving force for sublimation is low, and so most of the heat input is used to increase the temperature. As time progresses, the driving force for sublimation is higher, resulting in most of the heat used for sublimation instead of increasing the product temperature. The maximum error in temperature prediction is about 2–3 K, while that of drying time prediction is about 0.2 for Case 2a and 0.4 for Case 2b ($\sim$ 6%). This error could be partly because of the uncertainty of cake resistance, which is caused by a stochastic freezing process.
In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.0864 and 0.0739 kg water/kg solid). The concentration decreases exponentially following the linear driving force model, where the simulations are terminated when the concentration reaches 0.01 kg water/kg solid. It is important to note that the data used in Fig. 4C are based on the shelf temperature of 293 K. To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures such that the effects of temperature on the desorption process can be quantified.
Case Studies
The previous sections detail the development and validation of our mechanistic model for lyophilization. This section demonstrates various applications of our model via four different case studies.
Process design and optimization
The first case study focuses on model-based design and optimization of the entire continuous lyophilization process. One of the critical aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production. In the context of continuous lyophilization, our model can be used to study the evolution of critical process parameters, e.g., temperature, moisture content, throughout the process at various operating conditions, which can help guide the process design and optimization.
This work not only presents the first mechanistic model for continuous lyophilization, but also is one of the very few studies that proposes a complete model that includes all three steps in lyophilization, including freezing, primary drying, and secondary drying. This is important specifically for continuous manufacturing where the entire process should be designed and optimized simultaneously.
By using the default parameters in Table LABEL:Tab:DefaultParameters, we simulate a complete continuous lyophilization cycle including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig. 5). The evolution of the product temperature (Fig. 5A), mass of ice (Fig. 5B), and bound water concentration (Fig. 5C) can be obtained, given the operating pressure (Fig. 5D) and temperature (Fig. 5E). This information could be useful for various purposes. For example, it can guide the process and equipment design concerning chamber sizing and vial moving speed to accommodate the residence time of each step. Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time. It is important to note that the entire simulation shown in Fig. 5 can be computed in less than 1 second on a normal laptop, which makes the model very practical to be implemented for any purposes.
Figure 5: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) concentration of bound water, (D) operating pressure, and (E) operating temperature.
Studying vacuum-induced surface freezing
Our model can be simulated for the whole lyophilization process as shown in Section 5.1 or simulated separately for each step in the lyophilization cycle to provide insights into some part of the process. This section employs the model to study the freezing step, which usually receives little attention compared to the drying steps.
The proposed model can simulate different freezing protocols, namely uncontrolled (aka spontaneous) and controlled nucleation. Spontaneous nucleation was already investigate in detail by Refs. [11, 12], and so here we consider the controlled nucleation case via VISF, which is preferable for continuous lyophilization. As an example, the evolution of the product temperature and amount of water evaporating during VISF at different pressure profiles is simulated (Fig. 6). Varying among different systems, this information could be used to design a freezing protocol, e.g., pressure profile, that achieves the desired nucleation temperature.
Figure 6: Simulation results showing the evolution of the product during VISF at different pressure profiles. Panel A shows the pressure profiles that decrease linearly from $10^{5}$ Pa to $10^{4}$ Pa, 1,000, and 100 Pa, respectively. Panel B shows the evolution of the product temperature. Panel C shows the total mass of water, which implies the evaporation rate. External heat transfer with the wall and cryogenic gas is excluded so that only the contribution from VISF is considered.
Model modifications for conventional drying
Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization. The major modifications are to turn off the thermal radiation part in the model and recalculate/estimate the heat transfer coefficient and cake resistance. In this case study, we consider batch primary drying and secondary drying using the data from Refs. [18] and [21], with the parameters reported in Table 3.
Table 3: Parameters used for the conventional drying studies, including for Cases 4a, 4b, and 5.
Symbol Value Unit Source
Primary drying
[…]
$\rho_{\textrm{e}}$ 252 kg/m3 [18]
$\rho_{\textrm{f}}$ 917 kg/m3 [18]
$p_{\textrm{w,c}}$ 15 Pa [18]
$T_{0}$ 228.15 K [18]
$T_{\textrm{b}}$ see1 K [18]
$\mathcal{F}$ 0 turn off the radiation part
[…]
$R_{\textrm{p0}}$ 3.8$\times$104 m/s estimated from data in Ref. [18]
$R_{\textrm{p1}}$ 3.1$\times$107 1/s estimated from data in Ref. [18]
$R_{\textrm{p2}}$ 10 1/m estimated from data in Ref. [18]
$H$ 7.15$\times$10-3 m [18]
Secondary drying
$c_{\textrm{w,0}}$ 0.0603 kg water/kg solid [21]
$T_{0}$ 273 K [21]
$T_{\textrm{b}}$ see2 K estimated from data in Ref. [21]
$f_{\textrm{a}}$ 1.2$\times$10-3 1/s estimated from data in Ref. [21]
[…]
$\mathcal{F}$ 0 turn off the radiation part
[…]
$V_{\textrm{l}}$ 1.5$\times$10-6 m3 [21]
$d$ 0.01425 m 2R vial, [21]
  • 1
    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 (Case 4b).
  • 2
    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264.09 K to 312 K for Case 5.
Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig. 7). The model can also effectively predict the product temperature and bound water concentration during the secondary drying step (Fig. 8). With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
Figure 7: Comparison between our model prediction and the experimental/simulation data from Ref. [18] for the primary drying step in conventional lyophilization. Panel A shows the model prediction and the experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and reference simulation data for the sublimation front position. The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15, while other conditions are identical.
Figure 8: Comparison between our model prediction and the experimental data from Ref. [21] for the secondary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and experimental data for the concentration of bound water.
Condenser failure analysis
Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily at its desired conditions. A better process model should be able to simulate important abnormal conditions, e.g., equipment failure, that could occur due to various unexpected scenarios, which is critical in continuous manufacturing as these abnormal operations could affect the reliability, availability, and maintainability (RAM) as well as the safety of the plant/process.
In lyophilization, one of the crucial design considerations is associated with a condenser. During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1]. This is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase. This scenario may not lead to safety-related issues because the operating pressure in lyophilization is usually low and the amount of ice/water is not high enough to create overpressure. Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (31)), which then prolongs the primary drying stage. As a result, the final product quality could be affected. This section explores how to incorporate those dynamic behaviors into our model.
Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption. To incorporate the condenser and dynamics of vapor flow into the model, consider the mass balance equation for water vapor in a chamber [22]
$\frac{dp_{\textrm{w,c}}}{dt}=\frac{(j_{\textrm{w}}-j_{\textrm{w,max}})R\overline{T}}{V_{\textrm{c}}M_{\textrm{w}}},$ (51)
where $j_{\textrm{w}}$ is the total mass flow rate of water vapor, $j_{\textrm{w,max}}$ is the maximum condenser capacity, $V_{\textrm{c}}$ is the volume of a chamber, and $T_{\textrm{avg}}$ is the average temperature in the chamber assumed to be constant for simplification. The mass flow rate of water vapor can be calculated by
$j_{m,\textrm{w}}=n_{\textrm{vial}}A_{z}N_{w},$ (52)
where $n_{\textrm{vial}}$ is the number of vials in the chamber. By coupling (51) and (52) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
Here we consider a situation where there is some issue with the condenser, and so the total condenser capacity cannot handle the water vapor produced via sublimation during primary drying, with the parameters given in Table 4. The simulation result shows that there is a pressure increase from the normal operating value of 3 Pa to 20 Pa (Fig. 9A). This pressure increase results from the total vapor flow that exceeds the maximum condenser capacity, leading vapor accumulation in the chamber. When the pressure increases, the driving force for sublimation (also the sublimation flux) decreases. Finally, when the sublimation flux balances with the condenser capacity, the system pressure becomes constant. It is also shown that the drying time increases by about an hour (Fig. 9B), and the product temperature is higher (Fig. 9C).
Table 4: Parameters used for the condenser failure analysis.
Symbol Value Unit Source
$T_{0}$ 228 K
$T_{\textrm{b}}$ 270 K
$\overline{T}$ 260 K assumed to be constant
$n_{\textrm{vial}}$ 200
$V_{\textrm{c}}$ 0.118 m3 [22]
$j_{m,\textrm{max}}$ 1.8$\times$10-5 m3 [22]
Figure 9: Simulation results comparing the normal operation and condenser failure scenario for the (A) system pressure, (B) sublimation front position, and (C) product average temperature during primary drying.
Conclusion
This work proposes the first mechanistic model for the suspended-vial continuous lyophilization technology, in which vials containing the pharmaceutical products are suspended and moved continuously along the process. The model can be used to simulate the entire lyophilization process, including the freezing, primary drying, and secondary drying steps. The freezing model considers preconditioning, vacuum-induced surface freezing (VISF), spontaneous and controlled nucleation, solidification, and final cooling. The primary drying model captures heat transfer in the frozen region and ice sublimation driven by mass transfer at the sublimation front. The secondary drying model describes simultaneous heat transfer in the dried region and bound water desorption. The overall model can describe the evolution of product temperature, ice/water fraction, and residual moisture throughout the process.
Our model is validated using experimental data from the literature, in which the model predictions are consistent with the experimental measurements in all cases. The validated model is applied for four case studies that entail continuous lyophilization process design and optimization, detailed investigation of VISF, model modifications for conventional drying, and condenser failure analysis. The framework and results presented in this work are promising and could be used to guide the design and development of future continuous lyophilization technology.
With a well-developed mechanistic model, future work could consider utilizing the model for several crucial tasks, e.g., uncertainty quantification, state estimation, and control design, which will ultimately lead to the development of a high-quality digital twin for continuous lyophilization that can greatly facilitate the process design and development in this area.
Data Availability
All software and data will be made available upon publication of the manuscript.
Acknowledgements
This research was supported by the U.S. Food and Drug Administration under the FDA BAA-22-00123 program, Award Number 75F40122C00200.
Appendices
Appendix A Supplemental equations
This appendix summarizes additional equations that are not crucial part of the model but required for calculating some model parameters.
In Section 3.2.1, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
$\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (A.1)
Consequently, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated by
$m_{\textrm{s}}=x_{\textrm{s}}\rho_{\textrm{l}}V_{\textrm{l}},$ (A.2)
$m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (A.3)
Then, (1) can be used as usual.
During the freezing step, the surface area $A_{r}$ changes with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
$A_{r}(t)=\frac{4\!\left(\dfrac{m_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{m_{\textrm{w}}(t)}{\rho_{\textrm{w}}}+\dfrac{m_{\textrm{i}}(t)}{\rho_{\textrm{i}}}\right)}{d},$ (A.4)
where the product is assumed to be a cylinder.
For the primary drying model (Section 3.3), the volume of the frozen region is
$V_{\textrm{f}}=\pi d^{2}(H-S).$ (A.5)
The properties of the frozen region (subscript ‘f’) can be calculated from those of the ice and solid material or solute as follows:
$\rho_{\textrm{f}}=\frac{1}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{i}}}},$ (A.6)
$C_{p,\textrm{f}}=x_{\textrm{s}}C_{p,\textrm{s}}+(1-x_{\textrm{s}})C_{p,\textrm{i}},$ (A.7)
$k_{\textrm{f}}=x_{\textrm{s}}k_{\textrm{s}}+(1-x_{\textrm{s}})k_{\textrm{i}}.$ (A.8)
Here we assume that the final mass of ice is equal to the initial mass of water (complete freezing), meaning that the amount of water loss due to evaporation during VISF and unfrozen water (bound water) is negligible, which is generally true.
Finally, the height/thickness of the product $H$ can be calculated by
$H=\frac{(m_{\textrm{s}}+m_{\textrm{w,0}})}{\rho_{\textrm{f}}A_{z}}.$ (A.9)
Note that, for some experiments that focus on primary drying, $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and $H$ are reported, and so the above calculations may not be needed.
For the secondary drying model (Section 3.4), the volume of the dried region is
$V_{\textrm{e}}=\pi d^{2}H,$ (A.10)
which is constant because there is no moving boundary in this case.
Appendix B Numerical methods
This appendix comprehensively describes numerical techniques required for solving the model equations and summarizes the final equations implemented in MATLAB.
For the freezing model, since a lumped thermal capacity approximation is used, the resulting equations are all ODEs, and so the equations can be solved directly with the selected ODE solver, which is ode15s in this case.
For the primary drying model, the PDE (28) is discretized spatially to produce a system of ODEs, which is then fed to the ODE solver; the technique is referred to as the method of lines [41]. Since the PDE (28) involves a moving interface $S(t)$, define a new variable
$\xi=\frac{z-S}{H-S},$ (B.1)
where $\xi$ represents is dimensionless position with respect to the moving interface (aka sublimation front). This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$. Consequently, the PDE (28) and boundary conditions (34) and (35) become
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (B.2)
$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (B.3)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad\xi=0.$ (B.4)
To spatially discretize the PDE, define
$\Delta\xi=\frac{1}{n_{z}-1},$ (B.5)
$\xi=j\Delta\xi,$ (B.6)
where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$. Discretizing (B.2)–(B.4) using the second-order finite difference scheme results in
$\begin{split}\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(j\Delta\xi-1)}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (B.7)
$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.8)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.9)
where $T_{j}$ is the product temperature at position $j$. The ghost point at $j=n_{z}+1$ in (B.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=n_{z}$. Similarly, $T_{j=0}$ in (B.9) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=1$. The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system, which is then integrated using ode15s.
Numerical treatment for the secondary drying model are simpler than those described for the primary drying model as there is no moving interface. Define
$\Delta z=\frac{H}{n_{z}-1},$ (B.10)
$z=j\Delta z,$ (B.11)
where $\Delta z$ is the distance between each grid point. Discretizing (38), (39), (41), and (42) using the second-order finite difference scheme results in
$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (B.12)
$\dfrac{dc_{\textrm{w},j}}{dt}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (B.13)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.14)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.15)
where the ghost points can be treated similarly as done for the primary drying model.
Table 5: *
Nomenclature
Symbols
$A$ area (m2)
$A_{z}$ cross sectional area of the product (m2)
$A_{r}$ side surface area of the product (m2)
$b_{\textrm{n}}$ nucleation kinetics parameter (–)
[…]
$c_{\textrm{w}}$ concentration of bound water (kg water/kg solid)
$c^{*}_{\textrm{w}}$ equilibrium concentration of bound water (kg water/kg solid)
$d$ vial diameter (m)
[…]
$\mathcal{F}$ transfer factor (–)
$F_{1-2}$ view factor from surface 1 to surface 2 (–)
$f_{\textrm{a}}$ frequency factor (1/s)
$H$ height of the product (m)
$\Delta H_{\textrm{des}}$ heat of desorption (J/kg)
$\Delta H_{\textrm{fus}}$ heat of fusion (J/kg)
$\Delta H_{\textrm{sub}}$ heat of sublimation (J/kg)
$\Delta H_{\textrm{vap}}$ heat of vaporization (J/kg)
[…]
$h_{m}$ mass transfer coefficient (kg/m2$\cdots$)
$j$ integer index for discretization (–)
$j_{\textrm{w}}$ mass flow rate of water vapor (kg/s)
$j_{\textrm{w,max}}$ maximum mass flow rate of a condenser (kg/s)
[…]
$k_{\textrm{d}}$ desorption constant (1/s)
[…]
$M$ molar mass (kg/mol)
$m$ mass (kg)
$m_{\textrm{i,n}}$ mass of ice resulting from the first nucleation (kg)
[…]
$n_{\textrm{vial}}$ number of vials (–)
$n_{z}$ number of grid points in the $z$-direction (–)
$P$ probability
$p$ partial pressure (Pa)
$p_{\textrm{t}}$ total pressure (Pa)
$Q$ heat transfer rate (W)
$Q_{\textrm{rad}}$ radiative heat transfer rate (W)
$q$ heat flux (W/m2)
[…]
$R_{\textrm{p}}$ mass transfer resistance of the product during sublimation (m/s)
$R_{\textrm{p0}}$ parameter for the mass transfer resistance (m/s)
$R_{\textrm{p1}}$ parameter for the mass transfer resistance (1/s)
$R_{\textrm{p2}}$ parameter for the mass transfer resistance (1/m)
$r$ radial position (m)
$S$ sublimation front position (m)
$T$ product temperature (K)
$T_{\textrm{b}}$ bottom shelf temperature (K)
$T_{\textrm{c}}$ chamber wall temperature (K)
$T_{\textrm{f}}$ freezing point (K)
$T_{\textrm{g}}$ cryogenic gas temperature (K)
$T_{\textrm{n}}$ nucleation temperature (K)
$\overline{T}$ average chamber temperature (K)
$t$ time (s)
$t_{\textrm{0}}$ initial time, (s)
$t_{\textrm{f1}}$ time when preconditioning completes (s)
$t_{\textrm{f2}}$ time when VISF completes (s)
$t_{\textrm{f3}}$ time when first nucleation completes (s)
$t_{\textrm{f4}}$ time when solidification completes (s)
$t_{\textrm{f5}}$ time when final cooling completes (s)
$t_{\textrm{d1}}$ time when primary drying completes (s)
$t_{\textrm{d2}}$ time when secondary drying completes (s)
$V$ volume (m3)
$x$ mass fraction (–)
$z$ position in vertical direction (m)
$\Delta z$ distance between each grid point in the $z$-direction (m)
$\alpha$ thermal diffusivity, $k/\rho C_{p}$ (m2/s)
$\varepsilon$ emissivity (–)
$\xi$ normalized vertical position (–)
$\Delta\xi$ distance between each grid point in the $\xi$-direction (–)
$\rho$ density (kg/m3)
$\rho_{\textrm{d}}$ density of the dried product (kg/m3)
[…]
Common subscripts
00 initial condition
$1$ surface 1
$2$ surface 2
avg average
c chamber (environment)
b bottom shelf
e effective properties of the dried product (solid + gas)
f frozen material, freezing
gl glass vial
i ice
l liquid solution (water + solute/solid)
n nucleation
s solute/solid
sat saturated condition
s1 top surface of the product
s2 bottom surface of the product
s3 side surface of the product
w water
in inert gas

### v2

@@
-Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing
+Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization
 Abstract
-Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products, e.g., mRNA vaccines, in pharmaceutical manufacturing.
-While extensive efforts have been dedicated to shift the pharmaceutical industry towards continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode.
-This article proposes the first mechanistic model for a complete continuous lyophilization process, which includes freezing, primary drying, and secondary drying.
-The state-of-the-art lyophilization technology is considered, in which vials are suspended and moved continuously through the process.
-The model can describe the evolution of several critical process parameters, namely the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire lyophilization process.
-The model is also demonstrated for several applications related to process design and optimization.
-Ultimately, the framework and results presented in this work can serve as a solid foundation to guide the design and development of future continuous lyophilization processes.
+Lyophilization (aka freeze drying) is a typical process in pharmaceutical manufacturing used for improving the stability of various drug products, including its recent applications to mRNA vaccines.
+While extensive efforts have been dedicated to shifting the pharmaceutical industry toward continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode.
+This article presents the first mechanistic model for a complete continuous lyophilization process, which comprehensively incorporates and describes key transport phenomena in all three steps of lyophilization, namely freezing, primary drying, and secondary drying.
+The proposed model considers the state-of-the-art lyophilization technology, in which vials are suspended and move continuously through the process.
+The validated model can accurately predict the evolution of critical process parameters, including the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire process.
+Several applications related to model-based process design and optimization of continuous lyophilization are also demonstrated.
+The final model is made available in MATLAB and Julia as an open-source software package called ContLyo, which can ultimately be leveraged for guiding the design and development of future continuous lyophilization processes.
 **footnotetext: Corresponding author. Email: braatz@mit.edu
-Keywords— Lyophilization, Freeze Drying, Continuous Manufacturing, Continuous Lyophilization, Pharmaceutical Manufacturing, Suspended Vials
+Keywords— Lyophilization, Freeze Drying, Suspended Vials, Continuous Manufacturing, Pharmaceuticals
 Introduction
@@
 A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying.
-During freezing, the product and liquid (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with a small amount of liquid (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals.
+During freezing, the product and liquid solvent (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with the remaining fraction (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals [4].
 In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum.
@@
 A novel lyophilization technology has been recently proposed by Ref.
-[6], in which vials are suspended and continuously moved along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
+[7], in which vials are suspended and continuously move along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
 Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes.
-Since the phenomena in lyophilization are underpinned by several theories in the area of heat and mass transfer, models for lyophilization are mostly physics-based rather than data-driven.
+Since the key phenomena in lyophilization are underpinned by heat and mass transfer theories, process models for lyophilization are mostly physics-based rather than data-driven.
 Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [9, 10, 11, 12, 13], primary drying [14, 15, 16, 17, 18, 9, 19, 20, 21, 4, 22], and secondary drying [14, 15, 16, 17, 18, 23, 4, 24, 25].
@@
 This article presents the first mechanistic model for continuous lyophilization of suspended vials.
-The model is developed to capture important phenomena of the entire lyophilization process, including cooling and stochastic/controlled ice nucleation during the freezing step, heat and mass transfer with sublimation during the primary drying step, and heat transfer with desorption during the secondary drying step.
-The proposed model is validated and employed to predict the evolution of important process parameters, with several applications associated with process design and optimization demonstrated.
+The model is developed to capture important transport phenomena in the process, including cooling and stochastic/controlled ice nucleation during the freezing step, sublimation of ice crystals during the primary drying step, and desorption of bound water during the secondary drying step.
+The proposed model is validated and employed to predict the evolution of critical process parameters for the entire lyophilization process, which is subsequently demonstrated for model-based design and optimization.
 This article is organized as follows.
@@
 Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature.
-Section 4 validates the proposed model with the experimental data from the relevant literature.
-Section 5 showcases four case studies to demonstrate various applications of our model.
+Section 4 details the model validation and showcases various applications of the validated model.
 Finally, Section 5 summarizes the study and briefly discusses possible future work.
@@
 Batch lyophilization
-The majority of lyophilization processes in the pharmaceutical industry has been operated in a batch mode for many decades.
+The majority of lyophilization processes in the pharmaceutical industry have been operated in a batch mode for many decades.
 The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig.
@@
 The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps.
+In this case, the freezing and drying processes take place in the same space but at different times.
 Figure 1: (A) Conventional batch lyophilization of unit doses.
@@
 (B) Continuous lyophilization of suspended vials.
-A number of vials are suspended and moved continuously along the lyophilizer.
+A number of vials are suspended and move continuously through the lyophilizer.
 Several drawbacks associated with batch lyophilization are extensively discussed in Refs.
@@
 The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30].
-Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with high surface-to-volume ratio, greatly improving heat and mass transfer in the process [7, 28].
-Both aforementioned approaches are proposed for lyophilization of bulk material.
+Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with a high surface-to-volume ratio, greatly improving heat and mass transfer in the process [8, 28].
+Both aforementioned approaches have been proposed for lyophilization of bulk material.
 For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [5].
@@
 This technology employs the suspended-vial configuration as illustrated in Fig.
-1B, where all vials are suspended and continuously moved through the freezing and drying chambers, without any contact between the vials and shelf.
+1B, where all vials are suspended and continuously move through the freezing and drying chambers, without any contact between the vials and shelf.
+Unlike conventional lyophilizers, the freezing and drying processes occur in different locations/chambers of the equipment.
 This suspended-vial technology has several benefits over the existing ones.
@@
 In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation.
-Second, the process does not involve with complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control.
+Second, the process does not involve any complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control.
 Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications.
 Finally, with the automatic filling and load-lock systems, the process is fully continuous.
-Therefore, our mechanistic model developed in this work is based on this suspended-vial configuration.
-We refer to Ref.
-[6] for detailed information and discussion related to the equipment design.
+The mechanistic model developed in this work is based on this suspended-vial configuration as illustrated in Fig.
+1B, the current state-of-the-art continuous lyophilization technology for unit doses.
+Detailed modeling strategies and model equations are discussed in the next section (Section 3), while information and discussion related to the equipment design can be found in Ref.
+[7].
 Mechanistic Modeling
-This section develops a mechanistic model for continuous lyophilization based on the suspended-vial configuration (see Fig.
-1B for a simple schematic and Ref.
-[6] for detailed discussion).
-The section starts by discussing the existing models in the literature and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
-Review of the existing models
-Mechanistic modeling for conventional batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry.
+This section details the development of a mechanistic model for continuous lyophilization of suspended vials (Fig.
+1B).
+The section starts by discussing existing models and modeling strategies in the literature, and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
+Review of the existing models and modeling strategies
+Mechanistic modeling for conventional/batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry.
 Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials.
-Although there are some previous studies briefly discussing the literature on mechanistic modeling of lyophilization [18, 21], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation.
-Hence, this section firstly gives a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes the optimal model strategies to be used.
-Note that published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
+While some past reviews briefly discuss the literature on mechanistic modeling of lyophilization [19, 4], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation.
+Hence, this section first provides a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes with the optimal modeling strategies used in this work.
+The published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
 Figure 2: Modeling strategies for the freezing, primary drying, and secondary drying steps in lyophilization. The strategies used in this article are highlighted.
-Figure 2 summarizes the modeling strategies for the three different steps in lyophilization, including freezing, primary drying, and secondary drying.
-Published models mostly focus on primary drying because it is recognized as the most time-consuming, hazardous, and expensive step, and thus the first target for optimization and improvement.
+Figure 2 summarizes various modeling strategies for the key steps of lyophilization, namely freezing, primary drying, and secondary drying.
+Published models mostly focus on primary drying because this step is recognized as the most time-consuming and expensive step, thus the first target for process optimization and improvement.
 A variety of modeling strategies are available, ranging from the simplest lumped capacity model to the high-fidelity multidimensional model.
-Models for secondary drying have also been well developed as the final product quality (i.e., moisture content) is governed by this step.
-The mechanistic modeling of freezing receives the least attention in the lyophilization literature due to its complicated behavior, including subcooling and stochastic ice nucleation.
+Models for secondary drying have also been well developed because the final product quality (i.e., moisture content) is governed by this step.
+The mechanistic modeling of freezing receives less attention in the lyophilization literature due to its complicated behavior, including supercooling and stochastic ice nucleation.
 As a result, the number of available models for freezing is more limited compared to that of the drying models.
-Modeling the freezing step is most complicated among the three steps of localization; the blue column in Fig.
-2 highlights modeling strategies for freezing.
+Modeling strategies for the freezing step
+Modeling the freezing step is most complicated among the three steps of lyophilization; the blue column in Fig.
+2 highlights notable modeling strategies for this step.
 Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights.
 A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31].
-Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails subcooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point.
-A more common but complicated strategy is to consider the subcooling period and incorporate the heat transfer associated with ice nucleation as done in Ref.
+Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails supercooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point [10, 12].
+A more common but more complicated strategy is to consider the supercooling period and incorporate the heat transfer associated with ice nucleation as done in Ref.
 [9].
 Recently, Ref.
-[11] introduced a state-of-the-art freezing model which incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization.
+[12] introduced a state-of-the-art freezing model that incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization using a lumped capacity model.
 The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [13].
-The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the mass transfer resistance during sublimation in primary drying.
+In Appendix A, we provide a detailed analysis using the relevant dimensionless group to compare between the lumped capacity model and model with thermal gradients and show that the lumped capacity model is sufficiently accurate for lyophilization of unit doses in general.
+The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the solid structure and mass transfer resistance during sublimation in primary drying.
 While some recent studies have built mechanistic models with the goal of predicting pore size distribution [11, 32], how to best model the specific complex interacting phenomena remains an open research question.
+Modeling strategies for the primary drying step
 Many models for primary drying are available in the literature (see the green column in Fig.
@@
 The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system.
-In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and the temperature gradient in the vertical direction is substantial.
+In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and hence the temperature gradient in the vertical direction is significant.
 A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 9, 35].
@@
 However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer.
-A clever way of incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref.
+An approach to incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front/interface between the frozen and dried regions, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref.
 [19]).
-Of course, there are a number of models considering detailed heat and mass transfer, both in 1D [14, 15] and in higher dimensions [16].
-Nevertheless, implementing those high-fidelity model for practical applications, e.g., state estimation and control, is not easy due to its high complexity, computational time, and number of parameters [18, 21].
-Besides, the accuracy of complex models is not significantly different from 1D or simplified models in most cases [21].
+In this case, the driving force for mass transfer is dependent on the saturation (equilibrium) pressure at the sublimation interface, which is a function of temperature.
+Hence, the model should be able to accurately predict the spatial variation of the product temperature, including the interface temperature, and hence modeling the heat transfer in the vertical direction is needed.
+Moreover, one of the key considerations during the drying steps is to ensure that the product temperature does not exceed the upper limit (e.g., collapse temperature and glass transition temperature) [1], and thus spatially distributed temperature data are valuable for accurately determining the maximum temperature in the product.
+There are models that consider detailed heat and mass transfer by including full energy and continuity equations, both in 1D [15, 16] and in higher dimensions [17].
+Implementing such high-fidelity models for practical applications, e.g., state estimation and control, is not easy due to their high complexity, computation time, and number of parameters [19, 4].
+Besides, the accuracy of the complex models for primary drying is not significantly different from 1D or simplified models in most cases [4].
 Consequently, these high-fidelity models are not commonly used.
-Note that it is widely acknowledged that the amount of liquid removed via sublimation during primary drying is much higher than that of desorption, and so the effect of desorption can be omitted [15, 16, 18, 21].
+Note that, in primary drying, the water vapor is removed via both sublimation (frozen region) and desorption (dried region).
+Nevertheless, it is widely acknowledged that the amount of water vapor removed via sublimation is much higher than that of desorption, and so the effect of desorption can be omitted [16, 17, 19, 4].
+Modeling strategies for the secondary drying step
 Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation.
-Several modeling strategies are available, including a lumped capacity model [1], 1D heat transfer [25], 1D heat and mass transfer [15], and multidimensional models [16].
-In the study by Ref.
+Several modeling strategies are available, ranging from the simplest lumped capacity model [1] to the high-fidelity model considering multidimensional heat and mass transfer in detail [17].
+In Ref.
 [25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors.
-This observation is similar to the primary drying case, where a vertical thermal gradient is usually dominant.
-Here we select modeling strategies for our continuous lyophilization model that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum.
-From the above discussion, the selected strategies are (1) freezing with nucleation and subcooling for the freezing step, (2) 1D heat transfer with mass transfer incorporated as a boundary condition for the primary drying step, and (3) 1D heat transfer with bound water desorption for the secondary drying step, which are also highlighted in Fig.
+Therefore, 1D modeling is preferable.
+Nevertheless, typical measurement techniques such as Karl Fischer titration measure the total amount of bound water, not the spatial variation.
+Consequently, a lumped capacity model is sometimes considered acceptable.
+In primary drying, the drying rate is mainly governed by the rate of sublimation, which makes the modeling typically more straightforward than for secondary drying.
+In secondary drying, the two important mechanisms are (1) vapor phase transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product.
+Some previous models for secondary drying consider both phenomena [14, 15, 16, 17], whereas the more recently published models tend to consider only the desorption part [23, 4, 24, 25, 36].
+Simulation results in these studies show that the model prediction is highly accurate even when the vapor phase transport is omitted, suggesting that desorption is the actual limiting step.
+In Ref.
+[37], a detailed experimental study was conducted, concluding that desorption is the rate-limiting step for mass transfer in secondary drying.
+Evidence from both simulation and experiment suggest that only the desorption part is necessary.
+Although it is evident from both simulation and experiment in the literature that desorption is the limiting step, there is no systematic analysis from the theoretical perspective of transport phenomena.
+In Appendix A, we provide a simple and systematic way of analyzing the effect/contribution of each transport process via scale analysis, which can be used to justify the contribution of each transport process for different systems/conditions other than those considered in this work and in the literature.
+Optimal modeling strategies
+High-quality mechanistic modeling should balance between the model accuracy and complexity such that the resulting model can provide reliable results and be practically implemented for different purposes with ease, e.g., optimization, state estimation, and model-based control.
+High-fidelity models should have the best accuracy in theory, but these models typically entail a large number of equations and parameters.
+Simulating these model equations in real time could be challenging.
+Besides, incomplete knowledge of those parameters and associated uncertainty could negatively impact the model accuracy instead of improving it.
+On the other hand, low-fidelity models are much simpler to implement, but the model prediction might not be sufficiently accurate.
+Our goal is to select modeling strategies that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum.
+From the discussion in Sections 3.1.13.1.3 and Appendix A, the selected modeling strategies are highlighted in Fig.
 2.
+The freezing model relies on the lumped capacity method that considers nucleation and supercooling.
+The primary drying model simulates heat transfer in 1D (vertical direction) with mass transfer incorporated as a boundary condition at the sublimation front.
+Finally, the secondary drying model captures heat transfer in 1D (vertical direction) and desorption of bound water.
 The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration.
@@
 Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
-Figure 3: Schematic diagram showing the mechanistic modeling of suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
+Figure 3: Schematic diagram showing the mechanistic modeling of continuous lyophilization via suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
 Model for freezing
@@
 With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf.
-Given that the vial size is not large and heat transfer at all surfaces is almost identical, it is reasonable to use a lumped capacity model for the freezing step [11].
-Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification, and (5) final cooling.
-3.2.1 Preconditioning
-In our model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$.
-Preconditioning entails cooling the product in a vial such that the product temperature is uniform at the target value below the freezing point (subcooling).
-In terms of modeling, this is the simplest step because there is no phase change and nucleation, and so only sensible heat is important.
+Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification or ice formation, and (5) final cooling.
+Preconditioning
+In this model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$.
+Preconditioning entails cooling the product in a vial such that the product temperature is uniform at the target value below the freezing point (supercooling).
+In terms of modeling, this step is the simplest because there is no phase change and nucleation, and so only sensible heat is important.
 The energy balance equation for the liquid solution in a vial is
@@
 where $T_{0}$ is the initial product temperature.
-During the freezing step in general, the amount of water changes with time due to phase transition, and so we denote $m_{\textrm{w0}}$ as the initial mass of water at $t_{0}$.
+During the freezing step in general, the amount of water changes with time due to phase transition, and so we denote $m_{\textrm{w,0}}$ as the initial mass of water at $t_{0}$.
 On the other hand, the amount of solid does not change, meaning that $m_{\textrm{s}}$ is a constant.
@@
 In some cases, experimental data are reported as the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$, where the subscript ‘l’ denotes the liquid phase (solvent + solute).
-In such cases, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated from $V_{\textrm{l}}$ and the densities $\rho_{\textrm{s}}$, $\rho_{\textrm{w}}$ (see Appendix A for the relevant equations).
+In such cases, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated from $V_{\textrm{l}}$ and the densities $\rho_{\textrm{s}}$, $\rho_{\textrm{w}}$ (see Appendix B for the relevant equations).
 The next step is to define the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$, which could vary among systems.
-For example, based on the data reported in Ref.
-[6], the bottom shelf is used to cool the gas inside the chamber, which subsequently cools the vials, hence natural convection.
-Besides, there exists thermal radiation between the vials and chamber walls.
-Therefore, both natural convection and radiation effects are modeled in this case.
-Consequently, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are
-$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{c}}-T),$ (4)
-$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T),$ (5)
-$Q_{\textrm{s3}}=h_{\textrm{s3}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{r}\mathcal{F}(T_{\textrm{c}}^{4}-T^{4}),$ (6)
-where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, and $\mathcal{F}$ is the transfer factor.
+Here, we start by writing general expressions and then simplify them to match our system of interest.
+In suspended vials, there exists thermal radiation between the vials, chamber walls, and heating/cooling shelves.
+Therefore,
+$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{z}\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),$ (4)
+$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{z}\mathcal{F}_{\textrm{s2}}(T_{\textrm{b}}^{4}-T^{4}),$ (5)
+$Q_{\textrm{s3}}=h_{\textrm{s3}}A_{r}(T_{\textrm{g}}-T)+\sigma A_{r}\mathcal{F}_{\textrm{s3}}(T_{\textrm{c}}^{4}-T^{4}),$ (6)
+where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, $\mathcal{F}$ is the transfer factor (see the definition in Section 3.5), and the subscripts ‘u’, ‘b’, ‘c’, and ‘g’ denote the upper surface of the chamber, bottom shelf, chamber walls, and cold gas, respectively.
 In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter.
-Here there are two important temperatures: chamber wall temperature $T_{\textrm{c}}$ and cold gas temperature $T_{\textrm{g}}$.
-The top surface (chamber wall) and local gas in that area are not cooled, so we assume the gas temperature is equal to the wall temperature $T_{\textrm{c}}$, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection.
+In lyophilization, the bottom shelf and gas temperatures ($T_{\textrm{b}}$ and $T_{\textrm{g}}$) typically vary with time and are specified by the freezing/drying protocol.
+The upper surface and wall temperatures ($T_{\textrm{u}}$ and $T_{\textrm{c}}$), however, are not usually measured and thus can be estimated from data [39].
+In this article, we assume that both temperatures are approximately constant as reported by Refs.
+[40, 19].
+In addition, $T_{\textrm{u}}\approx T_{\textrm{c}}$ due to no additional heat source/sink at these locations.
+Nevertheless, our model is not restricted to these assumptions; i.e., temperature-dependent data can be fed to the model (if available).
+Next, (4)–(6) are simplified further to match the experimental setup in Ref.
+[7], whose data are used for our model validation.
+In Ref.
+[7], the gas inside the chamber was cooled by the bottom shelf, and this cold gas was subsequently used to cool the vials, hence natural convection.
+In this setup, the upper surface and local gas in that area are not cooled, so the gas temperature is assumed to be equal to the surface temperature $T_{\textrm{u}}$.
+Consequently, the two terms can be combined and written in the form of Newton’s law of cooling, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection, which can be done by linearizing the fourth-order term of the radiation part (see more details in Section 1.3.2 of Ref.
+[31] and Section 3.5).
 Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection.
-At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and thus $h_{\textrm{s3}}$ takes into account of the convection part only.
-As mentioned before, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system.
+As a result, $Q_{\textrm{s1}}$ and $Q_{\textrm{s2}}$ are
+$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{u}}-T),$ (7)
+$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T).$ (8)
+At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and so no simplification is needed for (6).
+Note that the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system.
 Another design proposed by Ref.
@@
 In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region.
-Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6).
-3.2.2 Vacuum-induced surface freezing
-At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until VISF completes at $t_{\textrm{f2}}$.
-The key idea of VISF is to reduce the total pressure of a system to evaporate the liquid solution, abruptly decreasing the temperature and promoting nucleation.
-A model for VISF can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref.
+Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6) even further.
+Vacuum-induced surface freezing
+At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until its completion at $t_{\textrm{f2}}$.
+The key idea of VISF is to reduce the total pressure to evaporate a small amount of liquid solution from the product, abruptly decreasing the temperature and promoting nucleation.
+A model for VISF is not readily available in the literature but can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref.
 [31].
 First, consider the mass transfer part.
-The evaporation rate of water at the surface for nonvolatile solute is
-$\frac{dm_{\textrm{w}}}{dt}=h_{m}A_{z}(x_{\textrm{w,sat}}-x_{\textrm{w,c}}),$ (7)
+The evaporation rate of water at the surface for a nonvolatile solute is
+$\frac{dm_{\textrm{w}}}{dt}=-h_{m}A_{z}(x_{\textrm{w,sat}}-x_{\textrm{w,c}}),$ (9)
 where $h_{m}$ is the mass transfer coefficient, $x_{\textrm{w,sat}}$ is the mole fraction of water at the liquid-vapor interface (equilibrium), and $x_{\textrm{w,c}}$ is the mole fraction of water in the chamber (environment).
@@
 $p_{\textrm{w,sat}}=10^{3}\exp\!\left(\!16.3872-\frac{3885.7}{T-42.98}\right).$ (12)
-The initial condition for (7) is simply the final temperature of the preconditioning step.
+Note that, if the VISF process is done properly, the amount of water that vaporizes is usually very small and thus does not significantly impact the overall thermophysical properties.
 Next, consider the heat transfer part. The energy balance equation can be modified from (1) as
-$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}(t)+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (11)
-where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization.
-At low temperature, $\Delta H_{\textrm{vap}}$ can be approximated by
+$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (13)
+where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization, which can be approximated by [41]
 $\Delta H_{\textrm{vap}}=2.257\mbox{$\times$}10^{6}\!\left(\frac{1-T/647.1}{1-373.15/647.1}\right)^{\!\!0.38}.$ (14)
-In (11), $m_{\textrm{w}}(t)$ and $Q_{\textrm{s3}}(t)$ are time-dependent because the amount of water (and hence the side surface area $A_{r}$) changes with time.
-The initial condition for (11) is the final mass of water at the end of the preconditioning step, which is $m_{\textrm{w,0}}$.
-Note that, if the VISF process is done properly, the amount of water vaporized is usually very small and so can be safely neglected.
-3.2.3 Nucleation
+In (13), $m_{\textrm{w}}(t)$ is time-dependent due to evaporation, which also means that the side surface area $A_{r}$ changes with time.
+The initial conditions for (9) and (13) are the final mass of water and temperature at the end of the preconditioning step.
+Nucleation
 The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$.
-Our strategy to treat the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref.
+Our modeling strategy for the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref.
 [12].
-In this work, we use the term nucleation to denote the first nucleation where the temperature of a subcooled liquid almost instantaneously increases to the equilibrium/freezing point, whereas the term solidification is used to denote the phase transition from liquid to solid after that first nucleation.
-Consider the controlled nucleation case first. The energy balance during nucleation, assuming the process is instantaneous and adiabatic, is
+In this work, we use the term nucleation to denote the first nucleation where the temperature of supercooled liquid almost instantaneously increases to the equilibrium/freezing point, which is caused by the heat released from the fraction of liquid being frozen/solidified.
+On the other hand, we use the term solidification to denote the phase transition from liquid to solid (ice formation) after that first nucleation.
+First, consider the case of controlled nucleation with VISF. The energy balance during nucleation, assuming the process is instantaneous and adiabatic, is
 $(T_{\textrm{f,l}}-T_{\textrm{n}})(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})=m_{\textrm{i,n}}\Delta H_{\textrm{fus}},$ (15)
@@
 $T_{\textrm{f,w}}-T_{\textrm{f,l}}=\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}-m_{\textrm{i,n}}}\right),$ (16)
-where $T_{\textrm{f,l}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute.
-The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (13) and (14) simultaneously, both of which become the initials conditions for the next (solidification) step.
+where $T_{\textrm{f,w}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute.
+The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (15) and (16) simultaneously.
 Finally, define $t_{\textrm{f3}}$ as the time when the nucleation process completes.
@@
 $T(t_{\textrm{f3}})=T_{\textrm{f,l}}.$ (21)
-For the uncontrolled nucleation case (no VISF), the entire calculation is Section 3.2.2 is skipped.
-By nature, ice nucleation is stochastic and interpreted as a Poisson process.
-The rate constant of the Poison process $\lambda$ is [32]
+For the case of stochastic nucleation, the entire calculation described for VISF in Section 3.2.2 is skipped.
+By nature, ice nucleation is stochastic and can be interpreted as a Poisson process.
+The rate constant of the Poison process $\lambda$ is
 $\lambda=k_{\textrm{n}}(T_{\textrm{f,l}}-T)^{b_{\textrm{n}}}V_{\textrm{l}},$ (22)
@@
 The probability defined by (24) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$.
-At this point, the exact same procedure described for the controlled nucleation case, i.e.,(13)–(19), is followed, in which $t_{\textrm{f2}}$ is replaced with $t_{\textrm{f1}}$.
-3.2.4 Solidification
-After the first nucleation completes, the solidification step starts at $t_{\textrm{f3}}$.
-During this step, the product temperature is equal to the freezing point of the solution, which follows the freezing-point depression relation
-$T(t)=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}(t)}\right),$ (23)
-The energy balance during the freezing step is
-$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}}+m_{\textrm{i}}(t)C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}(t)+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (24)
+Subsequently, follow the exact same procedure described by (15)–(21), with $t_{\textrm{f1}}$ replacing $t_{\textrm{f2}}$.
+Solidification
+After the first nucleation completes, the solidification (aka ice formation) step starts at $t_{\textrm{f3}}$.
+For solidification, the lumped capacity model was shown to provide accurate prediction of the temperature and freezing time for typical vial sizes [12]; we also discuss this in the context of the Biot number in Appendix A.
+The lumped capacity model is sufficiently accurate and highly computationally efficient, but it does not represent the correct physics of ice formation.
+By treating the entire solid/liquid as a lumped object during solidification, it implies that homogeneous nucleation is assumed.
+Nevertheless, it is commonly known that heterogeneous nucleation is a more frequently observed phenomenon during ice formation [42].
+Consequently, the solidification process generally starts from the vial surfaces, i.e., the outer and bottom surfaces in this case.
+To take this fact into account, a rigorous 2D heat transfer model could be considered [13].
+In such cases, the physics is incorporated more accurately, but the major drawback of 2D modeling is its high computational cost and numerical complexity.
+Our strategy here is to combine the advantages of the lumped capacity and 2D models together, resulting in a model that is computationally efficient, provides accurate prediction, and represents the correct physics; we denote it as a hybrid lumped capacity model.
+In this case, ice formation is assumed to initiate from the side and bottom surfaces (similar to a 2D model), in which the ice layer is treated as an additional heat transfer resistance between the cold gas and unfrozen water.
+Additionally, heat conduction in the ice layer is assumed to be quasi steady, which is a reasonable assumption for problems associated with phase change in general as the sensible heat is much smaller than the latent heat [31].
+Physically, this is equivalent to assuming that all the heat input is used for phase transition only.
+The unfrozen water and solute are treated as a lumped object as its temperature follows the equilibrium temperature.
+A schematic diagram showing our model for the solidification step is shown in Fig.
+4
+Figure 4: Schematic diagram showing the mechanistic modeling of the solidification step.
+For simplification, it is assumed that the liquid part retains a cylindrical shape with the same aspect ratio as the initial solution before nucleation starts.
+The energy balance during the solidification step is
+$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (25)
 with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively.
-During the solidification step, the amount of water and ice changes with time, which affects the volume and surface area of the product, and so $Q_{\textrm{s2}}(t)$ is time-dependent.
-The relation between the amount of water and ice is
-$m_{\textrm{w}}(t)=m_{\textrm{w}}(t_{\textrm{f3}})-m_{\textrm{i}}.$ (25)
+During the solidification step, the amount of water $m_{\textrm{w}}(t)$ and ice $m_{\textrm{i}}(t)$ varies with time, which also affects the volume and surface area of the product.
+The relation between $m_{\textrm{w}}$ and $m_{\textrm{i}}$ is
+$m_{\textrm{w}}=m_{\textrm{w}}(t_{\textrm{f3}})-m_{\textrm{i}}.$ (26)
+The temperature of the liquid phase and the solid-liquid interface follows the freezing-point depression relation
+$T=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}}\right).$ (27)
+The ice layer becomes an additional heat transfer resistance, and so the heat transfer coefficients $h_{\textrm{s2}}$ and $h_{\textrm{s3}}$ in (8) and (6) can be replaced by the overall heat transfer coefficients
+$U_{\textrm{s2}}=\frac{1}{1/h_{\textrm{s2}}+l/k_{\textrm{i}}},$ (28)
+$U_{\textrm{s3}}=\frac{1}{1/h_{\textrm{s3}}+\frac{r_{\textrm{o}}\ln(r_{\textrm{o}}/r)}{k_{\textrm{i}}}}.$ (29)
+where $r_{\textrm{o}}$, $r$ and $l$ are as defined in Fig.
+4 and $k_{\textrm{i}}$ is the thermal conductivity of ice.
+For the radiation part, the linearization technique described in Section 3.2.1 can be applied, and so the equation is written in the form of Newton’s law of cooling.
+Note that the overall heat transfer coefficients in (28) and (29) are defined based on the outer surface area to be consistent with the original expressions of $Q_{\textrm{s2}}$ and $Q_{\textrm{s3}}$
 The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely (17), (19), and (21).
@@
 where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [12].
-3.2.5 Cooling
+Our hybrid lumped capacity model presented in this section has the same model complexity and computational cost as those of the original lumped capacity model (e.g., in Ref [12]) while describing the physics in a more accurate way as in the 2D model (e.g., in Ref [13]).
+Final simulations results from these models are almost identical for lyophilization of unit doses, and so there is no concern in terms of model accuracy.
+Cooling
 The final step is to ensure that the product temperature is at the desire value before starting the drying process.
@@
 with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively.
-Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, taking the final conditions of the solidification step.
+Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, following the final conditions of the solidification step.
 The initial condition for (31) is the final temperature of the solidification step.
 It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (31) without any significant error.
-By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature and amount of substances (e.g., water and ice) during the freezing
+By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature, phase transition, and amount of ice/water during the freezing step can be predicted.
 Model for primary drying
-In conventional lyophilization where vials are placed on the heating shelf, there are three heat transfer mechanisms associated with the bottom surface, namely conduction, convection, and radiation, whereas only radiation dominates at the side and top surfaces [17].
+In conventional lyophilization, vials are placed on the heating shelf, and thus heat transfer at the bottom surface is driven by conduction (at the point of contact), convection (gas), and radiation, whereas only radiation dominates at the side and top surfaces [18].
 For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf.
@@
 Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$.
-If the primary drying model is simulated following the freezing step, then set $t_{0}=t_{\textrm{f4}}$.
+If the primary drying model is simulated consecutively after the freezing step, then set $t_{0}=t_{\textrm{f4}}$.
 Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
-The governing equations for the primary drying step consists of (1) the energy balance in the frozen region and (2) the mass balance at the sublimation front.
-By assuming that the supplied heat is used in the frozen region only, the energy balance for the frozen region is
-$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=k_{\textrm{f}}\dfrac{\partial^{2}T}{\partial z^{2}}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad S(t)<z<H,$ (28)
+The governing equations for the primary drying step consist of the (1) energy balance in the frozen region and (2) mass balance at the sublimation front.
+By assuming that the supplied heat is used in the frozen region only, the energy balance for the frozen region can be described by the partial differential equation (PDE)
+$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=k_{\textrm{f}}\dfrac{\partial^{2}T}{\partial z^{2}}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad S<z<H,$ (32)
 where $T(z,t)$ is the temperature, $S(t)$ is the sublimation front/interface position, $k$ is the thermal conductivity, $\rho$ is the density, $C_{p}$ is the heat capacity, $V$ is the volume, $H$ is the height of the product, and the subscript ‘f’ denotes the frozen region.
-We refer to Appendix A for the calculations of some relevant parameters.
+We refer to Appendix B for the calculations of some relevant parameters.
 The radiative heat transfer from the sidewall $Q_{\textrm{rad}}$ is
-$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\!\left(T_{\textrm{c}}^{4}-T^{4}\right),$ (29)
-where $\mathcal{F}$ is the transfer factor, $\sigma$ is the Stefan-Boltzmann constant, $T_{\textrm{c}}$ is the chamber wall temperature, and $A_{r}=\pi dH$ is the side area of the product.
+$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}_{\textrm{s3}}\!\left(T_{\textrm{c}}^{4}-T^{4}\right),$ (33)
+where $T_{\textrm{c}}$ is the chamber wall temperature and $A_{r}=\pi dH$ is the side area of the product.
 We refer to Section 3.5 for the detailed derivation of (33).
@@
 where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance.
-The saturation pressure for sublimation can be is described by [22]
+The saturation pressure for sublimation is described by [22]
 $p_{\textrm{w,sat}}=\exp\!\left(\frac{-6139.9}{T}+28.8912\right).$ (36)
-The variation of the resistance normally follows the empirical expression [17, 21, 22]
+The variation of the mass transfer resistance can be approximated by the empirical expression [18, 4, 22]
 $R_{\textrm{p}}=R_{\textrm{p0}}+\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S},$ (37)
-where $R_{\textrm{p}}$ is the mass transfer resistance (aka cake resistance) and $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, $R_{\textrm{p2}}$ are the constants estimated from data.
-The partial differential equation (PDE) (28) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
-$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H,$ (34)
-where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient which combines the effects of thermal radiation and natural convection.
-At the top surface, the energy balance associated with sublimation and thermal radiation from the above shelf/wall is
-$N_{\textrm{w}}\Delta H_{\textrm{sub}}=k_{\textrm{f}}\dfrac{\partial T}{\partial z}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=S(t),$ (35)
-where $\Delta H_{\textrm{sub}}$ is the heat of sublimation and $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial.
+where $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, and $R_{\textrm{p2}}$ are the constants to be estimated from data.
+Physically, $R_{\textrm{p0}}$ represents the resistance associated with mass convection above the product, whereas $\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S}$ corresponds to the resistance associated with diffusion through the porous dried layer in the product.
+The PDE (32) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
+$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (38)
+where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient that combines the effects of thermal radiation and natural convection.
+At the top surface, the energy balance associated with sublimation and thermal radiation from the upper surface of the chamber is
+$N_{\textrm{w}}\Delta H_{\textrm{sub}}=k_{\textrm{f}}\dfrac{\partial T}{\partial z}+\sigma\mathcal{F}_{\textrm{s1}}\big{(}T_{\textrm{u}}^{4}-T^{4}\big{)},\qquad z=S,$ (39)
+where $\Delta H_{\textrm{sub}}$ is the heat of sublimation.
 The initial conditions for (32) and (34) are
@@
 $S(t_{0})$ $=0.$ (41)
-If the model is simulated following the freezing step, then set $T_{0}=T(t_{\textrm{f4}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone primary drying simulation.
-The primary drying model is simulated until the interface position is equal to the height of the product $S(t)=H$, indicating that there is no frozen material left, which marks the end of the primary drying step $t_{\textrm{d1}}$.
-In some cases, there might be an additional heating period at the end of primary drying to adjust the temperature and ensure complete sublimation before starting the secondary drying stage.
-However, this period is trivial to model; only a simple heat equation is required, and so is not described it here.
+For consecutive simulation with the freezing step, set $T_{0}=T(t_{\textrm{f4}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone primary drying simulation.
+The primary drying model is simulated until the interface position is equal to the height of the product, i.e., $S=H$, indicating that there is no frozen material left, which marks the end of the primary drying step at $t_{\textrm{d1}}$.
+In some cases, there might be an additional heating period at the end of primary drying to adjust the temperature and ensure complete sublimation before starting the secondary drying step.
+However, the model contains only a simple heat equation, and so it is not detailed here.
 Model for secondary drying
-The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) as done for the primary drying model (Fig.
-3C).
+The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig.
+3C), which is consistent with the primary drying model.
 The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$.
-If the secondary drying model is simulated following the primary drying step, then set $t_{0}=t_{\textrm{d2}}$.
+If the secondary drying model is simulated consecutively after the primary drying step, then set $t_{0}=t_{\textrm{d2}}$.
 Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
-The governing equations for the secondary drying step consists of (1) the energy balance in the dried region and (2) the desorption kinetics. The energy balance of the dried product is
+The governing equations for the secondary drying step comprise the (1) energy balance in the dried region and (2) desorption kinetics. The energy balance of the dried product is
 $\rho_{\textrm{e}}C_{p,\textrm{e}}\dfrac{\partial T}{\partial t}=k_{\textrm{e}}\dfrac{\partial^{2}T}{\partial z^{2}}+\rho_{\textrm{d}}\Delta H_{\textrm{des}}\dfrac{\partial c_{\textrm{w}}}{\partial t}+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad 0\leq z\leq H,$ (42)
-where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the bound water concentration (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the enthalpy of desorption, $Q_{\textrm{rad}}$ is as defined in (29), and the subscript ‘e’ denotes the effective properties considering both solid and gas in the pores.
+where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the concentration of bound water (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the heat of desorption, $Q_{\textrm{rad}}$ is as defined in (33), and the other parameters are as defined in (32), with the subscript ‘e’ denoting the effective properties considering both solid and gas in the pores.
 The desorption kinetics of bound water is described by
 $\dfrac{\partial c_{\textrm{w}}}{\partial t}=k_{\textrm{d}}(c^{*}_{\textrm{w}}-c_{\textrm{w}}),$ (43)
-where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{w}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [14, 15, 21]
+where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{d}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [15, 16, 4]
 $k_{\textrm{d}}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT},$ (44)
-where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency) and $E_{\textrm{a}}$ is the activation energy.
+where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency), $E_{\textrm{a}}$ is the activation energy, and $R$ is the gas constant.
 The above desorption kinetics (43) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [15, 43, 19, 23, 4].
-Further simplification that is very common in the literature can be done by setting $c^{*}_{\textrm{w}}=0$.
+Further simplification that is relatively common can be done by setting $c^{*}_{\textrm{w}}=0$.
 This simplification produces insignificant error as shown in Ref.
@@
 The governing PDE (42) requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
-$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H.$ (41)
+$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (45)
+where the value of $h_{\textrm{b}}$ is approximated to have the same value as that used in primary drying (38).
 Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
-$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=0.$ (42)
+$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\mathcal{F}_{\textrm{s1}}\big{(}T_{\textrm{u}}^{4}-T^{4}\big{)},\qquad z=0.$ (46)
 The initial conditions for (42) and (43) are
@@
 $c_{\textrm{w}}(z,t_{0})=c_{\textrm{w,0}},\qquad 0\leq z\leq H.$ (48)
-If the model is simulated following the primary drying step, then set $T_{0}=T(z,t_{\textrm{d1}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone secondary drying simulation.
-The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step $t_{\textrm{d2}}$.
+For consecutive simulation with the primary drying step, set $T_{0}=T(z,t_{\textrm{d1}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone secondary drying simulation.
+The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step at $t_{\textrm{d2}}$.
 Modeling heat transfer in suspended vials
-Convection and radiation are two key heat transfer mechanisms in the suspended-vial configuration as discussed before.
+Convection and radiation are the two important modes of heat transfer in suspended vials as discussed extensively in the previous sections.
 During the freezing step, there could be a combination of natural/forced convection and radiation, depending on how the cryogenic gas is fed to the system.
@@
 Convection can be simply modeled using Newton’s law of cooling.
-A heat transfer coefficient for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection).
+Heat transfer coefficients for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection).
 Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref.
@@
 [31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization.
-Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; please refer to the aforementioned references for more details.
+Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; more detailed discussion can be found in the aforementioned references.
 Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively.
 In general, the net radiant energy receiving by surface 1 can be written as
-$Q_{\textrm{rad}}=\sigma A_{1}\mathcal{F}(T_{2}^{4}-T_{1}^{4}),$ (45)
-where $A$ is the surface area, $\varepsilon$ is the emissivity, $\mathcal{F}$ is the transfer factor, and the subscripts ‘1’ and ‘2’ denote surfaces 1 and 2, respectively.
-The transfer factor is dependent on geometry and material type (emissivity).
+$Q_{\textrm{rad}}=\sigma A_{1}\mathcal{F}\big{(}T_{2}^{4}-T_{1}^{4}\big{)},$ (49)
+where $A$ is the surface area, $\mathcal{F}$ is the transfer factor, and the subscripts ‘1’ and ‘2’ denote surfaces 1 and 2, respectively.
+The transfer factor is dependent on the geometry and material properties (emissivity) of both surfaces.
 For any two diffuse, gray surfaces that form an enclosure,
-$Q_{\textrm{rad}}=\dfrac{\sigma(T_{2}^{4}-T_{1}^{4})}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (46)
-where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 ($F_{1-2}=1$) and surface 2 is much larger than surface 1, (46) can be simplified as
-$Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}(T^{4}_{2}-T^{4}_{1}).$ (47)
-In (47), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Now consider the case when surface 1 is the vial and surface 2 is the chamber walls. In this case, (47) becomes
-$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}(T^{4}_{\textrm{c}}-T^{4})$ (48)
-for the side surface and
-$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{z}(T^{4}_{\textrm{c}}-T^{4})$ (49)
-for the top and bottom surfaces.
+$Q_{\textrm{rad}}=\dfrac{\sigma\big{(}T_{2}^{4}-T_{1}^{4}\big{)}}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (50)
+where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 (i.e., $F_{1-2}=1$) and surface 2 is much larger than surface 1, (50) can be simplified as
+$Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}\big{(}T^{4}_{2}-T^{4}_{1}\big{)}.$ (51)
+In (51), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Next, consider the case where a single vial (surface 1) is surrounded by the chamber walls (surface 2). In this case, (51) becomes
+$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}\big{(}T^{4}_{\textrm{c}}-T^{4}\big{)},$ (52)
+where $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial.
 Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$.
-The amount of radiative heat for the single-vial case represented by (48) and (49) defines the upper bound on the radiative heat that the vial can receive from the chamber walls.
-This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among vials.
+The amount of radiative heat for the single-vial case represented by (52) defines the upper bound on the radiative heat that one vial can receive from the chamber walls.
+This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among the vials.
 As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$.
-Therefore, (48) needs to be modified to account for interactions between vials, whereas (49) can be used directly because thermal radiation at the top and bottom surfaces is not influenced by the other vials.
+Therefore, (52) needs to be modified to account for interactions between vials.
 Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials.
@@
 In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls.
-A rigorous way of modeling this complicated radiation is known as the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref.
+A rigorous way of modeling this complicated radiation is the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref.
 [39].
@@
 Therefore, instead of using the radiation network approach, we modify (52) to follow (49) as
-$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}(T^{4}_{\textrm{c}}-T^{4}),$ (50)
+$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\big{(}T^{4}_{\textrm{c}}-T^{4}\big{)},$ (53)
 where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph.
 This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (50).
-This technique is similar to the hybrid approach presented in Ref.
-[39].
-Note that this approximation is for the suspended-vial configuration, in which heat transfer between vials is not significant; otherwise, the radiation network approach is recommended.
-When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; for the temperature of about 300K, the error is about 0.1% [31].
-The linearized equation can then be combined with the convection part, with the overall heat transfer coefficient taking into account of both convection and radiation, e.g., (34).
+In our model, the transfer factor $\mathcal{F}_{\textrm{s3}}$ appears in (6), (32), and (42).
+If the chamber design and vial configuration are identical for both freezing and drying steps, the same value of $\mathcal{F}_{\textrm{s3}}$ can be used for all equations.
+A similar analysis can also be done for $\mathcal{F}_{\textrm{s1}}$ in (39) and (46).
+When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; at the temperature of about 300 K, the error caused by linearization is about 0.1% for the temperature difference of 20 K and 2% for the temperature difference of 100 K (see Section 1.3.2 of Ref.
+[31]), which is tiny.
+The linearized equation can then be combined with the convection part, with the corresponding heat transfer coefficient taking into account of both convection and radiation, e.g., (8), (38), and (45).
+This approximation helps simplify the equations and also facilitates the use of an overall heat transfer coefficient (e.g., (29)) as both convection and radiation parts are written in the form of Newton’s law of cooling.
 […]
 Default model parameters
-The proposed model developed in Sections 3.2, 3.3, and 3.4 require a large number of input parameters.
-This section defines the default parameter values for our model to be used in the subsequent sections in this work.
+This section defines the default parameter values for the model developed in Sections 3.2, 3.3, and 3.4.
 These parameter values are either obtained from the literature or set based on the typical values used in lyophilization.
 This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively.
-Table LABEL:Tab:DefaultParameters shows the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
+Table LABEL:Tab:DefaultParameters lists the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
 Table 1: Default model parameters for a complete continuous lyophilization simulation.
@@
 $\rho_{\textrm{d}}$ 212.21 kg/m3 [16]
-$\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix A)
+$\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix B)
 $\rho_{\textrm{i}}$ 917 kg/m3 [32]
-$\rho_{\textrm{s}}$ 1,587.9 kg/m3 [32]
-$\rho_{\textrm{w}}$ 1,000 kg/m3 [12]
+$\rho_{\textrm{s}}$ 1587.9 kg/m3 [32]
+$\rho_{\textrm{w}}$ 1000 kg/m3 [13]
 Operating conditions
@@
 $c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [1]
-$p_{\textrm{t}}$ see Fig. 5D Pa
+$p_{\textrm{t}}$ $10^{5}$ (freezing) Pa
+$10^{4}$ (VISF) Pa
 $p_{\textrm{w,c}}$ 3 Pa [7]
@@
 295 (secondary drying) K
-$T_{\textrm{c}}$ 273 (before VISF) K
+$T_{\textrm{c}},T_{\textrm{u}}$ 273 (before VISF) K
 240 (after VISF) K
@@
 Heat and mass transfer
-$\mathcal{F}$ 0.78
+$\mathcal{F}_{\textrm{s1}}$ 0.8
+$\mathcal{F}_{\textrm{s3}}$ 0.624
 $f_{\textrm{a}}$ 1.5$\times$10-3 1/s
@@
 […]
-$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
-$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [6]
-$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [6]
+$R_{\textrm{p0}}$ 1.5$\times$104 m/s
+$R_{\textrm{p1}}$ 3.0$\times$107 1/s
+$R_{\textrm{p2}}$ 10 1/m
 […]
 Product and formulation
-$H$ 7.2$\times$10-3 m calculated (see Appendix A)
-$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix A)
-$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix A)
+$H$ 7.2$\times$10-3 m calculated (see Appendix B)
+$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix B)
+$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix B)
 $V_{\textrm{l}}$ 3$\times$10-6 m3 [7]
@@
 […]
-$M_{\textrm{in}}$ 28$\times$10-3 kg/mol [40]
+$M_{\textrm{in}}$ 0.028 kg/mol [44]
 $M_{\textrm{s}}$ 0.3423 kg/mol [13]
-$M_{\textrm{w}}$ 18$\times$10-3 kg/mol [40]
+$M_{\textrm{w}}$ 0.018 kg/mol [44]
 […]
-Parameters that should be estimated from experimental data include heat transfer coefficients, mass transfer coefficients and cake resistance, and desorption-related parameters (frequency factor and activation energy).
+The parameters that should be estimated from experimental data include heat transfer coefficients, mass transfer coefficients and cake resistance, and desorption-related parameters (frequency factor and activation energy).
+Operating conditions that are not measured, e.g., wall temperatures, could also be estimated from data.
 Implementation
-In this article, the mechanistic model was implemented in MATLAB R2023a, and so all results and plots were generated using MATLAB.
-A system of ordinary differential equations (ODEs) was integrated with ode15s.
-Partial differential equations (PDEs) were discretized spatially to produce a system of ODEs, which was then fed to the ODE solver (see Appendix B for the detailed numerical methods).
+The mechanistic model for continuous lyophilization developed in this article is aimed to be used for various purposes, ranging process design and development to the actual operation (e.g., real-time control, state estimation, and optimization).
+In continuous operation, the process runs continuously with minimal human intervention, and thus efficient (ideally real-time) simulation and computation are important.
+Therefore, appropriate numerical methods must be applied.
+The detailed numerical methods and implementation are given in Appendix C.
+All simulations, calculations, and results presented in this article were performed and generated using MATLAB R2023a.
+The model, data, and MATLAB code used in this work are made available (see the Data Availability section).
+In addition to the MATLAB implementation, we also provide an equivalent software package in Julia, a high-level, open-source programming language for scientific computing, for the selected simulations.
+Results and Discussion
 Model validation
@@
 [7].
-The model is validated for each lyophilization step separately, with the parameters reported in Table 2.
-The default values defined in Table LABEL:Tab:DefaultParameters are used for parameters not shown in Table 2.
+Our mdel validation is carried out for each lyophilization step separately, with the specific parameters reported in Table 2.
 Table 2: Parameters used for the model validation, including Cases 1, 2a, 2b, 3a, and 3b.
@@
 $T_{0}$ 280 K [7]
-$T_{\textrm{c}}$ 272 K estimated from data in Ref. [6]
+$T_{\textrm{c}},T_{\textrm{u}}$ 272 K estimated from data in Ref. [7]
 $T_{\textrm{g}}$ see Fig. 5A K [7]
@@
 313 (Case 2b) K [7]
-$T_{\textrm{b}}$ 275 K estimated from data in Ref. [6]
+$T_{\textrm{c}},T_{\textrm{u}}$ 275 K estimated from data in Ref. [7]
 […]
-$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
 $R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [7]
@@
 $T_{\textrm{b}}$ 293 K [7]
-$T_{\textrm{c}}$ 285 K estimated from data in Ref. [6]
+$T_{\textrm{c}},T_{\textrm{u}}$ 285 K estimated from data in Ref. [7]
 $f_{\textrm{a}}$ 0.42 1/s estimated from data in Ref. [7]
@@
 The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively.
-Panel C shows the model prediction and experimental data for the bound water concentration during the secondary drying step.
-The initial concentrations for Cases 3a and 3b are 0.0864 and 0.0739 kg water/kg solid, respectively.
+Panel C shows the model prediction and experimental data for the average concentration of bound water during the secondary drying step.
+The initial concentrations for Cases 3a and 3b are 0.088 and 0.075 kg water/kg solid, respectively.
 For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig.
-4A).
-During the preconditioning phase, the temperature decreases from its initial value of about 280 K to the nucleation temperature of about 263 K, where nucleation starts.
-At the first nucleation, the temperature instantaneously increases from 263 K to about the equilibrium freezing point of about 272 K.
-The temperature slowly decreases during the solidification phase, which takes about 30 min, following the freezing-point depression.
-Subsequently, the product temperature decreases and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
-The model can accurately predict the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig.
-4B).
-The product temperature increases fast at the beginning because the driving force for sublimation is low, and so most of the heat input is used to increase the temperature.
-As time progresses, the driving force for sublimation is higher, resulting in most of the heat used for sublimation instead of increasing the product temperature.
-The maximum error in temperature prediction is about 2–3 K, while that of drying time prediction is about 0.2 for Case 2a and 0.4 for Case 2b ($\sim$ 6%).
-This error could be partly because of the uncertainty of cake resistance, which is caused by a stochastic freezing process.
-In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.0864 and 0.0739 kg water/kg solid).
-The concentration decreases exponentially following the linear driving force model, where the simulations are terminated when the concentration reaches 0.01 kg water/kg solid.
+5A).
+The product temperature decreases from its initial value of 280 K to the nucleation temperature of about 263 K, where nucleation starts.
+Upon the first nucleation, the temperature instantaneously rises up from 263 K to the freezing point of about 272 K.
+The temperature slowly decreases during the solidification phase for about 30 min, following the freezing-point depression.
+Subsequently, the temperature decreases faster and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
+The model accurately predicts the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig.
+5B).
+The temperature rises up rapidly at the beginning and gradually increases at a slower rate toward the end of the process.
+This is because the driving force for sublimation is initially low, and so most of the heat input is used to increase the temperature.
+As time progresses, the driving force for sublimation is larger; thus, most of the heat input is used for sublimation.
+The maximum error in temperature prediction is about 2–3 K, while the error in drying time prediction is about 0.2 h for Case 2a and 0.4 h for Case 2b ($\sim$6%).
+This error could be partly because of the uncertainty in cake resistance, which results from the stochastic nature of the freezing process.
+In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.088 and 0.075 kg water/kg solid).
+The concentration decreases exponentially from its initial value to the final concentration of 0.01 kg water/kg solid, the threshold set for terminating the simulations.
+This concentration profile is expected given the linear driving force model is used to describe the desorption process.
 It is important to note that the data used in Fig.
-4C are based on the shelf temperature of 293 K.
-To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures such that the effects of temperature on the desorption process can be quantified.
-Case Studies
-The previous sections detail the development and validation of our mechanistic model for lyophilization. This section demonstrates various applications of our model via four different case studies.
-Process design and optimization
-The first case study focuses on model-based design and optimization of the entire continuous lyophilization process.
-One of the critical aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production.
-In the context of continuous lyophilization, our model can be used to study the evolution of critical process parameters, e.g., temperature, moisture content, throughout the process at various operating conditions, which can help guide the process design and optimization.
-This work not only presents the first mechanistic model for continuous lyophilization, but also is one of the very few studies that proposes a complete model that includes all three steps in lyophilization, including freezing, primary drying, and secondary drying.
+5C are based on the shelf temperature of 293 K.
+To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures so that the effects of temperature on the desorption process can be quantified.
+Prediction of critical process parameters for a complete continuous lyophilization cycle
+One of the important aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production.
+In the context of continuous lyophilization, our model can be used to study the evolution of critical process parameters – e.g., temperature, moisture content – throughout the process at various operating conditions, which can help guide process design and optimization.
+This work not only presents the first mechanistic model for continuous lyophilization of suspended vials, but also is one of the very few studies that develops a complete model incorporating all three steps of lyophilization, including freezing, primary drying, and secondary drying.
 This is important specifically for continuous manufacturing where the entire process should be designed and optimized simultaneously.
-By using the default parameters in Table LABEL:Tab:DefaultParameters, we simulate a complete continuous lyophilization cycle including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig.
-5).
-The evolution of the product temperature (Fig.
-5A), mass of ice (Fig.
-5B), and bound water concentration (Fig.
-5C) can be obtained, given the operating pressure (Fig.
-5D) and temperature (Fig.
-5E).
+With the default model parameters in Table LABEL:Tab:DefaultParameters, a complete continuous lyophilization cycle can be simulated, including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig.
+6).
+The time evolution of the product temperature (Fig.
+6A), mass of ice (Fig.
+6B), and concentration of bound water (Fig.
+6C) can be obtained, given the operating pressure (Fig.
+6D) and temperature (Fig.
+6E).
 This information could be useful for various purposes.
-For example, it can guide the process and equipment design concerning chamber sizing and vial moving speed to accommodate the residence time of each step.
+For example, it can guide process and equipment design concerning the chamber sizing and velocity profile of each vial, ensuring that the target residence time is met for each chamber.
 Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time.
 It is important to note that the entire simulation shown in Fig.
-5 can be computed in less than 1 second on a normal laptop, which makes the model very practical to be implemented for any purposes.
-Figure 5: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) concentration of bound water, (D) operating pressure, and (E) operating temperature.
-Studying vacuum-induced surface freezing
-Our model can be simulated for the whole lyophilization process as shown in Section 5.1 or simulated separately for each step in the lyophilization cycle to provide insights into some part of the process.
-This section employs the model to study the freezing step, which usually receives little attention compared to the drying steps.
-The proposed model can simulate different freezing protocols, namely uncontrolled (aka spontaneous) and controlled nucleation.
-Spontaneous nucleation was already investigate in detail by Refs.
-[11, 12], and so here we consider the controlled nucleation case via VISF, which is preferable for continuous lyophilization.
-As an example, the evolution of the product temperature and amount of water evaporating during VISF at different pressure profiles is simulated (Fig.
-6).
-Varying among different systems, this information could be used to design a freezing protocol, e.g., pressure profile, that achieves the desired nucleation temperature.
-Figure 6: Simulation results showing the evolution of the product during VISF at different pressure profiles.
-Panel A shows the pressure profiles that decrease linearly from $10^{5}$ Pa to $10^{4}$ Pa, 1,000, and 100 Pa, respectively.
-Panel B shows the evolution of the product temperature.
-Panel C shows the total mass of water, which implies the evaporation rate.
-External heat transfer with the wall and cryogenic gas is excluded so that only the contribution from VISF is considered.
-Model modifications for conventional drying
+6 can be computed in less than 1 s on a normal laptop, which makes the model very practical to be implemented for any purposes.
+Figure 6: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) average concentration of bound water, (D) operating pressure, and (E) operating temperature.
+Visualizing spatiotemporal data
+As discussed in Section 3.1, one of the key considerations during the drying steps is to ensure that the product temperature at any location does not exceed the upper limit.
+Consequently, spatiotemporal data are needed.
+Spatial variation of temperature and concentration in the product are dependent on several factors.
+For instance, temperature gradients are larger when the sample thickness and heat transfer coefficient increase (see the analysis in Appendix A).
+[…]
+Figure 7: Spatiotemporal evolution of the product temperature and sublimation front during primary drying. Only the frozen region is shown in the figure.
+Figure 8: Spatiotemporal evolution of the concentration of bound water during secondary drying.
+The initial concentration is assumed to vary linearly from 5% (0.05 kg water/solid) at the top to 20% (0.2 kg water/kg solid) at the bottom.
+Application to freezing: Optimizing freezing protocols
+This section employs the model to provide insights into the freezing step, which usually receives little attention compared to the drying steps.
+As spontaneous nucleation was already investigated in detail by Refs.
+[12, 13], controlled nucleation, which is preferable for continuous lyophilization, is considered instead.
+The evolution of the product temperature and amount of water during VISF at different pressures predicted by our model is shown in Fig.
+9.
+By varying the pressure profile, it is possible to control the rate of change in the product temperature, hence nucleation.
+The model can also be used to estimate the amount of water that evaporates during VISF as well as the rate of evaporation.
+This information could be useful for the development of a freezing protocol for controlled nucleation.
+Figure 9: Simulation results showing the (A) evolution of the product during temperature and (B) amount of water during VISF at three different pressures, namely $10^{4}$, 5000, and 1000 Pa.
+At $t=0$, the pressure decreases linearly from $10^{5}$ Pa to the target pressure in 30 s before being held constant.
+The nucleation temperature is set to 263 K.
+External heat transfer with the wall and cryogenic gas is omitted from the model so that only the contribution from VISF is observed.
+Other model parameters are kept at the default values.
+Application to drying: Modifying the model for conventional lyophilization
 Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization.
@@
 In this case study, we consider batch primary drying and secondary drying using the data from Refs.
-[18] and [21], with the parameters reported in Table 3.
+[19] and [4], with the specific parameters reported in Table 3.
 Table 3: Parameters used for the conventional drying studies, including for Cases 4a, 4b, and 5.
@@
 $T_{\textrm{b}}$ see1 K [19]
-$\mathcal{F}$ 0 turn off the radiation part
+$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
 […]
@@
 $c_{\textrm{w,0}}$ 0.0603 kg water/kg solid [4]
-$T_{0}$ 273 K [21]
+$T_{0}$ 264 K [4]
 $T_{\textrm{b}}$ see2 K estimated from data in Ref. [4]
@@
 […]
-$\mathcal{F}$ 0 turn off the radiation part
+$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
 […]
@@
   • 1
-    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 (Case 4b).
+    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 K (Case 4b).
   • 2
-    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264.09 K to 312 K for Case 5.
+    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264 K to 312 K for Case 5.
 Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig.
-7).
-The model can also effectively predict the product temperature and bound water concentration during the secondary drying step (Fig.
-8).
+10).
+The model can also accurately predict the product temperature and concentration of bound water during the secondary drying step (Fig.
+11).
 With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
-Figure 7: Comparison between our model prediction and the experimental/simulation data from Ref.
+Figure 10: Comparison of our model prediction and the experimental/simulation data from Ref.
 [19] for the primary drying step in conventional lyophilization.
-Panel A shows the model prediction and the experimental data for the product temperature at the bottom surface.
+Panel A shows the model prediction and experimental data for the product temperature at the bottom surface.
 Panel B shows the model prediction and reference simulation data for the sublimation front position.
-The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15, while other conditions are identical.
-Figure 8: Comparison between our model prediction and the experimental data from Ref.
+The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15 K, while other conditions are identical.
+Figure 11: Comparison of our model prediction and the experimental data from Ref.
 [4] for the secondary drying step in conventional lyophilization.
 Panel A shows the model prediction and experimental data for the product temperature at the bottom surface.
-Panel B shows the model prediction and experimental data for the concentration of bound water.
+Panel B shows the model prediction and experimental data for the average concentration of bound water.
 Condenser failure analysis
-Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily at its desired conditions.
+Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily under the desired conditions.
 A better process model should be able to simulate important abnormal conditions, e.g., equipment failure, that could occur due to various unexpected scenarios, which is critical in continuous manufacturing as these abnormal operations could affect the reliability, availability, and maintainability (RAM) as well as the safety of the plant/process.
@@
 During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1].
-This is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase.
-This scenario may not lead to safety-related issues because the operating pressure in lyophilization is usually low and the amount of ice/water is not high enough to create overpressure.
-Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (31)), which then prolongs the primary drying stage.
+This operational constraint is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase.
+This scenario may not lead to safety-related issues because the operating pressure in lyophilization is low and the amount of ice/water is generally not high enough to create overpressure.
+Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (35)), which then prolongs the primary drying step.
 As a result, the final product quality could be affected.
@@
 Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption.
-To incorporate the condenser and dynamics of vapor flow into the model, consider the mass balance equation for water vapor in a chamber [22]
+To incorporate the condenser dynamics into the model, we consider cases where the total condenser capacity is not enough to condense the water vapor produced via sublimation during primary drying.
+In such cases, the mass balance for the water vapor in a chamber is [22]
 $\frac{dp_{\textrm{w,c}}}{dt}=\frac{(j_{\textrm{w}}-j_{\textrm{w,max}})R\overline{T}}{V_{\textrm{c}}M_{\textrm{w}}},$ (54)
-where $j_{\textrm{w}}$ is the total mass flow rate of water vapor, $j_{\textrm{w,max}}$ is the maximum condenser capacity, $V_{\textrm{c}}$ is the volume of a chamber, and $T_{\textrm{avg}}$ is the average temperature in the chamber assumed to be constant for simplification.
+where $j_{\textrm{w}}$ is the total mass flow rate of water vapor resulting from sublimation, $j_{\textrm{w,max}}$ is the maximum condenser capacity, $V_{\textrm{c}}$ is the chamber volume, and $T_{\textrm{avg}}$ is the average temperature in the chamber assumed to be constant for simplification.
 The mass flow rate of water vapor can be calculated by
-$j_{m,\textrm{w}}=n_{\textrm{vial}}A_{z}N_{w},$ (52)
+$j_{\textrm{w}}=n_{\textrm{vial}}A_{z}N_{w},$ (55)
 where $n_{\textrm{vial}}$ is the number of vials in the chamber.
 By coupling (54) and (55) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
-Here we consider a situation where there is some issue with the condenser, and so the total condenser capacity cannot handle the water vapor produced via sublimation during primary drying, with the parameters given in Table 4.
-The simulation result shows that there is a pressure increase from the normal operating value of 3 Pa to 20 Pa (Fig.
-9A).
-This pressure increase results from the total vapor flow that exceeds the maximum condenser capacity, leading vapor accumulation in the chamber.
-When the pressure increases, the driving force for sublimation (also the sublimation flux) decreases.
-Finally, when the sublimation flux balances with the condenser capacity, the system pressure becomes constant.
-It is also shown that the drying time increases by about an hour (Fig.
-9B), and the product temperature is higher (Fig.
-9C).
+With the specific parameters given in Table 4, the total pressure increases from the normal operating value of 3 Pa to 20 Pa (Fig.
+12A).
+This pressure increase results from the total vapor flow exceeding the maximum condenser capacity, thus vapor accumulation in the chamber.
+When the system pressure increases, the driving force for sublimation (also the sublimation flux) decreases.
+After about 1 h, the system pressure becomes constant, indicating the equalization of the sublimation flux and condenser capacity.
+In this abnormal operation, the drying time and product temperature are higher than for the normal operation (Figs.
+12BC).
 Table 4: Parameters used for the condenser failure analysis.
 Symbol Value Unit Source
-$T_{0}$ 228 K
-$T_{\textrm{b}}$ 270 K
 $\overline{T}$ 260 K assumed to be constant
@@
 $V_{\textrm{c}}$ 0.118 m3 [22]
-$j_{m,\textrm{max}}$ 1.8$\times$10-5 m3 [22]
+$j_{\textrm{w,max}}$ 1.8$\times$10-5 m3/s [22]
 Figure 12: Simulation results comparing the normal operation and condenser failure scenario for the (A) system pressure, (B) sublimation front position, and (C) product average temperature during primary drying.
 Conclusion
-This work proposes the first mechanistic model for the suspended-vial continuous lyophilization technology, in which vials containing the pharmaceutical products are suspended and moved continuously along the process.
-The model can be used to simulate the entire lyophilization process, including the freezing, primary drying, and secondary drying steps.
-The freezing model considers preconditioning, vacuum-induced surface freezing (VISF), spontaneous and controlled nucleation, solidification, and final cooling.
-The primary drying model captures heat transfer in the frozen region and ice sublimation driven by mass transfer at the sublimation front.
-The secondary drying model describes simultaneous heat transfer in the dried region and bound water desorption.
+This article proposes the first mechanistic model for continuous lyophilization, in which the state-of-the-art suspended-vial technology is considered.
+The model can simulate the entire lyophilization process, including the freezing, primary drying, and secondary drying steps.
+The freezing model considers preconditioning, vacuum-induced surface freezing (VISF), spontaneous and controlled nucleation, solidification, and cooling.
+The primary drying model captures heat transfer in the frozen region and mass transfer resulting from sublimation of ice crystals at the sublimation front.
+The secondary drying model describes simultaneous heat transfer in the dried region and desorption of bound water.
 The overall model can describe the evolution of product temperature, ice/water fraction, and residual moisture throughout the process.
-Our model is validated using experimental data from the literature, in which the model predictions are consistent with the experimental measurements in all cases.
-The validated model is applied for four case studies that entail continuous lyophilization process design and optimization, detailed investigation of VISF, model modifications for conventional drying, and condenser failure analysis.
-The framework and results presented in this work are promising and could be used to guide the design and development of future continuous lyophilization technology.
-With a well-developed mechanistic model, future work could consider utilizing the model for several crucial tasks, e.g., uncertainty quantification, state estimation, and control design, which will ultimately lead to the development of a high-quality digital twin for continuous lyophilization that can greatly facilitate the process design and development in this area.
+The proposed model is validated for all three steps including freezing, primary drying, and secondary drying, in which the model predictions are consistent with the experimental measurements in all cases.
+The validated model is demonstrated for a variety of applications related to process design and optimization.
+Every single simulation can be run within less than 1 s on a normal laptop, allowing the model to be employed for any purpose.
+The framework and results presented in this work are suitable for guiding the design and development of future continuous lyophilization technology.
+With a well-developed mechanistic model, future work could consider utilizing the model for several important tasks, e.g., state estimation, optimal control design, and uncertainty quantification, which will ultimately support the development of a high-quality digital twin for continuous lyophilization to advance the manufacturing process.
 Data Availability
-All software and data will be made available upon publication of the manuscript.
-Acknowledgements
-This research was supported by the U.S. Food and Drug Administration under the FDA BAA-22-00123 program, Award Number 75F40122C00200.
+All software and data will be made available upon publication of the manuscript. Both MATLAB and Julia implementations will be provided.
+Acknowledgments
+RDB and PS were supported by the U.S.
+Food and Drug Administration under the FDA BAA-22-00123 program, Award Number 75F40122C00200.
+The authors would like to thank Pitipat Wongsittikan for helping optimize part of the Julia code.
 Appendices
-Appendix A Supplemental equations
-This appendix summarizes additional equations that are not crucial part of the model but required for calculating some model parameters.
+Appendix A Analysis of transport phenomena
+Biot number and spatial dimensions
+As discussed in the main text, the freezing step is modeled using the lumped capacity method, whereas the drying steps are analyzed in one spatial dimension, a vertical direction.
+Various studies have shown via simulation that variation in the radial direction is negligible [17, 13].
+This section provides a detailed analysis from the theoretical perspective to justify this assumption and briefly discusses strategies to modify the model for situations where this assumption is not true.
+Introduce the Biot number
+$\textrm{Bi}=\frac{hL}{k},$ (A.1)
+where $L$ is the characteristic length or length scale, $h$ is the heat transfer coefficient, and $k$ is the thermal conductivity.
+In general, the length scale is defined as the ratio of the volume to the surface area, which results in $r_{\textrm{o}}/2$ for a long cylinder.
+However, for conservative analysis, it is recommended to use the actual radius $r_{\textrm{o}}$ instead [45].
+The Biot number is defined as the ratio of the internal conduction resistance to the external convection resistance, and so a small Bi value implies that the resulting temperature gradient should also be small [31].
+In various heat transfer textbooks, the value of 0.1 is recommended as a guideline; if Bi $<0.1$, a lumped capacity model can be used.
+However, in practice, a lumped capacity model could still be valid even for Bi $>0.1$, depending on the system, which is analyzed here.
+[…]
+Figure A.1: Comparison between the 1D and lumped capacity models for the (A) volume-averaged temperature and (B) total heat removed for the default model parameters.
+[…]
+Figure A.2: Comparison between the 1D and lumped capacity model for the (A) volume-averaged temperature and (B) total heat removed for the conservative case.
+Finally, the errors (differences) between both models for different values of Bi are shown in Fig.
+A.3.
+Of course, if the Biot is too high, the lumped capacity model will not be accurate.
+That situation could happen, e.g., when a very large vial is used or the heat transfer coefficient is much higher than usual.
+In such cases, a high-fidelity model might be needed [13].
+However, the analysis in this section suggests that a lumped capacity model is highly accurate for typical lyophilization applications, in particular for unit doses.
+Figure A.3: Errors between the 1D and lumped capacity model for the (A) volume-averaged temperature and (B) total heat removed for different Bi.
+Mass transport in the dried product
+In secondary drying, the two important mechanisms for mass transport are (1) vapor phase or water vapor transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product.
+Previous simulation and experimental data suggest that desorption is the rate-limiting step [37, 23, 4, 24, 25, 36], and so the equations/terms related to vapor phase transport are typically omitted from recently published models, including our model.
+This section provides an analysis from the perspective of transport phenomena to justify this assumption.
+First, consider the vapor phase transport through the porous dried product.
+Modeling the transport of liquids and gases in porous media is challenging as it could entail a variety of mechanisms besides ordinary diffusion (e.g., Knudsen diffusion), and thus no fully satisfactory theory currently exists [46].
+Nevertheless, relative success has been achieved by using the concept of an effective diffusivity; in this analysis, we follow the procedure described in Ref.
+[31].
+Define the effective diffusivity $D_{\textrm{e}}$ by
+$\frac{1}{D_{\textrm{e}}}=\frac{1}{D_{\textrm{e,g}}}+\frac{1}{D_{\textrm{e,K}}},$ (A.2)
+where $D_{\textrm{e,g}}$ is the effective diffusion coefficient for ordinary diffusion, i.e., diffusion of the molecules in the vapor/gas phase, $D_{\textrm{e,K}}$ is the effective diffusion coefficient for Knudsen diffusion, and $D_{\textrm{e}}$ is the effective diffusivity that combines the effects of ordinary diffusion and Knudsen diffusion.
+For ordinary diffusion,
+$D_{\textrm{e,g}}=\frac{\epsilon}{\tau}D_{\textrm{g}},$ (A.3)
+where $D_{\textrm{g}}$ is the binary diffusion coefficient, e.g., water vapor in air, $\epsilon$ is the porosity, and $\tau$ is the tortuosity.
+For Knudsen diffusion, a similar expression can be used, that is,
+$D_{\textrm{e,K}}=\frac{\epsilon}{\tau}D_{\textrm{K}},$ (A.4)
+where $D_{\textrm{K}}$ is the Knudsen diffusion coefficient. This coefficient could be calculated from
+$D_{\textrm{K}}=97\bar{r}\sqrt{\frac{T}{M}},$ (A.5)
+where $\bar{r}$ is the average pore radius, $T$ is the temperature, and $M$ is the molar mass in g/mol.
+With the effective diffusivity calculated by (A.2)–(A.5), the total mass flux can be calculated using Fick’s law of diffusion, with the binary diffusion coefficient replaced by the effective diffusivity.
+Then, a continuity (mass balance) equation for the water vapor can be written and coupled to the mechanistic model proposed in the main paper.
+The key benefit of this technique is that the implementation closely resembles the approach for solving problems entailing ordinary diffusion of a binary mixture, which is relatively straightforward.
+Nevertheless, additional information is needed, including the porosity, tortuosity, and average pore radius.
+A more complicated strategy that is used in the literature is to apply the dusty gas model, which is derived from the Maxwell-Stefan Equation (see examples in [16, 19]).
+This approach, however, requires many more parameters and equations.
+Next, we compare the significance of vapor phase transport and desorption.
+Scale analysis can be used to analyze the contribution of each transport process (see various examples in [31]).
+Properly estimating the time scale of each phenomenon can help identify the key process that should be considered in the model and omit unimportant phenomena.
+The time scale (aka time constant) for mass diffusion can be approximated by [47, 31]
+$t_{\text{dff}}=\frac{L^{2}}{D_{\textrm{e}}},$ (A.6)
+where $t_{\text{dff}}$ is the time scale for diffusion and $L$ is the proper length scale. Similarly, the time scale for desorption can be approximated by
+$t_{\text{des}}=\frac{1}{k_{\textrm{d}}},$ (A.7)
+where $t_{\text{des}}$ is the time scale for desorption and $k_{\textrm{d}}$ is the desorption constant defined in (43).
+By comparing $t_{\text{dff}}$ and $t_{\text{des}}$, the limiting step for mass transfer in the secondary drying step can be identified, without any simulation or experimental data.
+For numerical comparison, a set of parameter values are obtained from Refs.
+[15, 17], namely $\bar{r}=5$$\times$$10^{-6}$ m, $\epsilon=0.815$, $\tau=1.2$, and $k_{\textrm{d}}=7.8$$\times$$10^{-5}$ s-1.
+The binary coefficient of water in air is obtained from Ref.
+[31], which is $D_{\textrm{g}}=1.97$$\times$$10^{-5}$ m2/s at 256 K and 1 atm.
+In secondary drying, the operating pressure is much lower than 1 atm, and so the actual $D_{\textrm{g}}$ is lower as the diffusion coefficient of gas mixtures at low pressure is inversely proportional to the pressure [31, 46], which makes our analysis more conservative.
+The length scale for this analysis is set to 0.01 m, a typical vial/sample size (thickness/diameter).
+Substituting these values into (A.6) and (A.7) gives $t_{\text{dff}}=7.5$ s and $t_{\text{des}}=3.6$ h.
+Obviously, the time scale for diffusion is much smaller than for desorption by several orders of magnitude, indicating that desorption is the rate-limiting step.
+A typical drying time of secondary drying is on the order of several hours, agreeing with the calculated $t_{\text{des}}$ of about 3.6 h.
+On the other hand, the time scale for diffusion is only 7.5 s, meaning that diffusion is much faster and so can be safely omitted from the model.
+Even some of the parameters are changed by an order of magnitude, this conclusion does not change.***This analysis holds for lyophilization of unit doses only; for lyophilization of bulk material (e.g., in food applications), the length scale of the product could be much larger, and so the conclusion might be different.
+Finally, there is another mass transfer mechanism that is not commonly discussed or modeled, namely diffusion of water in the solid matrix (not in the pores), i.e., solid-phase diffusion [37].
+Diffusion in the solid matrix is very slow, and so it could limit the mass transport in secondary drying.
+Modeling this phenomenon is challenging as it requires detailed knowledge of the solid matrix, particle size, and locations of bound water in the matrix, which is generally not available.
+A similar approximation to (A.6) can be carried out.
+By using the reported length scale of about $5$$\times$$10^{-7}$ m and diffusion coefficient of water in the solid phase of about $7$$\times$$10^{-16}$ m2/s [37], the time scale for solid-phase diffusion is about 10 min, which is still much lower than for desorption.
+To summarize, this section details the scale analysis that can be efficiently used to evaluate the effects of various mass transfer mechanisms.
+For lyophilization of pharmaceuticals in unit doses, our analysis shows that desorption of bound water is the rate-limiting step for mass transfer, and so other mechanisms can be omitted from the model, agreeing with both simulation and experimental data in the literature.
+Appendix B Supplemental equations
+This appendix summarizes equations that are not a crucial part of the model but required for calculating some model parameters.
 In Section 3.2.1, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
-$\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (A.1)
+$\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (B.1)
 Consequently, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated by
-$m_{\textrm{s}}=x_{\textrm{s}}\rho_{\textrm{l}}V_{\textrm{l}},$ (A.2)
-$m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (A.3)
-Then, (1) can be used as usual.
-During the freezing step, the surface area $A_{r}$ changes with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
-$A_{r}(t)=\frac{4\!\left(\dfrac{m_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{m_{\textrm{w}}(t)}{\rho_{\textrm{w}}}+\dfrac{m_{\textrm{i}}(t)}{\rho_{\textrm{i}}}\right)}{d},$ (A.4)
+$m_{\textrm{s}}=x_{\textrm{s}}\rho_{\textrm{l}}V_{\textrm{l}},$ (B.2)
+$m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (B.3)
+Then (1) can be used as usual.
+During the freezing step, the surface area $A_{r}$ varies with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
+$A_{r}(t)=\frac{4\!\left(\dfrac{m_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{m_{\textrm{w}}(t)}{\rho_{\textrm{w}}}+\dfrac{m_{\textrm{i}}(t)}{\rho_{\textrm{i}}}\right)}{d},$ (B.4)
 where the product is assumed to be a cylinder.
 For the primary drying model (Section 3.3), the volume of the frozen region is
-$V_{\textrm{f}}=\pi d^{2}(H-S).$ (A.5)
-The properties of the frozen region (subscript ‘f’) can be calculated from those of the ice and solid material or solute as follows:
-$\rho_{\textrm{f}}=\frac{1}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{i}}}},$ (A.6)
-$C_{p,\textrm{f}}=x_{\textrm{s}}C_{p,\textrm{s}}+(1-x_{\textrm{s}})C_{p,\textrm{i}},$ (A.7)
-$k_{\textrm{f}}=x_{\textrm{s}}k_{\textrm{s}}+(1-x_{\textrm{s}})k_{\textrm{i}}.$ (A.8)
-Here we assume that the final mass of ice is equal to the initial mass of water (complete freezing), meaning that the amount of water loss due to evaporation during VISF and unfrozen water (bound water) is negligible, which is generally true.
-Finally, the height/thickness of the product $H$ can be calculated by
-$H=\frac{(m_{\textrm{s}}+m_{\textrm{w,0}})}{\rho_{\textrm{f}}A_{z}}.$ (A.9)
-Note that, for some experiments that focus on primary drying, $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and $H$ are reported, and so the above calculations may not be needed.
+$V_{\textrm{f}}=\pi d^{2}(H-S).$ (B.5)
+The height/thickness of the product $H$ can be calculated by
+$H=\frac{m_{\textrm{s}}+m_{\textrm{w,0}}}{\rho_{\textrm{f}}A_{z}}.$ (B.6)
+The properties of the frozen region (subscript ‘f’) can be calculated from those of the ice and solid material or solute from
+$\rho_{\textrm{f}}=\frac{1}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{i}}}},$ (B.7)
+$C_{p,\textrm{f}}=x_{\textrm{s}}C_{p,\textrm{s}}+(1-x_{\textrm{s}})C_{p,\textrm{i}},$ (B.8)
+$k_{\textrm{f}}=x_{\textrm{s}}k_{\textrm{s}}+(1-x_{\textrm{s}})k_{\textrm{i}}.$ (B.9)
+These expressions assume that the final mass of ice is equal to the initial mass of water (complete freezing), meaning that the amount of water loss due to evaporation during VISF and unfrozen water (bound water) is negligible, which is generally true.
+Note that some experimental primary drying studies report $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and/or $H$, in which case the above calculations may not be needed.
 For the secondary drying model (Section 3.4), the volume of the dried region is
-$V_{\textrm{e}}=\pi d^{2}H,$ (A.10)
-which is constant because there is no moving boundary in this case.
-Appendix B Numerical methods
-This appendix comprehensively describes numerical techniques required for solving the model equations and summarizes the final equations implemented in MATLAB.
-For the freezing model, since a lumped thermal capacity approximation is used, the resulting equations are all ODEs, and so the equations can be solved directly with the selected ODE solver, which is ode15s in this case.
-For the primary drying model, the PDE (28) is discretized spatially to produce a system of ODEs, which is then fed to the ODE solver; the technique is referred to as the method of lines [41].
-Since the PDE (28) involves a moving interface $S(t)$, define a new variable
-$\xi=\frac{z-S}{H-S},$ (B.1)
-where $\xi$ represents is dimensionless position with respect to the moving interface (aka sublimation front).
+$V_{\textrm{e}}=\pi d^{2}H,$ (B.10)
+which is constant because there is no moving boundary (i.e., sublimation front) in this case.
+Appendix C Numerical methods
+This appendix comprehensively describes numerical techniques required for solving the model equations efficiently.
+With the given procedure, our model can be simulated accurately within 1 s on a normal laptop.
+For the freezing model, since the lumped capacity approximation is used, the resulting equations are ordinary differential equations (ODEs), and so the equations can be integrated directly with a proper ODE solver.
+This work uses MATLAB’s ode15s.
+For PDEs in the primary drying and secondary drying models, the method of lines [48] is recommended.
+The method of lines consists of two steps.
+First, the PDEs are discretized spatially to produce a system of ODEs.
+Second, the resulting ODEs are integrated with a proper ODE solver (MATLAB’s ode15s in this work).
+For the primary drying model, first consider the PDE (32). As the equation involves a moving interface $S(t)$, define a new variable
+$\xi=\frac{z-S}{H-S},$ (C.1)
+where $\xi$ is the dimensionless position with respect to the moving interface (aka sublimation front).
 This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$.
 Consequently, the PDE (32) and boundary conditions (38) and (39) become
-$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (B.2)
-$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (B.3)
-$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad\xi=0.$ (B.4)
+$\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (C.2)
+$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (C.3)
+$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),\qquad\xi=0.$ (C.4)
 To spatially discretize the PDE, define
-$\Delta\xi=\frac{1}{n_{z}-1},$ (B.5)
-$\xi=j\Delta\xi,$ (B.6)
+$\Delta\xi=\frac{1}{n_{z}-1},$ (C.5)
+$\xi=(j-1)\Delta\xi,$ (C.6)
 where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$.
-Discretizing (B.2)–(B.4) using the second-order finite difference scheme results in
-$\begin{split}\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(j\Delta\xi-1)}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (B.7)
-$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.8)
-$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.9)
+Discretizing (C.2)–(C.4) using the second-order finite difference scheme results in
+$\begin{split}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{j\Delta\xi-1}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (C.7)
+$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.8)
+$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.9)
 where $T_{j}$ is the product temperature at position $j$.
-The ghost point at $j=n_{z}+1$ in (B.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=n_{z}$.
-Similarly, $T_{j=0}$ in (B.9) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=1$.
-The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system, which is then integrated using ode15s.
-Numerical treatment for the secondary drying model are simpler than those described for the primary drying model as there is no moving interface. Define
-$\Delta z=\frac{H}{n_{z}-1},$ (B.10)
-$z=j\Delta z,$ (B.11)
-where $\Delta z$ is the distance between each grid point. Discretizing (38), (39), (41), and (42) using the second-order finite difference scheme results in
-$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (B.12)
-$\dfrac{dc_{\textrm{w},j}}{dt}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (B.13)
-$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.14)
-$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.15)
+The ghost point at $j=n_{z}+1$ in (C.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (C.7) for $j=n_{z}$.
+Similarly, $T_{j=0}$ in (C.9) can be eliminated by substituting $T_{j=0}$ into the governing equation (C.7) for $j=1$.
+The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system.
+Numerical treatment for the secondary drying model is simpler than those described for the primary drying model as there is no moving interface. Define
+$\Delta z=\frac{H}{n_{z}-1},$ (C.10)
+$z=(j-1)\Delta z,$ (C.11)
+where $\Delta z$ is the distance between each grid point, $j=1$ at $z=0$, and $j=n_{z}$ at $z=H$. Discretizing (42), (43), (45), and (46) using the second-order finite difference scheme results in
+$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (C.12)
+$\dfrac{dc_{\textrm{w},j}}{dt}=-f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (C.13)
+$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.14)
+$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.15)
 where the ghost points can be treated similarly as done for the primary drying model.
@@
 $c^{*}_{\textrm{w}}$ equilibrium concentration of bound water (kg water/kg solid)
+$D$ diffusivity (m2/s)
+$D_{\textrm{e}}$ effective diffusivity (m2/s)
+$D_{\textrm{e,g}}$ effective diffusivity for ordinary diffusion of gases (m2/s)
+$D_{\textrm{e,K}}$ effective diffusivity for Knudsen diffusion (m2/s)
+$D_{\textrm{g}}$ binary diffusion coefficient of gases (m2/s)
+$D_{\textrm{K}}$ Knudsen diffusion coefficient (m2/s)
 $d$ vial diameter (m)
@@
 $f_{\textrm{a}}$ frequency factor (1/s)
-$H$ height of the product (m)
+$H$ product height/thickness (m)
 $\Delta H_{\textrm{des}}$ heat of desorption (J/kg)
@@
 $j_{\textrm{w}}$ mass flow rate of water vapor (kg/s)
-$j_{\textrm{w,max}}$ maximum mass flow rate of a condenser (kg/s)
+$j_{\textrm{w,max}}$ condenser maximum capacity (kg/s)
 […]
@@
 […]
+$L$ length scale (m)
+$l$ thickness of the ice layer (m)
 $M$ molar mass (kg/mol)
 $m$ mass (kg)
-$m_{\textrm{i,n}}$ mass of ice resulting from the first nucleation (kg)
 […]
@@
 $r$ radial position (m)
+$r_{\textrm{o}}$ radius of the vial (m)
+$\bar{r}$ average pore radius (m)
 $S$ sublimation front position (m)
 $T$ product temperature (K)
-$T_{\textrm{b}}$ bottom shelf temperature (K)
-$T_{\textrm{c}}$ chamber wall temperature (K)
-$T_{\textrm{f}}$ freezing point (K)
-$T_{\textrm{g}}$ cryogenic gas temperature (K)
-$T_{\textrm{n}}$ nucleation temperature (K)
 $\overline{T}$ average chamber temperature (K)
 $t$ time (s)
-$t_{\textrm{0}}$ initial time, (s)
+$t_{\textrm{0}}$ initial time (s)
 $t_{\textrm{f1}}$ time when preconditioning completes (s)
@@
 $t_{\textrm{d2}}$ time when secondary drying completes (s)
+$t_{\textrm{dff}}$ time scale for diffusion (s)
+$t_{\textrm{des}}$ time scale for desorption (s)
 $V$ volume (m3)
 $x$ mass fraction (–)
+[…]
 $z$ position in vertical direction (m)
 $\Delta z$ distance between each grid point in the $z$-direction (m)
-$\alpha$ thermal diffusivity, $k/\rho C_{p}$ (m2/s)
+$\epsilon$ porosity
 $\varepsilon$ emissivity (–)
@@
 […]
+$\tau$ tortuosity
 Common subscripts
@@
 avg average
-c chamber (environment)
+c chamber (environment), chamber wall (side)
 b bottom shelf
@@
 s3 side surface of the product
+u upper surface of the chamber
 w water

### v3

@@
-Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization
+Mechanistic Modeling of Continuous Lyophilization for Biopharmaceutical Manufacturing
 Abstract
-Lyophilization (aka freeze drying) is a typical process in pharmaceutical manufacturing used for improving the stability of various drug products, including its recent applications to mRNA vaccines.
-While extensive efforts have been dedicated to shifting the pharmaceutical industry toward continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode.
+Lyophilization (aka freeze drying) is a typical process in (bio)pharmaceutical manufacturing used for improving the stability of various drug products, including its recent applications to mRNA vaccines.
+While extensive efforts have been dedicated to shifting the (bio)pharmaceutical industry toward continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in batch mode.
 This article presents the first mechanistic model for a complete continuous lyophilization process, which comprehensively incorporates and describes key transport phenomena in all three steps of lyophilization, namely freezing, primary drying, and secondary drying.
@@
 Several applications related to model-based process design and optimization of continuous lyophilization are also demonstrated.
-The final model is made available in MATLAB and Julia as an open-source software package called ContLyo, which can ultimately be leveraged for guiding the design and development of future continuous lyophilization processes.
-**footnotetext: Corresponding author. Email: braatz@mit.edu
-Keywords— Lyophilization, Freeze Drying, Suspended Vials, Continuous Manufacturing, Pharmaceuticals
-Introduction
-Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products in pharmaceutical manufacturing [1].
-Various promising applications of lyophilization have been proven, including its recent applications that provide long-term stability for mRNA vaccines (e.g., for COVID-19) to allow for higher storage temperature [2, 3].
-Advancement in lyophilization technology could therefore play a crucial role in the future of the pharmaceutical industry.
+The final model is made available as an open-source software package that can be leveraged for guiding the design and development of future continuous lyophilization processes.
+keywords:
+Lyophilization, Freeze Drying, Suspended Vials, Continuous Manufacturing, Pharmaceuticals
+Prakitr Srisuma Gang Chen Richard D. Braatz*
+{affiliations}
+P. Srisuma, G. Chen, R. D. Braatz
+Massachusetts Institute of Technology, Cambridge, MA 02139, USA
+Email: braatz@mit.edu
+1 Introduction
+Lyophilization, also known as freeze drying, is a low-temperature, low-pressure dehydration process used for improving the stability of various drug products in the biopharmaceutical industry [Fissore2018Review].
+By removing the liquid component (usually water) from the product, the final lyophilized product becomes more stable, hence longer shelf life.
+One of the most recent applications of lyophilization is to provide long-term stability for the mRNA COVID-19 vaccines at room temperature [Muramatsu2022mRNA, Meulewaeter2023mRNA].
+This advancement eliminates the need for an ultra-cold supply chain for maintaining the stability of mRNA, which helps facilitate the storage and distribution of mRNA drug product across the world Advancement in lyophilization technology could therefore play a crucial role in the future of mRNA manufacturing and the biopharmaceutical industry in general.
 A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying.
-During freezing, the product and liquid solvent (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with the remaining fraction (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals [4].
+During freezing, the product and water in a vial are cooled such that most of the liquid (free water) is frozen, with the remaining fraction (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals [Fissore2015Review].
 In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum.
 Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
-Although the current trends in the pharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in a batch mode [5, 6], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [7, 6].
+Although the current trends in the biopharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in batch mode [Meyer2015SpinFreezing, Pisano2019ReviewContLyo], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [Capozzi2019ContLyo_SuspendedVials, Pisano2019ReviewContLyo].
 Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs.
-A number of continuous lyophilization concepts for pharmaceutical products have been proposed, including spray freeze drying [8] and spin freezing [5] (see a comprehensive review of continuous lyophilization technologies in Ref.
-[6]).
-A novel lyophilization technology has been recently proposed by Ref.
-[7], in which vials are suspended and continuously move along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
+A number of continuous lyophilization concepts for (bio)pharmaceutical products have been proposed, including spray freeze drying [Ishwarya2015SprayFD] and spin freezing [Meyer2015SpinFreezing] (see a comprehensive review of continuous lyophilization technologies in [Pisano2019ReviewContLyo]).
+A novel lyophilization technology has been recently proposed by [Capozzi2019ContLyo_SuspendedVials], in which vials are suspended and continuously move along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
 Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes.
 Since the key phenomena in lyophilization are underpinned by heat and mass transfer theories, process models for lyophilization are mostly physics-based rather than data-driven.
-Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [9, 10, 11, 12, 13], primary drying [14, 15, 16, 17, 18, 9, 19, 20, 21, 4, 22], and secondary drying [14, 15, 16, 17, 18, 23, 4, 24, 25].
+Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [Hottot2006Freezing, Nakagawa2007Freezing, Arsiccio2017CrystalSize, Deck2022FreezingLumped, Deck2024Freezing2D], primary drying [Litchfield1979Model, Liapis1994Original, Sadikoglu1997Modeling, Sheehan1998Modeling, Pikal2005Model, Hottot2006Freezing, Veraldi2008SimBatchModels, Pisano2010control, Chen2015FEMmodel, Fissore2015Review, Bano2020LumpedDrying], and secondary drying [Litchfield1979Model, Liapis1994Original, Sadikoglu1997Modeling, Sheehan1998Modeling, Pikal2005Model, Fissore2011SecDryingMonitor, Fissore2015Review, Sahni2017Simplified, Yoon2021Sec0D1D3D].
 Only a few models for continuous lyophilization are available.
-For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by Ref.
-[26].
-A model for the freezing step was also developed for spin freezing [27].
-For spray freeze drying, Ref.
-[28] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases.
+For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by [Bockstal2017ContLyo_ModelPrimary].
+A model for the freezing step was also developed for spin freezing [Nuytten2021SpinFreezing].
+For spray freeze drying, [Sebastiao2019SprayFD] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases.
 Nevertheless, there is no model for the suspended-vial configuration.
@@
 Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies.
-Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature.
+Section 3 extensively describes the development of our mechanistic model, supported by a comprehensive review on a variety of modeling strategies used in the literature.
+Section 4 discussed the numerical methods required for solving and simulating the model equations efficiently.
 Section 5 details the model validation and showcases various applications of the validated model.
 Finally, Section 6 summarizes the study and briefly discusses possible future work.
-Process Description
+2 Process Description
 This work considers lyophilization of unit doses, in which the product is introduced into a vial prior to being lyophilized.
-This type of lyophilization, compared to lyophilization of bulk material, is more preferable in (bio)pharmaceutical applications due to its accurate dosage and better control of sterility [6].
+This type of lyophilization, compared to lyophilization of bulk material, is more preferable in (bio)pharmaceutical applications due to its accurate dosage and better control of sterility [Pisano2019ReviewContLyo].
 Therefore, results and discussion presented in this article are based entirely on lyophilization of unit doses.
-Batch lyophilization
-The majority of lyophilization processes in the pharmaceutical industry have been operated in a batch mode for many decades.
-The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig.
-1A).
+2.1 Batch lyophilization
+The majority of lyophilization processes in the (bio)pharmaceutical industry have been operated in batch mode for many decades.
+The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Figure 1A).
 The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps.
@@
 A number of vials are suspended and move continuously through the lyophilizer.
-Several drawbacks associated with batch lyophilization are extensively discussed in Refs.
-[7, 6].
+Several drawbacks associated with batch lyophilization are extensively discussed in [Capozzi2019ContLyo_SuspendedVials, Pisano2019ReviewContLyo].
 Typical issues of any batch process include batch-to-batch variability, quality control, process downtime, and operational flexibility.
-Another disadvantage that is specific to the configuration illustrated in Fig.
-1A is non-uniformity in heat transfer.
+Another disadvantage that is specific to the configuration illustrated in Figure 1A is non-uniformity in heat transfer.
 For example, vials located on the side and corner of the shelf are typically affected by thermal radiation from the chamber walls, whereas the effect is almost negligible for center vials.
@@
 These known issues motivate the development of a continuous lyophilization concept that can be practically implemented in an industrial scale.
-Continuous lyophilization
-A comprehensive review of continuous lyophilization technologies is given in Ref.
-[6].
+2.2 Continuous lyophilization
+A comprehensive review of continuous lyophilization technologies is given in [Pisano2019ReviewContLyo].
 To summarize, the current continuous lyophilization technologies can be divided into four main categories.
-The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30].
-Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with a high surface-to-volume ratio, greatly improving heat and mass transfer in the process [8, 28].
+The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [Gottfried1973ContLyoPatent] or a revolving plate with holes [Bruttini1993ContLyoPatent].
+Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with a high surface-to-volume ratio, greatly improving heat and mass transfer in the process [Ishwarya2015SprayFD, Sebastiao2019SprayFD].
 Both aforementioned approaches have been proposed for lyophilization of bulk material.
-For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [5].
-Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by Ref.
-[7].
-This technology employs the suspended-vial configuration as illustrated in Fig.
-1B, where all vials are suspended and continuously move through the freezing and drying chambers, without any contact between the vials and shelf.
+For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [Meyer2015SpinFreezing].
+Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by [Capozzi2019ContLyo_SuspendedVials].
+This technology employs the suspended-vial configuration as illustrated in Figure 1B, where all vials are suspended and continuously move through the freezing and drying chambers, without any contact between the vials and shelf.
 Unlike conventional lyophilizers, the freezing and drying processes occur in different locations/chambers of the equipment.
@@
 Second, the process does not involve any complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control.
-Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications.
+Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in (bio)pharmaceutical applications.
 Finally, with the automatic filling and load-lock systems, the process is fully continuous.
-The mechanistic model developed in this work is based on this suspended-vial configuration as illustrated in Fig.
-1B, the current state-of-the-art continuous lyophilization technology for unit doses.
-Detailed modeling strategies and model equations are discussed in the next section (Section 3), while information and discussion related to the equipment design can be found in Ref.
-[7].
-Mechanistic Modeling
-This section details the development of a mechanistic model for continuous lyophilization of suspended vials (Fig.
-1B).
+The mechanistic model developed in this work is based on this suspended-vial configuration as illustrated in Figure 1B, the current state-of-the-art continuous lyophilization technology for unit doses.
+Detailed modeling strategies and model equations are discussed in the next section (Section 3), while information and discussion related to the equipment design can be found in [Capozzi2019ContLyo_SuspendedVials].
+3 Mechanistic Modeling
+This section details the development of a mechanistic model for continuous lyophilization of suspended vials (Figure 1B).
 The section starts by discussing existing models and modeling strategies in the literature, and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
-Review of the existing models and modeling strategies
+3.1 Review of the existing models and modeling strategies
 Mechanistic modeling for conventional/batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry.
 Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials.
-While some past reviews briefly discuss the literature on mechanistic modeling of lyophilization [19, 4], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation.
+While some past reviews briefly discuss the literature on mechanistic modeling of lyophilization [Veraldi2008SimBatchModels, Fissore2015Review], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation.
 Hence, this section first provides a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes with the optimal modeling strategies used in this work.
-The published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
+The published models for continuous lyophilization mentioned in Section 1 are based on the spin and spray freeze drying, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
 Figure 2: Modeling strategies for the freezing, primary drying, and secondary drying steps in lyophilization. The strategies used in this article are highlighted.
@@
 As a result, the number of available models for freezing is more limited compared to that of the drying models.
-Modeling strategies for the freezing step
-Modeling the freezing step is most complicated among the three steps of lyophilization; the blue column in Fig.
-2 highlights notable modeling strategies for this step.
+3.1.1 Modeling strategies for the freezing step
+Modeling the freezing step is most complicated among the three steps of lyophilization; the blue column in Figure 2 highlights notable modeling strategies for this step.
 Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights.
-A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31].
-Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails supercooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point [10, 12].
-A more common but more complicated strategy is to consider the supercooling period and incorporate the heat transfer associated with ice nucleation as done in Ref.
-[9].
-Recently, Ref.
-[12] introduced a state-of-the-art freezing model that incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization using a lumped capacity model.
-The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [13].
+A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [Mills1995HeatTransfer].
+Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails supercooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point [Nakagawa2007Freezing, Deck2022FreezingLumped].
+A more common but more complicated strategy is to consider the supercooling period and incorporate the heat transfer associated with ice nucleation as done in [Hottot2006Freezing].
+Recently, [Deck2022FreezingLumped] introduced a state-of-the-art freezing model that incorporates the stochastic nature of ice nucleation [Colucci2020CrystalSize] and considers heat transfer among multiple vials in batch lyophilization using a lumped capacity model.
+The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [Deck2024Freezing2D].
 In Appendix A, we provide a detailed analysis using the relevant dimensionless group to compare between the lumped capacity model and model with thermal gradients and show that the lumped capacity model is sufficiently accurate for lyophilization of unit doses in general.
 The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the solid structure and mass transfer resistance during sublimation in primary drying.
-While some recent studies have built mechanistic models with the goal of predicting pore size distribution [11, 32], how to best model the specific complex interacting phenomena remains an open research question.
-Modeling strategies for the primary drying step
-Many models for primary drying are available in the literature (see the green column in Fig.
-2).
-The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system.
+While some recent studies have built mechanistic models with the goal of predicting pore size distribution [Arsiccio2017CrystalSize, Colucci2020CrystalSize], how to best model the specific complex interacting phenomena remains an open research question.
+3.1.2 Modeling strategies for the primary drying step
+Many models for primary drying are available in the literature (see the green column in Figure 2).
+The simplest strategy is to use a lumped capacity model [Bano2020LumpedDrying], which does not capture the effects of spatial gradients in the system.
 In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and hence the temperature gradient in the vertical direction is significant.
-A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 9, 35].
-By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [35], and so the model can be implemented easily.
+A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [Dyer1968HeatTransferLimit, Jafar2003HeatTransferLimit, Hottot2006Freezing, Srisuma2023Analytical].
+By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [Srisuma2023Analytical], and so the model can be implemented easily.
 However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer.
-An approach to incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front/interface between the frozen and dried regions, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref.
-[19]).
+An approach to incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front/interface between the frozen and dried regions, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of [Veraldi2008SimBatchModels]).
 In this case, the driving force for mass transfer is dependent on the saturation (equilibrium) pressure at the sublimation interface, which is a function of temperature.
 Hence, the model should be able to accurately predict the spatial variation of the product temperature, including the interface temperature, and hence modeling the heat transfer in the vertical direction is needed.
-Moreover, one of the key considerations during the drying steps is to ensure that the product temperature does not exceed the upper limit (e.g., collapse temperature and glass transition temperature) [1], and thus spatially distributed temperature data are valuable for accurately determining the maximum temperature in the product.
-There are models that consider detailed heat and mass transfer by including full energy and continuity equations, both in 1D [15, 16] and in higher dimensions [17].
-Implementing such high-fidelity models for practical applications, e.g., state estimation and control, is not easy due to their high complexity, computation time, and number of parameters [19, 4].
-Besides, the accuracy of the complex models for primary drying is not significantly different from 1D or simplified models in most cases [4].
+Moreover, one of the key considerations during the drying steps is to ensure that the product temperature does not exceed the upper limit (e.g., collapse temperature and glass transition temperature) [Fissore2018Review], and thus spatially distributed temperature data are valuable for accurately determining the maximum temperature in the product.
+There are models that consider detailed heat and mass transfer by including full energy and continuity equations, both in 1D [Liapis1994Original, Sadikoglu1997Modeling] and in higher dimensions [Sheehan1998Modeling].
+Implementing such high-fidelity models for practical applications, e.g., state estimation and control, is not easy due to their high complexity, computation time, and number of parameters [Veraldi2008SimBatchModels, Fissore2015Review].
+Besides, the accuracy of the complex models for primary drying is not significantly different from 1D or simplified models in most cases [Fissore2015Review].
 Consequently, these high-fidelity models are not commonly used.
 Note that, in primary drying, the water vapor is removed via both sublimation (frozen region) and desorption (dried region).
-Nevertheless, it is widely acknowledged that the amount of water vapor removed via sublimation is much higher than that of desorption, and so the effect of desorption can be omitted [16, 17, 19, 4].
-Modeling strategies for the secondary drying step
+Nevertheless, it is widely acknowledged that the amount of water vapor removed via sublimation is much higher than that of desorption, and so the effect of desorption can be omitted [Sadikoglu1997Modeling, Sheehan1998Modeling, Veraldi2008SimBatchModels, Fissore2015Review].
+3.1.3 Modeling strategies for the secondary drying step
 Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation.
-Several modeling strategies are available, ranging from the simplest lumped capacity model [1] to the high-fidelity model considering multidimensional heat and mass transfer in detail [17].
-In Ref.
-[25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors.
+Several modeling strategies are available, ranging from the simplest lumped capacity model [Fissore2018Review] to the high-fidelity model considering multidimensional heat and mass transfer in detail [Sheehan1998Modeling].
+In [Yoon2021Sec0D1D3D], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors.
 Therefore, 1D modeling is preferable.
@@
 In secondary drying, the two important mechanisms are (1) vapor phase transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product.
-Some previous models for secondary drying consider both phenomena [14, 15, 16, 17], whereas the more recently published models tend to consider only the desorption part [23, 4, 24, 25, 36].
+Some previous models for secondary drying consider both phenomena [Litchfield1979Model, Liapis1994Original, Sadikoglu1997Modeling, Sheehan1998Modeling], whereas the more recently published models tend to consider only the desorption part [Fissore2011SecDryingMonitor, Fissore2015Review, Sahni2017Simplified, Yoon2021Sec0D1D3D, Srisuma2024Obs].
 Simulation results in these studies show that the model prediction is highly accurate even when the vapor phase transport is omitted, suggesting that desorption is the actual limiting step.
-In Ref.
-[37], a detailed experimental study was conducted, concluding that desorption is the rate-limiting step for mass transfer in secondary drying.
+In [Pikal1990SecDrying_Kinetics], a detailed experimental study was conducted, concluding that desorption is the rate-limiting step for mass transfer in secondary drying.
 Evidence from both simulation and experiment suggest that only the desorption part is necessary.
@@
 In Appendix A, we provide a simple and systematic way of analyzing the effect/contribution of each transport process via scale analysis, which can be used to justify the contribution of each transport process for different systems/conditions other than those considered in this work and in the literature.
-Optimal modeling strategies
+3.1.4 Optimal modeling strategies
 High-quality mechanistic modeling should balance between the model accuracy and complexity such that the resulting model can provide reliable results and be practically implemented for different purposes with ease, e.g., optimization, state estimation, and model-based control.
@@
 On the other hand, low-fidelity models are much simpler to implement, but the model prediction might not be sufficiently accurate.
-Our goal is to select modeling strategies that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum.
-From the discussion in Sections 3.1.13.1.3 and Appendix A, the selected modeling strategies are highlighted in Fig.
-2.
+Our goal is to select modeling strategies that can capture the critical phenomena in lyophilization and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum.
+The resulting models should be sufficiently accurate and efficient to be used for general process design (e.g., input/output design, heat/material balance, selecting operating conditions), process optimization, state estimation, and real-time model-based control (e.g., model predictive control).
+For detailed process and equipment design, a high-fidelity model is needed, which is beyond the scope of this work.
+From the discussion in Sections 3.1.13.1.3 and Appendix A, the selected modeling strategies are highlighted in Figure 2.
 The freezing model relies on the lumped capacity method that considers nucleation and supercooling.
@@
 Figure 3: Schematic diagram showing the mechanistic modeling of continuous lyophilization via suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
-Model for freezing
-For the design proposed by Ref.
-[7], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate.
-Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [38].
+3.2 Model for freezing
+For the design proposed by [Capozzi2019ContLyo_SuspendedVials], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate.
+Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [Kramer2002VISF].
 With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf.
-Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification or ice formation, and (5) final cooling.
-Preconditioning
+Our model (Figure 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification or ice formation, and (5) final cooling.
+3.2.1 Preconditioning
 In this model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$.
@@
 where $T(t)$ is the temperature, $t$ is time, $m$ is the total mass, $C_{p}$ is the specific heat capacity (per mass), $Q$ is the total heat transfer between the vial surface and environment, the subscripts ‘w’ and ‘s’ refer to the water (solvent) and solid (solute) phases, and the subscripts ‘s1’, ‘s2’ and ‘s3’ denote the top, bottom, and side of the vial, respectively.
-The initial conditions for (1) are
+The initial conditions for Equation (1) are
 $T(t_{0})=T_{0},$ (2)
@@
 In lyophilization, the bottom shelf and gas temperatures ($T_{\textrm{b}}$ and $T_{\textrm{g}}$) typically vary with time and are specified by the freezing/drying protocol.
-The upper surface and wall temperatures ($T_{\textrm{u}}$ and $T_{\textrm{c}}$), however, are not usually measured and thus can be estimated from data [39].
-In this article, we assume that both temperatures are approximately constant as reported by Refs.
-[40, 19].
+The upper surface and wall temperatures ($T_{\textrm{u}}$ and $T_{\textrm{c}}$), however, are not usually measured and thus can be estimated from data [Srisuma2024Rad].
+In this article, we assume that both temperatures are approximately constant as reported by [Gan2005WallTemp, Veraldi2008SimBatchModels].
 In addition, $T_{\textrm{u}}\approx T_{\textrm{c}}$ due to no additional heat source/sink at these locations.
 Nevertheless, our model is not restricted to these assumptions; i.e., temperature-dependent data can be fed to the model (if available).
-Next, (4)–(6) are simplified further to match the experimental setup in Ref.
-[7], whose data are used for our model validation.
-In Ref.
-[7], the gas inside the chamber was cooled by the bottom shelf, and this cold gas was subsequently used to cool the vials, hence natural convection.
+Next, Equations (4)–(6) are simplified further to match the experimental setup in [Capozzi2019ContLyo_SuspendedVials], whose data are used for our model validation.
+In [Capozzi2019ContLyo_SuspendedVials], the gas inside the chamber was cooled by the bottom shelf, and this cold gas was subsequently used to cool the vials, hence natural convection.
 In this setup, the upper surface and local gas in that area are not cooled, so the gas temperature is assumed to be equal to the surface temperature $T_{\textrm{u}}$.
-Consequently, the two terms can be combined and written in the form of Newton’s law of cooling, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection, which can be done by linearizing the fourth-order term of the radiation part (see more details in Section 1.3.2 of Ref.
-[31] and Section 3.5).
+Consequently, the two terms can be combined and written in the form of Newton’s law of cooling, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection, which can be done by linearizing the fourth-order term of the radiation part (see more details in Section 1.3.2 of [Mills1995HeatTransfer] and Section 3.5).
 Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection.
@@
 $Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T).$ (8)
-At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and so no simplification is needed for (6).
+At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and so no simplification is needed for Equation (6).
 Note that the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system.
-Another design proposed by Ref.
-[7] is to flush the cryogenic gas directly into the chamber.
+Another design proposed by [Capozzi2019ContLyo_SuspendedVials] is to flush the cryogenic gas directly into the chamber.
 In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region.
-Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6) even further.
-Vacuum-induced surface freezing
+Thus, the effect of thermal radiation might be omitted, simplifying Equations (4)–(6) even further.
+3.2.2 Vacuum-induced surface freezing
 At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until its completion at $t_{\textrm{f2}}$.
 The key idea of VISF is to reduce the total pressure to evaporate a small amount of liquid solution from the product, abruptly decreasing the temperature and promoting nucleation.
-A model for VISF is not readily available in the literature but can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref.
-[31].
+An example of available models for VISF can be found in [Pisano2017VISF], which relies on the concept of the condensing/evaporating efficiency.
+A similar approach is used in this work, but the model is derived from the fundamental of heat and mass transfer during evaporation, e.g., as described in [Mills1995HeatTransfer], and thus the associated parameters are the heat and mass transfer coefficients instead of the condensing/evaporating efficiency.
+In addition, the stochastic nature of ice nucleation is also incorporated into the VISF model (described later in Section 3.2.3), allowing for the comparison between spontaneous nucleation and VISF.
 First, consider the mass transfer part.
@@
 The subscripts ‘sat’ and ‘c’ here denote the equilibrium condition and environment, respectively.
-To calculate the mass fraction of water for (9), we first assume that there are two gas/vapor components during VISF, namely (1) water and (2) nitrogen or inert gas, denoted by the subscript ‘in’.
+To calculate the mass fraction of water for Equation (9), we first assume that there are two gas/vapor components during VISF, namely (1) water and (2) nitrogen or inert gas, denoted by the subscript ‘in’.
 By assuming the ideal gas law,
@@
 If the environment contains only nitrogen or inert gas, $p_{\textrm{w,c}}=0$.
-The saturation pressure $p_{\textrm{w,sat}}$ is a function of temperature, that is [41],
+The saturation pressure $p_{\textrm{w,sat}}$ is a function of temperature, that is [Smith2018Thermo],
 $p_{\textrm{w,sat}}=10^{3}\exp\!\left(\!16.3872-\frac{3885.7}{T-42.98}\right).$ (12)
-Note that, if the VISF process is done properly, the amount of water that vaporizes is usually very small and thus does not significantly impact the overall thermophysical properties.
-Next, consider the heat transfer part. The energy balance equation can be modified from (1) as
+Note that, if the VISF process is carried out properly, the amount of water that vaporizes is usually very small and thus does not significantly impact the overall thermophysical properties.
+Next, consider the heat transfer part. The energy balance equation can be modified from Equation (1) as
 $(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (13)
-where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization, which can be approximated by [41]
+where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization, which can be approximated by [Smith2018Thermo]
 $\Delta H_{\textrm{vap}}=2.257\mbox{$\times$}10^{6}\!\left(\frac{1-T/647.1}{1-373.15/647.1}\right)^{\!\!0.38}.$ (14)
-In (13), $m_{\textrm{w}}(t)$ is time-dependent due to evaporation, which also means that the side surface area $A_{r}$ changes with time.
-The initial conditions for (9) and (13) are the final mass of water and temperature at the end of the preconditioning step.
-Nucleation
+In Equation (13), $m_{\textrm{w}}(t)$ is time-dependent due to evaporation, which also means that the side surface area $A_{r}$ changes with time.
+The initial conditions for Equations (9) and (13) are the final mass of water and temperature at the end of the preconditioning step.
+3.2.3 Nucleation
 The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$.
-Our modeling strategy for the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref.
-[12].
+Our modeling strategy for the nucleation step described below is based on the state-of-the-art freezing model proposed by [Deck2022FreezingLumped].
 In this work, we use the term nucleation to denote the first nucleation where the temperature of supercooled liquid almost instantaneously increases to the equilibrium/freezing point, which is caused by the heat released from the fraction of liquid being frozen/solidified.
@@
 where $T_{\textrm{f,w}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute.
-The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (15) and (16) simultaneously.
+The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving Equations (15) and (16) simultaneously.
 Finally, define $t_{\textrm{f3}}$ as the time when the nucleation process completes.
-Since the nucleation process is nearly instantaneous as explained by Ref.
-[12], $t_{\textrm{f3}}$ is set to $t_{\textrm{f2}}$, with the relations
+Since the nucleation process is nearly instantaneous as explained by [Deck2022FreezingLumped], $t_{\textrm{f3}}$ is set to $t_{\textrm{f2}}$, with the relations
 $m_{\textrm{w}}(t_{\textrm{f3}})=m_{\textrm{w}}(t_{\textrm{f2}})-m_{\textrm{i,n}},$ (17)
@@
 $P=1-\exp(\lambda\Delta t).$ (24)
-The probability defined by (24) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$.
-Subsequently, follow the exact same procedure described by (15)–(21), with $t_{\textrm{f1}}$ replacing $t_{\textrm{f2}}$.
-Solidification
+The probability defined by Equation (24) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$.
+Subsequently, follow the exact same procedure described by Equations (15)–(21), with $t_{\textrm{f1}}$ replacing $t_{\textrm{f2}}$.
+3.2.4 Solidification
 After the first nucleation completes, the solidification (aka ice formation) step starts at $t_{\textrm{f3}}$.
-For solidification, the lumped capacity model was shown to provide accurate prediction of the temperature and freezing time for typical vial sizes [12]; we also discuss this in the context of the Biot number in Appendix A.
+For solidification, the lumped capacity model was shown to provide accurate prediction of the temperature and freezing time for typical vial sizes [Deck2022FreezingLumped]; we also discuss this in the context of the Biot number in Appendix A.
 The lumped capacity model is sufficiently accurate and highly computationally efficient, but it does not represent the correct physics of ice formation.
 By treating the entire solid/liquid as a lumped object during solidification, it implies that homogeneous nucleation is assumed.
-Nevertheless, it is commonly known that heterogeneous nucleation is a more frequently observed phenomenon during ice formation [42].
+Nevertheless, it is commonly known that heterogeneous nucleation is a more frequently observed phenomenon during ice formation [Glatz2018Nucleation].
 Consequently, the solidification process generally starts from the vial surfaces, i.e., the outer and bottom surfaces in this case.
-To take this fact into account, a rigorous 2D heat transfer model could be considered [13].
+To take this fact into account, a rigorous 2D heat transfer model could be considered [Deck2024Freezing2D].
 In such cases, the physics is incorporated more accurately, but the major drawback of 2D modeling is its high computational cost and numerical complexity.
@@
 In this case, ice formation is assumed to initiate from the side and bottom surfaces (similar to a 2D model), in which the ice layer is treated as an additional heat transfer resistance between the cold gas and unfrozen water.
-Additionally, heat conduction in the ice layer is assumed to be quasi steady, which is a reasonable assumption for problems associated with phase change in general as the sensible heat is much smaller than the latent heat [31].
+Additionally, heat conduction in the ice layer is assumed to be quasi steady, which is a reasonable assumption for problems associated with phase change in general as the sensible heat is much smaller than the latent heat [Mills1995HeatTransfer].
 Physically, this is equivalent to assuming that all the heat input is used for phase transition only.
 The unfrozen water and solute are treated as a lumped object as its temperature follows the equilibrium temperature.
-A schematic diagram showing our model for the solidification step is shown in Fig.
-4
+A schematic diagram showing our model for the solidification step is shown in Figure 4
 Figure 4: Schematic diagram showing the mechanistic modeling of the solidification step.
@@
 $(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (25)
-with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively.
+with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in Equations (7), (8), and (6), respectively.
 During the solidification step, the amount of water $m_{\textrm{w}}(t)$ and ice $m_{\textrm{i}}(t)$ varies with time, which also affects the volume and surface area of the product.
@@
 $T=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}}\right).$ (27)
-The ice layer becomes an additional heat transfer resistance, and so the heat transfer coefficients $h_{\textrm{s2}}$ and $h_{\textrm{s3}}$ in (8) and (6) can be replaced by the overall heat transfer coefficients
+The ice layer becomes an additional heat transfer resistance, and so the heat transfer coefficients $h_{\textrm{s2}}$ and $h_{\textrm{s3}}$ in Equations (8) and (6) can be replaced by the overall heat transfer coefficients
 $U_{\textrm{s2}}=\frac{1}{1/h_{\textrm{s2}}+l/k_{\textrm{i}}},$ (28)
 $U_{\textrm{s3}}=\frac{1}{1/h_{\textrm{s3}}+\frac{r_{\textrm{o}}\ln(r_{\textrm{o}}/r)}{k_{\textrm{i}}}}.$ (29)
-where $r_{\textrm{o}}$, $r$ and $l$ are as defined in Fig.
-4 and $k_{\textrm{i}}$ is the thermal conductivity of ice.
+where $_{\textrm{o}}, rr$ and $l$ are as defined in Figure 4 and $k_{\textrm{i}}$ is the thermal conductivity of ice.
 For the radiation part, the linearization technique described in Section 3.2.1 can be applied, and so the equation is written in the form of Newton’s law of cooling.
-Note that the overall heat transfer coefficients in (28) and (29) are defined based on the outer surface area to be consistent with the original expressions of $Q_{\textrm{s2}}$ and $Q_{\textrm{s3}}$
-The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely (17), (19), and (21).
-Finally, define $t_{\textrm{f4}}$ as the time when solidification completes. The criterion for complete solidification is defined as
+Note that the overall heat transfer coefficients in Equations (28) and (29) are defined based on the outer surface area to be consistent with the original expressions of $Q_{\textrm{s2}}$ and $Q_{\textrm{s3}}$ .
+The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely Equations (17), (19), and (21).
+Finally, define $t_{\textrm{f4}}$ as the time when solidification completes.
+The criterion for complete solidification is defined as
 $m_{\textrm{i}}(t_{\textrm{f4}})=0.95m_{\textrm{w}}(t_{\textrm{f3}}),$ (30)
-where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [12].
-Our hybrid lumped capacity model presented in this section has the same model complexity and computational cost as those of the original lumped capacity model (e.g., in Ref [12]) while describing the physics in a more accurate way as in the 2D model (e.g., in Ref [13]).
+where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [Deck2022FreezingLumped].
+Our hybrid lumped capacity model presented in this section has the same model complexity and computational cost as those of the original lumped capacity model (e.g., in [Deck2022FreezingLumped]) while describing the physics in a more accurate way as in the 2D model (e.g., in [Deck2024Freezing2D]).
 Final simulations results from these models are almost identical for lyophilization of unit doses, and so there is no concern in terms of model accuracy.
-Cooling
+3.2.5 Cooling
 The final step is to ensure that the product temperature is at the desire value before starting the drying process.
 This cooling step starts at $t_{\textrm{f4}}$ and completes at $t_{\textrm{f5}}$.
-The model for the final cooling step is relatively simple, with the only energy balance equation
+The energy balance is
 $(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}}+m_{\textrm{i}}C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}},$ (31)
-with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively.
+with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in Equations (7), (8), and (6), respectively.
 Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, following the final conditions of the solidification step.
-The initial condition for (31) is the final temperature of the solidification step.
-It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (31) without any significant error.
+The initial condition for Equation (31) is the final temperature of the solidification step.
+It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from Equation (31) without any significant error.
 By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature, phase transition, and amount of ice/water during the freezing step can be predicted.
-Model for primary drying
-In conventional lyophilization, vials are placed on the heating shelf, and thus heat transfer at the bottom surface is driven by conduction (at the point of contact), convection (gas), and radiation, whereas only radiation dominates at the side and top surfaces [18].
+3.3 Model for primary drying
+In conventional lyophilization, vials are placed on the heating shelf, and thus heat transfer at the bottom surface is driven by conduction (at the point of contact), convection (gas), and radiation, whereas only radiation dominates at the side and top surfaces [Pikal2005Model].
 For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf.
 As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
-The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig.
-3B).
+The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Figure 3B).
 Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$.
@@
 where $T_{\textrm{c}}$ is the chamber wall temperature and $A_{r}=\pi dH$ is the side area of the product.
-We refer to Section 3.5 for the detailed derivation of (33).
+We refer to Section 3.5 for the detailed derivation of Equation (33).
 The mass balance of water at the sublimation front gives
@@
 where $N_{\textrm{w}}$ is the sublimation flux and $\rho_{\textrm{e}}$ is the effective density of the dried region above the sublimation front.
-The driving force for mass transfer at the sublimation interface is [18, 1, 22]
+The driving force for mass transfer at the sublimation interface is [Pikal2005Model, Fissore2018Review, Bano2020LumpedDrying]
 $N_{\textrm{w}}=\frac{p_{\textrm{w,sat}}-p_{\textrm{w,c}}}{R_{\textrm{p}}},$ (35)
 where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance.
-The saturation pressure for sublimation is described by [22]
+The saturation pressure for sublimation is described by [Bano2020LumpedDrying]
 $p_{\textrm{w,sat}}=\exp\!\left(\frac{-6139.9}{T}+28.8912\right).$ (36)
-The variation of the mass transfer resistance can be approximated by the empirical expression [18, 4, 22]
+The variation of the mass transfer resistance can be approximated by the empirical expression [Pikal2005Model, Fissore2015Review, Bano2020LumpedDrying]
 $R_{\textrm{p}}=R_{\textrm{p0}}+\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S},$ (37)
 where $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, and $R_{\textrm{p2}}$ are the constants to be estimated from data.
-Physically, $R_{\textrm{p0}}$ represents the resistance associated with mass convection above the product, whereas $\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S}$ corresponds to the resistance associated with diffusion through the porous dried layer in the product.
-The PDE (32) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
-$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (38)
+Physically, $R_{\textrm{p0}}$ represents the resistance associated with mass convection above the product, which could include the presence of a skin layer on the cake surface and the effect of the stopper [Stratta2024Stopper], whereas $\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S}$ corresponds to the resistance associated with diffusion through the porous dried layer in the product.
+Modeling the heat transfer and sublimation front in one dimension, as described in Equations (32) and (34), implies an assumption that the sublimation front is flat.
+This assumption has been widely used for conventional lyophilization as the heat flux through the bottom of the vial is much stronger than that from the sidewall, mainly due to heat conduction.
+In lyophilization of suspended vials, due to the absence of conductive heat transfer at the bottom surface, the sidewall heat flux becomes more comparable to the bottom heat flux.
+However, the bottom heat flux still dominates for two key reasons.
+First, the bottom shelf remains the primary heat source in the system, which directly interacts with the bottom surface of the vial through thermal radiation and convection.
+Second, since most vials are surrounded by others at similar temperatures, the driving force for sidewall heat transfer is significantly weaker than that for bottom heat transfer.
+A systematic analysis for comparing different heat transfer modes can be done using relevant dimensionless numbers, e.g., in Appendix A.
+The PDE represented by Equation (32) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
+$k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (38)
 where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient that combines the effects of thermal radiation and natural convection.
@@
 where $\Delta H_{\textrm{sub}}$ is the heat of sublimation.
-The initial conditions for (32) and (34) are
+The initial conditions for Equations (32) and (34) are
 $T(z,t_{0})$ $=T_{0},\qquad 0\leq z\leq H,$ (40)
@@
 However, the model contains only a simple heat equation, and so it is not detailed here.
-Model for secondary drying
-The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig.
-3C), which is consistent with the primary drying model.
+3.4 Model for secondary drying
+The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Figure 3C), which is consistent with the primary drying model.
 The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$.
@@
 $\rho_{\textrm{e}}C_{p,\textrm{e}}\dfrac{\partial T}{\partial t}=k_{\textrm{e}}\dfrac{\partial^{2}T}{\partial z^{2}}+\rho_{\textrm{d}}\Delta H_{\textrm{des}}\dfrac{\partial c_{\textrm{w}}}{\partial t}+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad 0\leq z\leq H,$ (42)
-where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the concentration of bound water (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the heat of desorption, $Q_{\textrm{rad}}$ is as defined in (33), and the other parameters are as defined in (32), with the subscript ‘e’ denoting the effective properties considering both solid and gas in the pores.
+where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the concentration of bound water (aka moisture content, residual moisture, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the heat of desorption, $Q_{\textrm{rad}}$ is as defined in Equation (33), and the other parameters are as defined in Equation (32), with the subscript ‘e’ denoting the effective properties considering both solid and gas in the pores.
 The desorption kinetics of bound water is described by
 $\dfrac{\partial c_{\textrm{w}}}{\partial t}=k_{\textrm{d}}(c^{*}_{\textrm{w}}-c_{\textrm{w}}),$ (43)
-where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{d}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [15, 16, 4]
+where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{d}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [Liapis1994Original, Sadikoglu1997Modeling, Fissore2015Review]
 $k_{\textrm{d}}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT},$ (44)
 where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency), $E_{\textrm{a}}$ is the activation energy, and $R$ is the gas constant.
-The above desorption kinetics (43) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [15, 43, 19, 23, 4].
+The above desorption kinetics, Equation (43), is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [Liapis1994Original, Sircar2000LDF, Veraldi2008SimBatchModels, Fissore2011SecDryingMonitor, Fissore2015Review].
 Further simplification that is relatively common can be done by setting $c^{*}_{\textrm{w}}=0$.
-This simplification produces insignificant error as shown in Ref.
-[16] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [4].
-The governing PDE (42) requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
-$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (45)
-where the value of $h_{\textrm{b}}$ is approximated to have the same value as that used in primary drying (38).
+This simplification produces insignificant error as shown in [Sadikoglu1997Modeling] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [Fissore2015Review].
+The governing PDE, Equation (42), requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
+$k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (45)
+where the value of $h_{\textrm{b}}$ is approximated to have the same value as that used in primary drying, Equation (38).
 Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
-$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\mathcal{F}_{\textrm{s1}}\big{(}T_{\textrm{u}}^{4}-T^{4}\big{)},\qquad z=0.$ (46)
-The initial conditions for (42) and (43) are
+$k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\mathcal{F}_{\textrm{s1}}\big(T_{\textrm{u}}^{4}-T^{4}\big),\qquad z=0.$ (46)
+The initial conditions for Equations (42) and (43) are
 $T(z,t_{0})=T_{0},\qquad 0\leq z\leq H,$ (47)
@@
 The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step at $t_{\textrm{d2}}$.
-Modeling heat transfer in suspended vials
+3.5 Modeling heat transfer in suspended vials
 Convection and radiation are the two important modes of heat transfer in suspended vials as discussed extensively in the previous sections.
@@
 Heat transfer coefficients for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection).
-Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref.
-[4].
+Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in [Fissore2015Review].
 Thermal radiation is significantly more complicated in terms of modeling.
-The theories and model equations described below are mainly based on Refs.
-[31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization.
+The theories and model equations described below are mainly based on in [Mills1995HeatTransfer] and [Srisuma2024Rad]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization.
 Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; more detailed discussion can be found in the aforementioned references.
@@
 $Q_{\textrm{rad}}=\dfrac{\sigma\big(T_{2}^{4}-T_{1}^{4}\big)}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (50)
-where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 (i.e., $F_{1-2}=1$) and surface 2 is much larger than surface 1, (50) can be simplified as
+where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 (i.e., $F_{1-2}=1$) and surface 2 is much larger than surface 1, Equation (50) can be simplified as
 $Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}\big(T^{4}_{2}-T^{4}_{1}\big).$ (51)
-In (51), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Next, consider the case where a single vial (surface 1) is surrounded by the chamber walls (surface 2). In this case, (51) becomes
+In Equation (51), the transfer factor $\mathcal{F}=\varepsilon_{1}$.
+Next, consider the case where a single vial (surface 1) is surrounded by the chamber walls (surface 2).
+In this case, Equation (51) becomes
 $Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}\big(T^{4}_{\textrm{c}}-T^{4}\big),$ (52)
@@
 Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$.
-The amount of radiative heat for the single-vial case represented by (52) defines the upper bound on the radiative heat that one vial can receive from the chamber walls.
+The amount of radiative heat for the single-vial case represented by Equation (52) defines the upper bound on the radiative heat that one vial can receive from the chamber walls.
 This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among the vials.
 As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$.
-Therefore, (52) needs to be modified to account for interactions between vials.
+Therefore, Equation (52) needs to be modified to account for interactions between vials.
 Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials.
@@
 In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls.
-A rigorous way of modeling this complicated radiation is the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref.
-[39].
-In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by Ref.
-[7].
+A rigorous way of modeling this complicated radiation is the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in [Srisuma2024Rad].
+In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by [Capozzi2019ContLyo_SuspendedVials].
 As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials.
-Therefore, instead of using the radiation network approach, we modify (52) to follow (49) as
+Therefore, instead of using the radiation network approach, we modify Equation (52) to follow Equation (49) as
 $Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\big(T^{4}_{\textrm{c}}-T^{4}\big),$ (53)
 where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph.
-This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (50).
-In our model, the transfer factor $\mathcal{F}_{\textrm{s3}}$ appears in (6), (32), and (42).
+This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using Equation (50).
+In our model, the transfer factor $\mathcal{F}_{\textrm{s3}}$ appears in Equations (6), (32), and (42).
 If the chamber design and vial configuration are identical for both freezing and drying steps, the same value of $\mathcal{F}_{\textrm{s3}}$ can be used for all equations.
-A similar analysis can also be done for $\mathcal{F}_{\textrm{s1}}$ in (39) and (46).
-When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; at the temperature of about 300 K, the error caused by linearization is about 0.1% for the temperature difference of 20 K and 2% for the temperature difference of 100 K (see Section 1.3.2 of Ref.
-[31]), which is tiny.
-The linearized equation can then be combined with the convection part, with the corresponding heat transfer coefficient taking into account of both convection and radiation, e.g., (8), (38), and (45).
-This approximation helps simplify the equations and also facilitates the use of an overall heat transfer coefficient (e.g., (29)) as both convection and radiation parts are written in the form of Newton’s law of cooling.
+A similar analysis can also be done for $\mathcal{F}_{\textrm{s1}}$ in Equations (39) and (46).
+When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; at the temperature of about 300 K, the error caused by linearization is about 0.1% for the temperature difference of 20 K and 2% for the temperature difference of 100 K (see Section 1.3.2 of [Mills1995HeatTransfer]), which is tiny.
+The linearized equation can then be combined with the convection part, with the corresponding heat transfer coefficient taking into account of both convection and radiation, e.g., Equations (8), (38), and (45).
+This approximation helps simplify the equations and also facilitates the use of an overall heat transfer coefficient (e.g., Equation (29)) as both convection and radiation parts are written in the form of Newton’s law of cooling.
 […]
-Default model parameters
+3.6 Default model parameters
 This section defines the default parameter values for the model developed in Sections 3.2, 3.3, and 3.4.
@@
 This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively.
-Table LABEL:Tab:DefaultParameters lists the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
+Table LABEL:ch2-tab:DefaultParameters lists the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
 Table 1: Default model parameters for a complete continuous lyophilization simulation.
@@
 […]
-$\Delta H_{\textrm{des}}$ 2.68$\times$106 J/kg [16]
-$\Delta H_{\textrm{fus}}$ 3.34$\times$105 J/kg [41]
-$\Delta H_{\textrm{sub}}$ 2.84$\times$106 J/kg [19]
-$\Delta H_{\textrm{vap}}$ see (14) J/kg [41]
+[…]
+[…]
+[…]
+$\Delta H_{\textrm{vap}}$ see Equation (14) J/kg [Smith2018Thermo]
 […]
@@
 […]
-$\rho_{\textrm{e}}$ 215 kg/m3 [16]
-$\rho_{\textrm{d}}$ 212.21 kg/m3 [16]
+$\rho_{\textrm{e}}$ 215 kg/m3 [Sadikoglu1997Modeling]
+$\rho_{\textrm{d}}$ 212.21 kg/m3 [Sadikoglu1997Modeling]
 $\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix B)
-$\rho_{\textrm{i}}$ 917 kg/m3 [32]
-$\rho_{\textrm{s}}$ 1587.9 kg/m3 [32]
-$\rho_{\textrm{w}}$ 1000 kg/m3 [13]
+$\rho_{\textrm{i}}$ 917 kg/m3 [Colucci2020CrystalSize]
+$\rho_{\textrm{s}}$ 1587.9 kg/m3 [Colucci2020CrystalSize]
+$\rho_{\textrm{w}}$ 1000 kg/m3 [Deck2024Freezing2D]
 Operating conditions
-$c_{\textrm{w,0}}$ 0.088 kg water/kg solid [7]
-$c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [1]
+$c_{\textrm{w,0}}$ 0.088 kg water/kg solid [Capozzi2019ContLyo_SuspendedVials]
+$c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [Fissore2018Review]
 $p_{\textrm{t}}$ $10^{5}$ (freezing) Pa
 $10^{4}$ (VISF) Pa
-$p_{\textrm{w,c}}$ 3 Pa [7]
-$p_{\textrm{w,sat}}$ see (12) and (36) Pa [41, 22]
+$p_{\textrm{w,c}}$ 3 Pa [Capozzi2019ContLyo_SuspendedVials]
+$p_{\textrm{w,sat}}$ see Equations (12) and (36) Pa [Smith2018Thermo, Bano2020LumpedDrying]
 $T_{0}$ 298.15 K
@@
 290 (secondary drying) K
-$T_{\textrm{f,w}}$ 273.15 K [13]
+$T_{\textrm{f,w}}$ 273.15 K [Deck2024Freezing2D]
 $T_{\textrm{g}}$ 268 (before VISF) K
@@
 $\mathcal{F}_{\textrm{s3}}$ 0.624
-$f_{\textrm{a}}$ 1.5$\times$10-3 1/s
-[…]
-[…]
-$h_{\textrm{m}}$ 6.34$\times$10-3 kg/m2$\cdots$ [31]
 […]
@@
 […]
-$R_{\textrm{p0}}$ 1.5$\times$104 m/s
-$R_{\textrm{p1}}$ 3.0$\times$107 1/s
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 $R_{\textrm{p2}}$ 10 1/m
@@
 Product and formulation
-$H$ 7.2$\times$10-3 m calculated (see Appendix B)
-$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix B)
-$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix B)
-$V_{\textrm{l}}$ 3$\times$10-6 m3 [7]
-$x_{\textrm{s}}$ 0.05 [7]
+[…]
+[…]
+[…]
+[…]
+$x_{\textrm{s}}$ 0.05 [Capozzi2019ContLyo_SuspendedVials]
 Vial
 $d$ 0.024 m 10R vial
-$\epsilon_{\textrm{gl}}$ 0.8 [31]
+$\epsilon_{\textrm{gl}}$ 0.8 [Mills1995HeatTransfer]
 Others
 […]
-$M_{\textrm{in}}$ 0.028 kg/mol [44]
-$M_{\textrm{s}}$ 0.3423 kg/mol [13]
-$M_{\textrm{w}}$ 0.018 kg/mol [44]
+$M_{\textrm{in}}$ 0.028 kg/mol [Chang2010Chem]
+$M_{\textrm{s}}$ 0.3423 kg/mol [Deck2024Freezing2D]
+$M_{\textrm{w}}$ 0.018 kg/mol [Chang2010Chem]
 […]
@@
 Operating conditions that are not measured, e.g., wall temperatures, could also be estimated from data.
-Implementation
-The mechanistic model for continuous lyophilization developed in this article is aimed to be used for various purposes, ranging process design and development to the actual operation (e.g., real-time control, state estimation, and optimization).
-In continuous operation, the process runs continuously with minimal human intervention, and thus efficient (ideally real-time) simulation and computation are important.
-Therefore, appropriate numerical methods must be applied.
-The detailed numerical methods and implementation are given in Appendix C.
+4 Numerical Methods
+This section comprehensively describes the numerical techniques required for solving the model equations efficiently.
+For the freezing model, since the lumped capacity approximation is used, the resulting equations are ordinary differential equations (ODEs).
+In [Deck2022FreezingLumped, Deck2024Freezing2D], the ODEs were solved numerically using an explicit scheme with fixed time steps.
+The probability of nucleation (Equation (24)) was calculated at every time step in a discrete fashion.
+From the computational perspective, this strategy is not the most efficient approach as it requires small step size to ensure stability.
+In this work, we rely on an implicit method with adaptive time steps; MATLAB’s ode15s is selected.
+Nevertheless, since Equation (24) is an algebraic equation that depends on the time step $\Delta t$, solving an ODE system coupled with Equation (24) requires fixing $\Delta t$ via a predefined time span and then running the solver repeatedly and sequentially for each time step, with the probability of nucleation calculated at the end of each run, until the first nucleation is detected.
+A more efficient strategy is to consider a continuous version of Equation (24), which is
+$P=1-\exp\left(\int_{t_{0}}^{t}-\lambda dt\right),$ (54)
+where $t_{0}$ is the initial time and $P(t=t_{0})=0$. Differentiate Equation (54) results in
+$\frac{dP}{dt}=\lambda(1-P).$ (55)
+The resulting ODE, Equation (55), can now be coupled with a system of ODEs describing heat and mass transfer in the process (e.g., during preconditioning and VISF), which can then be integrated using the selected ODE solver.
+In MATLAB, the Event function can be used to detect the time of first nucleation, i.e., when the probability exceeds the given value.
+With this strategy, our freezing model can be integrated continuously in a single run (i.e., without running the solver repeatedly).
+Consequently, the model equations can be solved within 0.1 s on a normal laptop, which is highly efficient.
+For the PDEs in the primary drying and secondary drying models, the method of lines [Schiesser1991MOL] is recommended.
+The method of lines consists of two steps.
+First, the PDEs are discretized spatially to produce a system of ODEs.
+Second, the resulting ODEs are integrated with a proper ODE solver (MATLAB’s ode15s in this work).
+For the primary drying model, first consider Equation (32).
+As the equation involves a moving interface $S(t)$, define a new variable
+$\xi=\frac{z-S}{H-S},$ (56)
+where $\xi$ is the dimensionless position with respect to the moving interface (aka sublimation front).
+This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$.
+Consequently, Equations (32), (38) and (39) become
+$\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (57)
+$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (58)
+$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),\qquad\xi=0.$ (59)
+To spatially discretize the PDE, define
+$\Delta\xi=\frac{1}{n_{z}-1},$ (60)
+$\xi=(j-1)\Delta\xi,$ (61)
+where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$.
+Discretizing Equations (57)–(59) using the second-order finite difference scheme results in
+$\begin{split}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{j\Delta\xi-1}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (62)
+$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (63)
+$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (64)
+where $T_{j}$ is the product temperature at position $j$.
+The ghost point at $j=n_{z}+1$ in Equation (63) can be eliminated by substituting $T_{j=n_{z}+1}$ into Equation (62) for $j=n_{z}$.
+Similarly, $T_{j=0}$ in Equation (64) can be eliminated by substituting $T_{j=0}$ into Equation (62) for $j=1$.
+The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system.
+Numerical treatment for the secondary drying model is simpler than those described for the primary drying model as there is no moving interface. Define
+$\Delta z=\frac{H}{n_{z}-1},$ (65)
+$z=(j-1)\Delta z,$ (66)
+where $\Delta z$ is the distance between each grid point, $j=1$ at $z=0$, and $j=n_{z}$ at $z=H$.
+Discretizing Equations (42), (43), (45), and (46) using the second-order finite difference scheme results in
+$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (67)
+$\dfrac{dc_{\textrm{w},j}}{dt}=-f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (68)
+$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (69)
+$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (70)
+where the ghost points can be treated similarly as done for the primary drying model.
 All simulations, calculations, and results presented in this article were performed and generated using MATLAB R2023a.
+With the aforementioned numerical methods, our model can be simulated accurately within 1 s on a normal laptop.
 The model, data, and MATLAB code used in this work are made available (see the Data Availability section).
-In addition to the MATLAB implementation, we also provide an equivalent software package in Julia, a high-level, open-source programming language for scientific computing, for the selected simulations.
-Results and Discussion
-Model validation
-The mechanistic model presented in Section 3 is validated using the experimental data reported in Ref.
-[7].
-Our mdel validation is carried out for each lyophilization step separately, with the specific parameters reported in Table 2.
+5 Results and Discussion
+5.1 Model validation
+The mechanistic model presented in Section 3 is validated using the experimental data reported in [Capozzi2019ContLyo_SuspendedVials].
+Our model validation is carried out for each lyophilization step separately, with the specific parameters reported in Table LABEL:ch2-tab:ParametersValidation.
 Table 2: Parameters used for the model validation, including Cases 1, 2a, 2b, 3a, and 3b.
@@
 Freezing (no VISF)
-$T_{0}$ 280 K [7]
-$T_{\textrm{c}},T_{\textrm{u}}$ 272 K estimated from data in Ref. [7]
-$T_{\textrm{g}}$ see Fig. 5A K [7]
-$T_{\textrm{n}}$ 263.18 K [7]
+$T_{0}$ 280 K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{c}},T_{\textrm{u}}$ 272 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{g}}$ see Figure 5A K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{n}}$ 263.18 K [Capozzi2019ContLyo_SuspendedVials]
 […]
@@
 Primary drying
-$T_{0}$ 231 (Case 2a) K [7]
-235 (Case 2b) K [7]
-$T_{\textrm{b}}$ 263 (Case 2a) K [7]
-313 (Case 2b) K [7]
-$T_{\textrm{c}},T_{\textrm{u}}$ 275 K estimated from data in Ref. [7]
-[…]
-$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [7]
-$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [7]
-Secondary drying
-$c_{\textrm{w,0}}$ 0.088 (Case 3a) kg water/kg solid [7]
-0.075 (Case 3b) kg water/kg solid [7]
-$T_{0}$ 273 K [7]
-$T_{\textrm{b}}$ 293 K [7]
-$T_{\textrm{c}},T_{\textrm{u}}$ 285 K estimated from data in Ref. [7]
-$f_{\textrm{a}}$ 0.42 1/s estimated from data in Ref. [7]
+$T_{0}$ 231 (Case 2a) K [Capozzi2019ContLyo_SuspendedVials]
+235 (Case 2b) K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{b}}$ 263 (Case 2a) K [Capozzi2019ContLyo_SuspendedVials]
+313 (Case 2b) K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{c}},T_{\textrm{u}}$ 275 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
 […]
 […]
-$V_{\textrm{l}}$ 2$\times$10-6 m3 [7]
-Figure 5: Model validation for the lyophilization of suspended vials using the experimental data from Ref.
-[7].
+$R_{\textrm{p2}}$ 1 1/m estimated from data in [Capozzi2019ContLyo_SuspendedVials]
+Secondary drying
+$c_{\textrm{w,0}}$ 0.088 (Case 3a) kg water/kg solid [Capozzi2019ContLyo_SuspendedVials]
+0.075 (Case 3b) kg water/kg solid [Capozzi2019ContLyo_SuspendedVials]
+$T_{0}$ 273 K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{b}}$ 293 K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{c}},T_{\textrm{u}}$ 285 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
+$f_{\textrm{a}}$ 0.42 1/s estimated from data in [Capozzi2019ContLyo_SuspendedVials]
+[…]
+[…]
+[…]
+Figure 5: Model validation for the lyophilization of suspended vials using the experimental data from [Capozzi2019ContLyo_SuspendedVials].
 Panel A shows the model prediction and experimental data for the product temperature during the freezing step.
@@
 The initial concentrations for Cases 3a and 3b are 0.088 and 0.075 kg water/kg solid, respectively.
-For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig.
-5A).
+For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Figure 5A).
 The product temperature decreases from its initial value of 280 K to the nucleation temperature of about 263 K, where nucleation starts.
@@
 The temperature slowly decreases during the solidification phase for about 30 min, following the freezing-point depression.
-Subsequently, the temperature decreases faster and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
-The model accurately predicts the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig.
-5B).
+Subsequently, the temperature decreases faster and reaches equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
+The model accurately predicts the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Figure 5B).
 The temperature rises up rapidly at the beginning and gradually increases at a slower rate toward the end of the process.
@@
 This error could be partly because of the uncertainty in cake resistance, which results from the stochastic nature of the freezing process.
-In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.088 and 0.075 kg water/kg solid).
+The datasets in Figure 5B, which represent the lower and upper limits of the operating boundaries, demonstrate the validity and generalizability of the model across the full range of operating conditions.
+Similar to primary drying, in secondary drying, we consider two different concentration levels for the bound water (residual moisture): 0.088 and 0.075 kg water/kg solid.
+The simulated concentration of bound water closely aligns with the experimental data for both concentration levels (Figure 5C).
 The concentration decreases exponentially from its initial value to the final concentration of 0.01 kg water/kg solid, the threshold set for terminating the simulations.
 This concentration profile is expected given the linear driving force model is used to describe the desorption process.
-It is important to note that the data used in Fig.
-5C are based on the shelf temperature of 293 K.
-To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures so that the effects of temperature on the desorption process can be quantified.
-Prediction of critical process parameters for a complete continuous lyophilization cycle
+It is important to note that the data used in Figure 5C are based on the shelf temperature of 293 K.
+To rigorously estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures so that the effects of temperature on the desorption process can be quantified.
+Overall, our model is able to accurately predict the time evolution of the critical process parameters for all three steps of lyophilization.
+The maximum deviation between the predicted and measured temperature is about 3 K, which is about 1% compared to the absolute operating temperature used in lyophilization.
+This error is smaller than the measurement noise of some non-invasive temperature sensors, e.g., thermal imaging cameras [Srisuma2023Analytical] For the concentration of bound water, the maximum error is less than 0.01 kg water/kg solid, which is typically the threshold for identifying the endpoint of secondary drying [Fissore2018Review] Therefore, the accuracy of our model is sufficient for general process design, optimization, and control.
+We note that, in the suspended-vial lyophilization considered in this article, every single vial moves through the process following the same trajectory as shown in Figure 1B, ensuring identical heat transfer conditions as a function of time for all vials.
+For batch lyophilization designs where the vials are distributed in different locations (Figure 1A), the evolution of the critical process parameters in each vial also varies.
+For instance, the outermost vials in a batch lyophilizer dry the fastest due to thermal radiation from the environment, while the thermal radiation effect is much weaker on the inner vials [Srisuma2024Rad].
+5.2 Simulation of a complete lyophilization cycle
 One of the important aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production.
@@
 This is important specifically for continuous manufacturing where the entire process should be designed and optimized simultaneously.
-With the default model parameters in Table LABEL:Tab:DefaultParameters, a complete continuous lyophilization cycle can be simulated, including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig.
-6).
-The time evolution of the product temperature (Fig.
-6A), mass of ice (Fig.
-6B), and concentration of bound water (Fig.
-6C) can be obtained, given the operating pressure (Fig.
-6D) and temperature (Fig.
-6E).
+With the default model parameters in Table LABEL:ch2-tab:DefaultParameters, a complete continuous lyophilization cycle can be simulated, including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Figure 6).
+The time evolution of the product temperature (Figure 6A), mass of ice (Figure 6B), and concentration of bound water (Figure 6C) can be obtained, given the operating pressure (Figure 6D) and temperature (Figure 6E).
 This information could be useful for various purposes.
@@
 Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time.
-It is important to note that the entire simulation shown in Fig.
-6 can be computed in less than 1 s on a normal laptop, which makes the model very practical to be implemented for any purposes.
+It is important to note that the entire simulation shown in Figure 6 can be computed in less than 1 s on a normal laptop, which makes the model very practical to be implemented for any purposes.
 Figure 6: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) average concentration of bound water, (D) operating pressure, and (E) operating temperature.
-Visualizing spatiotemporal data
+5.3 Visualizing spatiotemporal data
 As discussed in Section 3.1, one of the key considerations during the drying steps is to ensure that the product temperature at any location does not exceed the upper limit.
@@
 The initial concentration is assumed to vary linearly from 5% (0.05 kg water/solid) at the top to 20% (0.2 kg water/kg solid) at the bottom.
-Application to freezing: Optimizing freezing protocols
-This section employs the model to provide insights into the freezing step, which usually receives little attention compared to the drying steps.
-As spontaneous nucleation was already investigated in detail by Refs.
-[12, 13], controlled nucleation, which is preferable for continuous lyophilization, is considered instead.
-The evolution of the product temperature and amount of water during VISF at different pressures predicted by our model is shown in Fig.
-9.
-By varying the pressure profile, it is possible to control the rate of change in the product temperature, hence nucleation.
-The model can also be used to estimate the amount of water that evaporates during VISF as well as the rate of evaporation.
-This information could be useful for the development of a freezing protocol for controlled nucleation.
-Figure 9: Simulation results showing the (A) evolution of the product during temperature and (B) amount of water during VISF at three different pressures, namely $10^{4}$, 5000, and 1000 Pa.
-At $t=0$, the pressure decreases linearly from $10^{5}$ Pa to the target pressure in 30 s before being held constant.
-The nucleation temperature is set to 263 K.
-External heat transfer with the wall and cryogenic gas is omitted from the model so that only the contribution from VISF is observed.
-Other model parameters are kept at the default values.
-Application to drying: Modifying the model for conventional lyophilization
-Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization.
-The major modifications are to turn off the thermal radiation part in the model and recalculate/estimate the heat transfer coefficient and cake resistance.
-In this case study, we consider batch primary drying and secondary drying using the data from Refs.
-[19] and [4], with the specific parameters reported in Table 3.
-Table 3: Parameters used for the conventional drying studies, including for Cases 4a, 4b, and 5.
-Symbol Value Unit Source
-Primary drying
+5.4 Understanding and optimizing VISF
+A key benefit of mechanistic models is that they can sometimes provide important insights about the process when the experimental data are limited.
+This section employs the model and simulation to study the freezing step, which usually receives little attention compared to the drying steps.
+Spontaneous (aka uncontrolled) ice nucleation was already investigated in detail by [Deck2022FreezingLumped, Deck2024Freezing2D].
+However, uncontrolled nucleation is not ideal for continuous lyophilization as described in [Capozzi2019ContLyo_SuspendedVials].
+Therefore, we focus on controlled nucleation, with the VISF technique as used in [Capozzi2019ContLyo_SuspendedVials].
+The input and parameter data are given in Tables 3 and 4.
+Table 3: Parameters used for the modeling of VISF in Figures 9A–D.
+Symbol Value Unit
+$p_{\textrm{t}}$ see1 Pa
+$T_{0}$ 280 K
+$T_{\textrm{g}},T_{\textrm{c}},T_{\textrm{u}}$ see2 K
 […]
 […]
-$\rho_{\textrm{e}}$ 252 kg/m3 [19]
-$\rho_{\textrm{f}}$ 917 kg/m3 [19]
-$p_{\textrm{w,c}}$ 15 Pa [19]
-$T_{0}$ 228.15 K [19]
-$T_{\textrm{b}}$ see1 K [19]
-$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
+  • 1
+    $p_{\textrm{t}}$ is initially constant at $10^{5}$ Pa and then decreases linearly to the VISF pressure (different values are used in this study) in 1 min after VISF starts at $t$ = 0.25 h.
+  • 2
+    $T_{\textrm{g}},T_{\textrm{c}}$, and $T_{\textrm{u}}$ are initially constant at 268 K and then decrease linearly to 260 K in 1 min after VISF starts at $t$ = 0.25 h.
+Table 4: Parameters used for the validation of our VISF model in Figure 9E.
+Symbol Value Unit Source
+$p_{\textrm{t}}$ see Figure 3 in [Capozzi2019ContLyo_SuspendedVials] Pa [Capozzi2019ContLyo_SuspendedVials]
+$T_{0}$ 268.27 K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{g}}$ see Figure 3 in [Capozzi2019ContLyo_SuspendedVials] K [Capozzi2019ContLyo_SuspendedVials]
+$T_{\textrm{c}},T_{\textrm{u}}$ 282 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
 […]
-$R_{\textrm{p0}}$ 3.8$\times$104 m/s estimated from data in Ref. [19]
-$R_{\textrm{p1}}$ 3.1$\times$107 1/s estimated from data in Ref. [19]
-$R_{\textrm{p2}}$ 10 1/m estimated from data in Ref. [19]
-$H$ 7.15$\times$10-3 m [19]
-Secondary drying
-$c_{\textrm{w,0}}$ 0.0603 kg water/kg solid [4]
-$T_{0}$ 264 K [4]
-$T_{\textrm{b}}$ see2 K estimated from data in Ref. [4]
-$f_{\textrm{a}}$ 1.2$\times$10-3 1/s estimated from data in Ref. [4]
-[…]
-$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
-[…]
-$V_{\textrm{l}}$ 1.5$\times$10-6 m3 [4]
-$d$ 0.01425 m 2R vial, [4]
-  • 1
-    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 K (Case 4b).
-  • 2
-    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264 K to 312 K for Case 5.
-Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig.
-10).
-The model can also accurately predict the product temperature and concentration of bound water during the secondary drying step (Fig.
-11).
-With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
-Figure 10: Comparison of our model prediction and the experimental/simulation data from Ref.
-[19] for the primary drying step in conventional lyophilization.
-Panel A shows the model prediction and experimental data for the product temperature at the bottom surface.
-Panel B shows the model prediction and reference simulation data for the sublimation front position.
-The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15 K, while other conditions are identical.
-Figure 11: Comparison of our model prediction and the experimental data from Ref.
-[4] for the secondary drying step in conventional lyophilization.
-Panel A shows the model prediction and experimental data for the product temperature at the bottom surface.
-Panel B shows the model prediction and experimental data for the average concentration of bound water.
-Condenser failure analysis
+Figure 9: Controlled nucleation via vacuum induced surface freezing (VISF).
+(A) Product temperature profiles at different VISF pressures when the nucleation temperature is fixed at 260 K.
+(B) Nucleation time and mass loss due to evaporation at different VISF pressures when the nucleation temperature is fixed at 260 K.
+(C) Product temperature profiles for uncontrolled nucleation and VISF at 5000 Pa and 100 Pa considering stochastic ice nucleation.
+(D) Nucleation time at different VISF pressures considering stochastic ice nucleation.
+(E) Comparison between the simulated temperature profile during VISF and experimental data from [Capozzi2019ContLyo_SuspendedVials].
+Note that the nucleation time here is defined as the time when the first nucleation occurs.
+Several experimental studies on VISF have been conduced in the literature.
+Here, we use our model and simulation results to provide insights into the process, and use that understanding to explain the previous experimental data.
+To clearly understand VISF, we first consider cases where that the nucleation temperature is fixed at 260 K; i.e., the effect of stochastic ice nucleation is excluded from the model (Figures 9AB).
+In VISF, the key idea is to reduce the system pressure, typically from the atmospheric pressure ($10^{5}$ Pa) to the VISF pressure, to evaporate a small amount of liquid, resulting in a fast decrease in the product temperature and hence nucleation.
+The product temperature decreases faster with a lower VISF pressure (Figure 9A).
+For example, when VISF starts at 0.25 h, the temperature drops from the initial value of 268 K to the nucleation temperature of 260 K almost instantaneously for the VISF pressure of 100 Pa, while it takes about 10 min when the VISF pressure is $10^{4}$ Pa.
+Consequently, the first nucleation occurs earlier at a lower VISF pressure (dash line in Figure 9B).
+During VISF, a small amount of liquid evaporates, which increases with a decrease in the VISF pressure (solid line in Figure 9B).
+This is expected because of a higher mass transfer driving force for evaporation at lower VISF pressure.
+In general, if VISF is carried out properly, the amount of liquid evaporating is small.
+In this case, the original liquid mass is about $2.9\times 10^{-3}$ kg, and so the mass loss of about $3\times 10^{-5}$ is almost negligible.
+Next, we include the effect of stochastic ice nucleation into our VISF model for a more realistic analysis (Figures 9CD), with a Monte Carlo simulation of $10^{4}$ runs for proper statistics.
+Our simulation result show that VISF can significantly reduce the variation in the product temperature, compared to the uncontrolled nucleation case (Figure 9C).
+A decrease in the VISF pressure also reduces the degree of variation.
+Besides the product temperature, The variation in the nucleation time111defined as the time when the first nucleation occurs also decreases with the VISF pressure, with the first nucleation occurring earlier (Figure 9D).
+These results indicate that the nucleation process is well controlled with VISF (i.e., less variation in the nucleation time and temperature), agreeing with general experimental observation in the literature.
+Finally, for model validation and implementation in real-world applications, the model can be fitted to experimental data obtained from the system of interest for more accurate results; the two important parameters are the mass transfer coefficient and heat transfer coefficient associated with evaporation.
+For example, by using the data from [Capozzi2019ContLyo_SuspendedVials], the temperature profile simulated by our model agrees well with the experimental data (Figure 9E).
+The results presented in this section assume minimal vial-to-vial thermal interactions because the vials are arranged in a single row, in which all the vials experience the same heat transfer conditions throughout the process, as described in Section 3.5.
+This condition is not true in conventional lyophilization where a large number of vials are arranged in a hexagonal or rectangular array.
+In such cases, vial-to-vial interactions could significantly affect the distribution of nucleation times and freezing rates.
+For example, when one vial nucleates, an increase in the temperature of that vial could slightly heat the adjacent vials, which subsequently delays the first nucleation of those vials.
+The effects of vial-to-vial interactions on freezing have been discussed in [Deck2022FreezingLumped, Deck2024Freezing2D, Pisano2025Freezing].
+In the context of modeling, these interactions can be incorporated straightforwardly by coupling our freezing model with the radiation network approach described in Section 3.5 and [Srisuma2024Rad].
+In any case, such effects become much less important in controlled nucleation.
+In summary, this section demonstrates how the model and simulation results elucidate the role of VISF and its operating conditions in promoting controlled nucleation and influencing the nucleation process, improving the uniformity of the product.
+Furthermore, we show that the model can explain real data relatively well.
+Hence, our model can be used to understand and guide the design and optimization of the VISF method for controlling the nucleation process in lyophilization.
+5.5 Analysis of condenser failure
 Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily under the desired conditions.
@@
 In lyophilization, one of the crucial design considerations is associated with a condenser.
-During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1].
+During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [Fissore2018Review].
 This operational constraint is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase.
 This scenario may not lead to safety-related issues because the operating pressure in lyophilization is low and the amount of ice/water is generally not high enough to create overpressure.
-Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (35)), which then prolongs the primary drying step.
+Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see Equation (35)), which then prolongs the primary drying step.
 As a result, the final product quality could be affected.
@@
 To incorporate the condenser dynamics into the model, we consider cases where the total condenser capacity is not enough to condense the water vapor produced via sublimation during primary drying.
-In such cases, the mass balance for the water vapor in a chamber is [22]
+In such cases, the mass balance for the water vapor in a chamber is [Bano2020LumpedDrying]
 $\frac{dp_{\textrm{w,c}}}{dt}=\frac{(j_{\textrm{w}}-j_{\textrm{w,max}})R\overline{T}}{V_{\textrm{c}}M_{\textrm{w}}},$ (71)
@@
 where $n_{\textrm{vial}}$ is the number of vials in the chamber.
-By coupling (54) and (55) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
-With the specific parameters given in Table 4, the total pressure increases from the normal operating value of 3 Pa to 20 Pa (Fig.
-12A).
+By coupling Equations (71) and (72) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
+With the specific parameters given in Table 5, the total pressure increases from the normal operating value of 3 Pa to 20 Pa (Figure 10A).
 This pressure increase results from the total vapor flow exceeding the maximum condenser capacity, thus vapor accumulation in the chamber.
@@
 After about 1 h, the system pressure becomes constant, indicating the equalization of the sublimation flux and condenser capacity.
-In this abnormal operation, the drying time and product temperature are higher than for the normal operation (Figs.
-12BC).
+In this abnormal operation, the drying time and product temperature are higher than for the normal operation (Figures 10BC).
 Table 5: Parameters used for the condenser failure analysis.
@@
 $n_{\textrm{vial}}$ 200
-$V_{\textrm{c}}$ 0.118 m3 [22]
-$j_{\textrm{w,max}}$ 1.8$\times$10-5 m3/s [22]
+$V_{\textrm{c}}$ 0.118 m3 [Bano2020LumpedDrying]
+[…]
 Figure 10: Simulation results comparing the normal operation and condenser failure scenario for the (A) system pressure, (B) sublimation front position, and (C) product average temperature during primary drying.
-Conclusion
+6 Conclusion
 This article proposes the first mechanistic model for continuous lyophilization, in which the state-of-the-art suspended-vial technology is considered.
@@
 With a well-developed mechanistic model, future work could consider utilizing the model for several important tasks, e.g., state estimation, optimal control design, and uncertainty quantification, which will ultimately support the development of a high-quality digital twin for continuous lyophilization to advance the manufacturing process.
-Data Availability
-All software and data will be made available upon publication of the manuscript. Both MATLAB and Julia implementations will be provided.
+Data and Code Availability
+Software and data used in this work are available at https://github.com/PrakitrSrisuma/ContLyo-modeling.
+All simulations, calculations, and results presented in this article were performed and generated using MATLAB, but an equivalent Julia implementation of the code is also provided.
 Acknowledgments
@@
 Appendix A Analysis of Transport Phenomena
-Biot number and spatial dimensions
+A.1 Biot Number and Spatial Dimensions
 As discussed in the main text, the freezing step is modeled using the lumped capacity method, whereas the drying steps are analyzed in one spatial dimension, a vertical direction.
-Various studies have shown via simulation that variation in the radial direction is negligible [17, 13].
+Various studies have shown via simulation that variation in the radial direction is negligible [Sheehan1998Modeling, Deck2024Freezing2D].
 This section provides a detailed analysis from the theoretical perspective to justify this assumption and briefly discusses strategies to modify the model for situations where this assumption is not true.
@@
 In general, the length scale is defined as the ratio of the volume to the surface area, which results in $r_{\textrm{o}}/2$ for a long cylinder.
-However, for conservative analysis, it is recommended to use the actual radius $r_{\textrm{o}}$ instead [45].
-The Biot number is defined as the ratio of the internal conduction resistance to the external convection resistance, and so a small Bi value implies that the resulting temperature gradient should also be small [31].
-In various heat transfer textbooks, the value of 0.1 is recommended as a guideline; if Bi $<0.1$, a lumped capacity model can be used.
-However, in practice, a lumped capacity model could still be valid even for Bi $>0.1$, depending on the system, which is analyzed here.
+However, for conservative analysis, it is recommended to use the actual radius $r_{\textrm{o}}$ instead [Incropera2007HeatTransfer].
+The Biot number is defined as the ratio of the internal conduction resistance to the external convection resistance, and so a small Bi value implies that the resulting temperature gradient should also be small [Mills1995HeatTransfer].
+In various heat transfer textbooks, the value of 0.1 is recommended as a guideline; if Bi <0.1<0.1, a lumped capacity model can be used.
+However, in practice, a lumped capacity model could still be valid even for Bi >0.1>0.1, depending on the system, which is analyzed here.
 […]
@@
 Figure A.2: Comparison between the 1D and lumped capacity model for the (A) volume-averaged temperature and (B) total heat removed for the conservative case.
-Finally, the errors (differences) between both models for different values of Bi are shown in Fig.
-A.3.
+Finally, the errors (differences) between both models for different values of Bi are shown in Figure A.3.
 Of course, if the Biot is too high, the lumped capacity model will not be accurate.
 That situation could happen, e.g., when a very large vial is used or the heat transfer coefficient is much higher than usual.
-In such cases, a high-fidelity model might be needed [13].
+In such cases, a high-fidelity model might be needed [Deck2024Freezing2D].
 However, the analysis in this section suggests that a lumped capacity model is highly accurate for typical lyophilization applications, in particular for unit doses.
 Figure A.3: Errors between the 1D and lumped capacity model for the (A) volume-averaged temperature and (B) total heat removed for different Bi.
-Mass transport in the dried product
+A.2 Mass Transport in the Dried Product
 In secondary drying, the two important mechanisms for mass transport are (1) vapor phase or water vapor transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product.
-Previous simulation and experimental data suggest that desorption is the rate-limiting step [37, 23, 4, 24, 25, 36], and so the equations/terms related to vapor phase transport are typically omitted from recently published models, including our model.
+Previous simulation and experimental data suggest that desorption is the rate-limiting step [Pikal1990SecDrying_Kinetics, Fissore2011SecDryingMonitor, Fissore2015Review, Sahni2017Simplified, Yoon2021Sec0D1D3D, Srisuma2024Obs], and so the equations/terms related to vapor phase transport are typically omitted from recently published models, including our model.
 This section provides an analysis from the perspective of transport phenomena to justify this assumption.
 First, consider the vapor phase transport through the porous dried product.
-Modeling the transport of liquids and gases in porous media is challenging as it could entail a variety of mechanisms besides ordinary diffusion (e.g., Knudsen diffusion), and thus no fully satisfactory theory currently exists [46].
-Nevertheless, relative success has been achieved by using the concept of an effective diffusivity; in this analysis, we follow the procedure described in Ref.
-[31].
+Modeling the transport of liquids and gases in porous media is challenging as it could entail a variety of mechanisms besides ordinary diffusion (e.g., Knudsen diffusion), and thus no fully satisfactory theory currently exists [Bird2002Transport].
+Nevertheless, relative success has been achieved by using the concept of an effective diffusivity; in this analysis, we follow the procedure described in [Mills1995HeatTransfer].
 Define the effective diffusivity $D_{\textrm{e}}$ by
@@
 where $\bar{r}$ is the average pore radius, $T$ is the temperature, and $M$ is the molar mass in g/mol.
-With the effective diffusivity calculated by (A.2)–(A.5), the total mass flux can be calculated using Fick’s law of diffusion, with the binary diffusion coefficient replaced by the effective diffusivity.
+With the effective diffusivity calculated by Equations (A.2)–(A.5), the total mass flux can be calculated using Fick’s law of diffusion, with the binary diffusion coefficient replaced by the effective diffusivity.
 Then, a continuity (mass balance) equation for the water vapor can be written and coupled to the mechanistic model proposed in the main paper.
@@
 Nevertheless, additional information is needed, including the porosity, tortuosity, and average pore radius.
-A more complicated strategy that is used in the literature is to apply the dusty gas model, which is derived from the Maxwell-Stefan Equation (see examples in [16, 19]).
+A more complicated strategy that is used in the literature is to apply the dusty gas model, which is derived from the Maxwell-Stefan Equation (see examples in [Sadikoglu1997Modeling, Veraldi2008SimBatchModels]).
 This approach, however, requires many more parameters and equations.
 Next, we compare the significance of vapor phase transport and desorption.
-Scale analysis can be used to analyze the contribution of each transport process (see various examples in [31]).
+Scale analysis can be used to analyze the contribution of each transport process (see various examples in [Mills1995HeatTransfer]).
 Properly estimating the time scale of each phenomenon can help identify the key process that should be considered in the model and omit unimportant phenomena.
-The time scale (aka time constant) for mass diffusion can be approximated by [47, 31]
+The time scale (aka time constant) for mass diffusion can be approximated by [Crank1979Diffusion, Mills1995HeatTransfer]
 $t_{\text{dff}}=\frac{L^{2}}{D_{\textrm{e}}},$ (A.6)
@@
 $t_{\text{des}}=\frac{1}{k_{\textrm{d}}},$ (A.7)
-where $t_{\text{des}}$ is the time scale for desorption and $k_{\textrm{d}}$ is the desorption constant defined in (43).
+where $t_{\text{des}}$ is the time scale for desorption and $k_{\textrm{d}}$ is the desorption constant defined in Equation (43).
 By comparing $t_{\text{dff}}$ and $t_{\text{des}}$, the limiting step for mass transfer in the secondary drying step can be identified, without any simulation or experimental data.
-For numerical comparison, a set of parameter values are obtained from Refs.
-[15, 17], namely $\bar{r}=5$$\times$$10^{-6}$ m, $\epsilon=0.815$, $\tau=1.2$, and $k_{\textrm{d}}=7.8$$\times$$10^{-5}$ s-1.
-The binary coefficient of water in air is obtained from Ref.
-[31], which is $D_{\textrm{g}}=1.97$$\times$$10^{-5}$ m2/s at 256 K and 1 atm.
-In secondary drying, the operating pressure is much lower than 1 atm, and so the actual $D_{\textrm{g}}$ is lower as the diffusion coefficient of gas mixtures at low pressure is inversely proportional to the pressure [31, 46], which makes our analysis more conservative.
+For numerical comparison, a set of parameter values are obtained from [Liapis1994Original, Sheehan1998Modeling], namely $\bar{r}=5$$\times$$10^{-6}$ m, $\epsilon=0.815$, $\tau=1.2$, and $k_{\textrm{d}}=7.8$$\times$$10^{-5}$ s-1.
+The binary coefficient of water in air is obtained from [Mills1995HeatTransfer], which is $D_{\textrm{g}}=1.97$$\times$$10^{-5}$ m2/s at 256 K and 1 atm.
+In secondary drying, the operating pressure is much lower than 1 atm, and so the actual $D_{\textrm{g}}$ is lower as the diffusion coefficient of gas mixtures at low pressure is inversely proportional to the pressure [Mills1995HeatTransfer, Bird2002Transport], which makes our analysis more conservative.
 The length scale for this analysis is set to 0.01 m, a typical vial/sample size (thickness/diameter).
-Substituting these values into (A.6) and (A.7) gives $t_{\text{dff}}=7.5$ s and $t_{\text{des}}=3.6$ h.
+Substituting these values into Equations (A.6) and (A.7) gives $t_{\text{dff}}=7.5$ s and $t_{\text{des}}=3.6$ h.
 Obviously, the time scale for diffusion is much smaller than for desorption by several orders of magnitude, indicating that desorption is the rate-limiting step.
@@
 On the other hand, the time scale for diffusion is only 7.5 s, meaning that diffusion is much faster and so can be safely omitted from the model.
-Even some of the parameters are changed by an order of magnitude, this conclusion does not change.***This analysis holds for lyophilization of unit doses only; for lyophilization of bulk material (e.g., in food applications), the length scale of the product could be much larger, and so the conclusion might be different.
-Finally, there is another mass transfer mechanism that is not commonly discussed or modeled, namely diffusion of water in the solid matrix (not in the pores), i.e., solid-phase diffusion [37].
-Diffusion in the solid matrix is very slow, and so it could limit the mass transport in secondary drying.
-Modeling this phenomenon is challenging as it requires detailed knowledge of the solid matrix, particle size, and locations of bound water in the matrix, which is generally not available.
-A similar approximation to (A.6) can be carried out.
-By using the reported length scale of about $5$$\times$$10^{-7}$ m and diffusion coefficient of water in the solid phase of about $7$$\times$$10^{-16}$ m2/s [37], the time scale for solid-phase diffusion is about 10 min, which is still much lower than for desorption.
+Even some of the parameters are changed by an order of magnitude, this conclusion does not change.
+[^2] This analysis holds for freeze drying of unit doses only; for freeze drying of bulk material (e.g., in food applications), the length scale of the product could be much larger, and so the conclusion might be different.
+[…]
 To summarize, this section details the scale analysis that can be efficiently used to evaluate the effects of various mass transfer mechanisms.
-For lyophilization of pharmaceuticals in unit doses, our analysis shows that desorption of bound water is the rate-limiting step for mass transfer, and so other mechanisms can be omitted from the model, agreeing with both simulation and experimental data in the literature.
+For freeze drying of (bio)pharmaceuticals in unit doses, our analysis shows that desorption of bound water is the rate-limiting step for mass transfer, and so other mechanisms can be omitted from the model, agreeing with both simulation and experimental data in the literature.
 Appendix B Supplemental Equations
 This appendix summarizes equations that are not a crucial part of the model but required for calculating some model parameters.
-In Section 3.2.1, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
+For the freezing model, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
 $\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (B.1)
@@
 $m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (B.3)
-Then (1) can be used as usual.
+Then, Equation (1) can be used as usual.
 During the freezing step, the surface area $A_{r}$ varies with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
@@
 where the product is assumed to be a cylinder.
-For the primary drying model (Section 3.3), the volume of the frozen region is
+For the primary drying model, the volume of the frozen region is
 $V_{\textrm{f}}=\pi d^{2}(H-S).$ (B.5)
@@
 Note that some experimental primary drying studies report $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and/or $H$, in which case the above calculations may not be needed.
-For the secondary drying model (Section 3.4), the volume of the dried region is
+For the secondary drying model, the volume of the dried region is
 $V_{\textrm{e}}=\pi d^{2}H,$ (B.10)
 which is constant because there is no moving boundary (i.e., sublimation front) in this case.
-Appendix C Numerical methods
-This appendix comprehensively describes numerical techniques required for solving the model equations efficiently.
-With the given procedure, our model can be simulated accurately within 1 s on a normal laptop.
-For the freezing model, since the lumped capacity approximation is used, the resulting equations are ordinary differential equations (ODEs), and so the equations can be integrated directly with a proper ODE solver.
-This work uses MATLAB’s ode15s.
-For PDEs in the primary drying and secondary drying models, the method of lines [48] is recommended.
-The method of lines consists of two steps.
-First, the PDEs are discretized spatially to produce a system of ODEs.
-Second, the resulting ODEs are integrated with a proper ODE solver (MATLAB’s ode15s in this work).
-For the primary drying model, first consider the PDE (32). As the equation involves a moving interface $S(t)$, define a new variable
-$\xi=\frac{z-S}{H-S},$ (C.1)
-where $\xi$ is the dimensionless position with respect to the moving interface (aka sublimation front).
-This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$.
-Consequently, the PDE (32) and boundary conditions (38) and (39) become
-$\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (C.2)
-$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (C.3)
-$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),\qquad\xi=0.$ (C.4)
-To spatially discretize the PDE, define
-$\Delta\xi=\frac{1}{n_{z}-1},$ (C.5)
-$\xi=(j-1)\Delta\xi,$ (C.6)
-where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$.
-Discretizing (C.2)–(C.4) using the second-order finite difference scheme results in
-$\begin{split}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{j\Delta\xi-1}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (C.7)
-$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.8)
-$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.9)
-where $T_{j}$ is the product temperature at position $j$.
-The ghost point at $j=n_{z}+1$ in (C.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (C.7) for $j=n_{z}$.
-Similarly, $T_{j=0}$ in (C.9) can be eliminated by substituting $T_{j=0}$ into the governing equation (C.7) for $j=1$.
-The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system.
-Numerical treatment for the secondary drying model is simpler than those described for the primary drying model as there is no moving interface. Define
-$\Delta z=\frac{H}{n_{z}-1},$ (C.10)
-$z=(j-1)\Delta z,$ (C.11)
-where $\Delta z$ is the distance between each grid point, $j=1$ at $z=0$, and $j=n_{z}$ at $z=H$. Discretizing (42), (43), (45), and (46) using the second-order finite difference scheme results in
-$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (C.12)
-$\dfrac{dc_{\textrm{w},j}}{dt}=-f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (C.13)
-$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.14)
-$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.15)
-where the ghost points can be treated similarly as done for the primary drying model.
 Table 6: *
@@
 Symbols
-$A$ area (m2)
+AA area (m2)
 $A_{z}$ cross sectional area of the product (m2)
@@
 […]
-$h_{m}$ mass transfer coefficient (kg/m2$\cdots$)
+[…]
 $j$ integer index for discretization (–)
@@
 0 initial condition
-$1$ surface 1
-$2$ surface 2
+11 surface 1
+22 surface 2
 avg average
@@
 in inert gas
+This article presents the first mechanistic model for continuous lyophilization (aka freeze drying).
+The model accurately describes the key phenomena and critical process parameters in continuous lyophilization, which can be used for process design, optimization, and control.
+Results from this work could drive the transition of (bio)pharmaceutical freeze drying from batch processes to continuous manufacturing.